### v1

# Quasi-Bayes empirical Bayes: a sequential approach to the Poisson compound decision problem

Abstract
The Poisson compound decision problem is a classical problem in statistics, for which parametric and nonparametric empirical Bayes methodologies are available to estimate the Poisson’s means in static or batch domains. In this paper, we consider the Poisson compound decision problem in a streaming or online domain. By relying on a quasi-Bayesian approach, often referred to as Newton’s algorithm, we obtain sequential Poisson’s mean estimates that are of easy evaluation, computationally efficient and with a constant computational cost as data increase, which is desirable for streaming data. Large sample asymptotic properties of the proposed estimates are investigated, also providing frequentist guarantees in terms of a regret analysis. We validate empirically our methodology, both on synthetic and real data, comparing against the most popular alternatives.
Keywords: Empirical Bayes; Newton’s algorithm; Poisson compound decision; quasi-Bayes; streaming data.
1 Introduction
Given $n\geq 1$ observations modeled as independent Poisson random variables $Y_{1},\ldots,Y_{n}$ with means $\theta_{1},\ldots,\theta_{n}$, respectively, the Poisson compound decision problem refers to the estimation of $(\theta_{1},\ldots,\theta_{n})$ under the squared-error loss. Empirical Bayes is a general approach to handle compound decision problems as first suggested by Robbins (1951, 1956); see also Efron (2003) and Zhang (2003). Assuming the parameters $\theta_{i}$’s to be i.i.d as a prior $G$ on $\Theta\subseteq\mathbb{R}^{+}$, the best estimate of $\theta_{i}$ under the squared-error loss, i.e. the Bayes estimate, is the posterior mean, namely for $y\in\mathbb{N}_{0}$
$\hat{\theta}_{G}(y)=E_{G}[\theta_{i}\mid Y_{i}=y]=\frac{\int_{\Theta}\theta\frac{\text{e}^{-\theta}\theta^{y}}{y!}G(\mathrm{d}\theta)}{\int_{\Theta}\frac{\text{e}^{-\theta}\theta^{y}}{y!}G(\mathrm{d}\theta)}=(y+1)\frac{p_{G}(y+1)}{p_{G}(y)},$ (1)
where $p_{G}$ is the probability mass function of the $Y_{i}$’s. The empirical Bayes approach proposes to bypass the knowledge of $G$, which might be unavailable in practice, estimating either $p_{G}$ or ${G}$. Methodologies to estimate $p_{G}$ and ${G}$ are referred to as $f$-modeling and $g$-modeling, respectively (Efron, 2014).
Under $f$-modeling, one obtains an estimate of $p_{G}$ from $(Y_{1},\ldots,Y_{n})$, and then apply (1). For example, Robbins (1956) proposed an estimate of $\hat{\theta}_{G}(y)$ by replacing $p_{G}$ in (1) with the empirical distribution
$\hat{p}_{G}(y)=\sum_{i=1}^{n}I(Y_{i}=y)\qquad y\in\mathbb{N}_{0},$ (2)
where $I$ denotes the indicator function. This is the celebrated nonparametric empirical Bayes estimate (Brown et al., 2013; Polyanskiy and Wu, 2021). Despite its simplicity and (minimax) optimality, Robbins’ method produces unstable estimates in practice; this occurs particularly when $\hat{p}_{G}$ is small or zero (Efron and Hastie, 2021). To alleviate the shortcomings of Robbins’ method, a number of modifications that enforce smoothness of $\hat{p}_{G}(y)$ have been proposed, including $g$-modeling.
Under $g$-modeling, one obtains an estimate of $G$ from $(Y_{1},\ldots,Y_{n})$, and then apply (1). An example is the estimate of $\hat{\theta}_{G}(y)$ obtained by replacing $G$ in (1) with the nonparametric maximum likelihood estimate
$\hat{G}=\operatorname*{arg\,max}_{G}\frac{1}{n}\sum_{i=1}^{n}\log p_{G}(Y_{i}),$ (3)
where the maximization is taken over all distributions on $\mathbb{R}^{+}$ (Kiefer and Wolfowitz, 1956). Much of $g$-modeling involves the estimate (3) (Laird, 1978; Jiang and Zhang, 2009) and nonparametric minimum distance estimates (Jana et al., 2024). Parametric $g$-modeling has been much less developed (Efron and Morris, 1978; Efron, 2016). A fully Bayesian approach is possible by placing a hyperprior distribution on $G$; this is the Bayes empirical Bayes method (Deely and Lindley, 1981).
The Poisson compound decision problem provides a flexible framework for count data, with a broad range of applications where static or batch processing have traditionally been sufficient, e.g. missing-species problems and microarrays (Efron, 2003), insurance analytics (Efron and Hastie, 2021, Chapter 6), deconvolution (Efron, 2016), data confidentiality (Zhang, 2005) and sports analytics (Jana et al., 2024). However, the emergence of streaming count data in some of these applications, as well as in social media analytics, network traffic monitoring, bio-surveillance and healthcare, and e-commerce web analytics, makes it natural and necessary to consider the Poisson compound decision problem in a streaming or online domain, where a timely analysis can enhance decision-making and responsiveness to new information. While it is indeed possible to adapt existing empirical Bayes methodologies to streaming data, the resulting estimates typically suffer from computational inefficiencies, especially when they involve optimizations and posterior computations.
1.1 Our contribution
We propose a principled sequential $g$-modeling methodology for the Poisson compound decision problem: starting from an initial guess $G_{0}$ for the prior $G$, for $n\geq 1$ we update $G_{n-1}$ at each new observation $Y_{n}$ following a procedure of Smith and Makov (1978), often referred to as Newton’s algorithm (Newton et al., 1998; Martin and Ghosh, 2008). This recursive procedure is however regarded not just as an algorithm, but as a quasi-Bayesian learning model (Fortini and Petrone, 2020), which provides a predictive construction for an asymptotically equivalent Bayesian model. The resulting quasi-Bayes $g$-modeling allows to estimate $\hat{\theta}_{G}(y)$, based on $(Y_{1},\dots,Y_{n})$, with the plug-in estimate $\hat{\theta}_{G_{n}}(y)$ obtained from (1) by replacing $G$ with $G_{n}$. Besides its ease of evaluation, the sequential estimate $\hat{\theta}_{G_{n}}(y)$ is computationally efficient and with constant computational cost as data increase, which is desirable for streaming data. By considering Newton’s algorithm as a learning model, it is possible to measure uncertainty of $\hat{\theta}_{G_{n}}(y)$; here, we obtain a Gaussian central limit theorem for $\hat{\theta}_{G_{n}}(y)$, $n\rightarrow+\infty$, leading to asymptotic credible intervals for $\hat{\theta}_{G}(y)$.
We provide the estimate $\hat{\theta}_{G_{n}}(y)$ with asymptotic frequentist guarantees. Assuming the “true” mixing distribution $G$ to be known, say equal to the oracle prior $G^{\ast}$, and that the observations $Y_{n}$’s are i.i.d. from a Poisson mixture model with mixing distribution $G^{*}$, we study the large $n$ asymptotic behavior of the Bayesian regret of $\hat{\theta}_{G_{n}}(y)$, which is defined as the difference between the Bayes risk of $\hat{\theta}_{G_{n}}(y)$ and the Bayes risk of the plug-in estimate $\hat{\theta}_{G^{\ast}}(y)$ obtained from (1) by replacing $G$ with $G^{\ast}$. In particular, we characterize the rate at which the Bayesian regret goes to zero as $n\rightarrow+\infty$.
An empirical validation of our sequential $g$-modeling is presented on synthetic data generated i.i.d from a “true” mixing distribution $G$, and on real data. The estimate $\hat{\theta}_{G_{n}}(y)$, at fixed $n\geq 1$, outperforms the nonparametric empirical Bayes method and competes with the nonparametric maximum likelihood and the minimum squared Hellinger distance methods. Further, $\hat{\theta}_{G_{n}}(y)$ competes with the parametric Bayes empirical Bayes method, assuming a Gamma hyperprior on $G$.
2 Quasi-Bayes empirical Bayes
We consider a stream of observations $(Y_{n})_{n\geq 1}$, such that the $Y_{n}$’s have Poisson distributions with means $\theta_{n}$’s, respectively. To estimate sequentially the parameters $\theta_{n}$’s, we adopt a $g$-modeling methodology that, on a probability space $(\Omega,\mathcal{F},\mathds{P})$, specifies sequentially the distribution of $(Y_{n})_{n\geq 1}$ by giving the conditional distribution of $Y_{n+1}$ given $(Y_{1},\dots,Y_{n})$, for each $n\geq 0$. In particular, denoting by $(\mathcal{G}_{n})_{n\geq 1}$ the natural filtration of $(Y_{n})_{n\geq 1}$, for $n\geq 0$ each conditional distribution with respect to $\mathcal{G}_{n}$ is assumed to be a mixture of Poisson kernels $k(\cdot\mid\theta)$, with $\theta$ defined in a sufficiently fine and wide grid of points $\Theta_{d}=\{\vartheta_{1},\dots,\vartheta_{d}\}$, with $\vartheta_{j}<\vartheta_{j+1}$ ($j=1,\dots,d-1$), namely
$p_{g_{n}}(y)=\mathds{P}(Y_{n+1}=y\mid\mathcal{G}_{n})=\sum_{j=1}^{d}k(y\mid\vartheta_{j})g_{n}(\vartheta_{j})\quad y\in\mathbb{N}_{0},$ (4)
where the mixing distribution $g_{n}$ is defined recursively by means of the following Newton’s algorithm as
$g_{n+1}(\vartheta_{j})=(1-\alpha_{n+1})g_{n}(\vartheta_{j})+\alpha_{n+1}\frac{k(Y_{n+1}\mid\vartheta_{j})g_{n}(\vartheta_{j})}{\sum_{i=1}^{d}k(Y_{n+1}\mid\vartheta_{i})g_{n}(\vartheta_{i})}\qquad j=1,\ldots,d,$ (5)
given an initial guess $g_{0}$ and a sequence $(\alpha_{n})_{n\geq 1}$ of real numbers in $(0,1)$ such that $\sum_{n\geq 1}\alpha_{n}=+\infty$ and $\sum_{n\geq 1}\alpha^{2}_{n}<+\infty$. From (5), $g_{n}$ is updated as $Y_{n+1}$ becomes available, by taking a weighted average, with respect to $\alpha_{n}$, of $g_{n}$ itself and of its posterior distribution based on $Y_{n+1}$. The sequence $(\alpha_{n})_{n\geq 1}$ is referred to as the learning rate, for which a standard choice is $\alpha_{n}=(\alpha+n)^{-\gamma}$ for $\alpha>0$ and $\gamma\in(1/2,1]$. The next result explains in which sense our sequential $g$-modeling is quasi-Bayesian.
Theorem 2.1.
[…]
See Supplementary Material C.1 for the proof of Theorem 2.1. Under (4)-(5), for each $n\geq 0$, the future observations of $(Y_{n})_{n\geq 1}$ are conditionally identically distributed given $\mathcal{G}_{n}$. From Berti et al. (2004), this implies that the $Y_{n}$’s are asymptotically i.i.d., given $\tilde{g}$, with law $p_{\tilde{g}}$, taking on the interpretation of a quasi-Bayesian framework (Fortini and Petrone, 2020). In particular, $p_{\tilde{g}}$ is the random distribution that would be assigned if we had access to an infinitely large sample, from which it is reasonable to consider $\tilde{g}$ as the parameter driving our sequential $g$-modeling. Differently from the quasi-Bayesian framework, the Bayesian framework relies on a prior distribution on $g$, such that the $Y_{n}$’s are i.i.d, given $g$. Under this framework, the computation cost to evaluate the posterior distribution of $g$ is typically cumbersome, and the posterior distribution must be recomputed for each new observation. The quasi-Bayesian framework offers a significantly lower computational cost.
2.1 Sequential estimation
Under the sequential $g$-modeling methodology of Section 2, the empirical Bayes estimate of $\hat{\theta}_{G}(y)$ in (1) that would be assigned if we had access to an infinitely large sample would be of the form
$\hat{\theta}_{\tilde{g}}(y)=(y+1)\frac{p_{\tilde{g}}(y+1)}{p_{\tilde{g}}(y)}\qquad y\in\mathbb{N}_{0}.$ (6)
An estimate of $\hat{\theta}_{G}(y)$ based on the sample $(Y_{1},\ldots,Y_{n})$ is obtained from (6) by replacing $\tilde{g}$ with $g_{n}$, i.e.
$\hat{\theta}_{g_{n}}(y)=(y+1)\frac{p_{g_{n}}(y+1)}{p_{g_{n}}(y)}\qquad y\in\mathbb{N}_{0}.$ (7)
Although the mixing distribution is concentrated on $\Theta_{d}=\{\vartheta_{1},\dots,\vartheta_{d}\}$, we are not restricting the parameter $\theta$ or its estimate to take values in $\Theta_{d}$. Under (4)-(5), $g_{n}$ converges $\mathds{P}$ a.s. to $\tilde{g}$, as $n\rightarrow+\infty$. Then, it follows that for every $y\in\mathbb{N}_{0}$, $\hat{\theta}_{g_{n}}(y)$ converges $\mathds{P}$ a.s. to $\hat{\theta}_{\tilde{g}}(y)$, as $n\rightarrow+\infty$. Next theorem gives an asymptotic Gaussian approximation of the posterior distribution of $\hat{\theta}_{\tilde{g}}(y)$ for $y\in\mathbb{N}_{0}$.
Theorem 2.2.
Let (4)-(5) hold with $(\alpha_{n})_{n\geq 1}$ non increasing and $\sum_{n\geq 1}(\alpha_{n}^{2}/\sum_{k\geq n}\alpha_{n}^{2})^{2}<+\infty$, and let $b_{n}=(\sum_{k\geq n}\alpha_{k}^{2})^{-1}$. For $y\in\mathbb{N}_{0}$, the conditional distribution of $b_{n}^{1/2}(\hat{\theta}_{\tilde{g}}(y)-\hat{\theta}_{g_{n}}(y)$, given $\mathcal{G}_{n}$, converges weakly $\mathds{P}$-a.s. to a centered (i.e., zero-mean) Gaussian kernel with the random variance
$W_{\tilde{g}}(y)=\hat{\theta}^{2}_{\tilde{g}}(y)E_{Z\sim p_{\tilde{g}}}\left[\left(\sum_{j=1}^{d}\frac{k(Z\mid\vartheta_{j})\tilde{g}(\vartheta_{j})}{p_{\tilde{g}}(Z)}\left(\frac{k(y+1\mid\vartheta_{j})}{p_{\tilde{g}}(y+1)}-\frac{k(y\mid\vartheta_{j})}{p_{\tilde{g}}(y)}\right)\right)^{2}\right].$ (8)
See Supplementary Material C.2 for the proof of Theorem 2.2. For the standard choice $\alpha_{n}=(\alpha+n)^{-\gamma}$ for $\alpha>0$ and $\gamma\in(1/2,1]$ then $b_{n}=(2\gamma-1)n^{2\gamma-1}$. Let $W_{g_{n}}(y)$ be defined as in (8), with $g_{n}$ in the place of $\tilde{g}$. Since $g_{n}$ converges to $\tilde{g}$ ($\mathds{P}$-a.s.), then $W_{g_{n}}(y)$ converges to $W_{\tilde{g}}(y)$, $\mathds{P}$-a.s. Therefore, the conditional distribution of $\hat{\theta}_{\tilde{g}}(y)$ given $\mathcal{G}_{n}$ is approximately Gaussian with mean $\hat{\theta}_{g_{n}}(y)$ and variance $W_{g_{n}}(y)$. This provides asymptotic credible intervals for $\hat{\theta}_{\tilde{g}}(y)$, for $y\in\mathbb{N}_{0}$.
Remark 2.3.
A sequential $g$-modeling methodology may be also considered from a frequentist perspective, by assuming i.i.d. observations $(Y_{1},\ldots,Y_{n})$ and applying Newton’s algorithm as a consistent estimate of a “true” $g$ (Tokdar et al., 2009; Martin and Tokdar, 2009). This approach would provide the same estimate $\hat{\theta}_{g_{n}}(y)$ of $\hat{\theta}_{G}(y)$ as in (7), and an upper bound for the rate of convergence (see Section 2.2). However, to the best of our knowledge, going beyond point estimation, for example by building asymptotic confidence intervals, is not feasible based on asymptotic results available for Newton’s algorithm. Indeed, these are based on stochastic approximation (Martin and Ghosh, 2008), but central limit theorems in this field does not apply to Newton’s algorithm.
2.2 Frequentist guarantees
The “true” Poisson compound model for $(Y_{n})_{n\geq 1}$ is the model under which, for any $n\geq 1$: i) $Y_{n}$ given $\theta^{\ast}_{n}$ is Poisson distributed with parameter $\theta^{\ast}_{n}$; ii) the $\theta^{\ast}_{n}$’s are independent and identically distributed from the “true” prior distribution, say the oracle prior $G^{\ast}$ with probability mass function $g^{\ast}$ on $\Theta_{d}$. If $G^{\ast}$ was known, then we could estimate $\hat{\theta}_{G}(y)$ by $\hat{\theta}_{g^{*}}(y)$, namely the plug-in estimate obtained from (1) by replacing $g$ with $g^{\ast}$. Under the “true” model, the Bayes risk of $\hat{\theta}_{g^{*}}(y)$ is
$\mathcal{R}(g^{*},\hat{\theta}_{g^{*}})=E^{*}(\hat{\theta}_{g^{*}}(Y_{n})-\theta_{n}^{*})^{2}=\sum_{j=1}^{d}\sum_{y\in\mathbb{N}_{0}}^{\infty}(\hat{\theta}_{g^{*}}(y)-\vartheta_{j})^{2}k(y\mid\vartheta_{j})g^{*}(\vartheta_{j}),$
with $E^{*}$ denoting the expected value with respect to the true model. Not knowing the oracle $g^{*}$, we employ the estimate $\hat{\theta}_{g_{n}}(y)$ in (7). In particular, under the “true” model, the Bayes risk of $\hat{\theta}_{g_{n}}(y)$ is
$\mathcal{R}(g^{*},\hat{\theta}_{g_{n}})=E^{*}(\hat{\theta}_{g_{n}}(Y_{n})-\theta^{*}_{n})^{2}=\mathcal{R}(g^{*},\hat{\theta}_{g^{*}})+E^{*}(\hat{\theta}_{g_{n}}(Y_{n})-\hat{\theta}_{g^{*}}(Y_{n}))^{2}.$
The Bayesian regret (of not knowing $g^{*}$ and using $g_{n}$ instead) is the difference between $\mathcal{R}(g^{*},\hat{\theta}_{g_{n}})$ and $\mathcal{R}(g^{*},\hat{\theta}_{g^{*}})$, i.e.
$\text{Regret}(g_{n},g^{*})=E^{*}((\hat{\theta}_{g_{n}}(Y_{n})-\hat{\theta}_{g^{*}}(Y_{n}))^{2}=\sum_{j=1}^{d}\sum_{y\in\mathbb{N}_{0}}(\hat{\theta}_{g_{n}}(y)-\hat{\theta}_{g^{*}}(y))^{2}k(y\mid\vartheta_{j})g^{*}(\vartheta_{j}).$ (9)
See Efron (2019) for details on (9). The next theorem provides the large $n$ asymptotic behaviour of the regret.
Theorem 2.4.
Let $g_{n}$ be as in (5) with $\alpha_{n}=(\alpha+n)^{-\gamma}$ for $\alpha>0$ and $\gamma\in(1/2,1)$, and let $g^{\ast}$ be the probability mass function of the oracle prior $G^{\ast}$. Denote by $\mathds{P}^{*}$ a probability measure under which the “true” Poisson compound model for $(Y_{n})_{n\geq 1}$ holds true. Then, for every $\delta<2-1/\gamma$, it holds that
$\text{Regret}(g_{n},g^{*})=o(n^{-\delta})\quad\mathds{P}^{\ast}\,\text{-}\,a.s.$
See Supplementary Material C.3 for the proof of Theorem 2.4. The regret analysis of Theorem 2.4 is provided under the standard choice $\alpha_{n}=(1/(\alpha+n))^{\gamma}$, though other learning rates may be considered.
3 Numerical illustrations
We validate empirically our sequential $g$-modeling on synthetic and real data. We compare the quasi-Bayes (NP-QB) estimate $\hat{\theta}_{g_{n}}$ of $\hat{\theta}_{G}$, at fixed $n\geq 1$, with the estimates obtained by the following methods: i) nonparametric empirical Bayes (NP-EB) (Robbins, 1956); ii) nonparametric maximum likelihood (NP-ML) (Jiang and Zhang, 2009); iii) nonparametric minimum squared Hellinger distance (NP-MD) (Jana et al., 2024); parametric Bayes empirical Bayes (P-EB) with a Gamma hyperprior on $G$ (Deely and Lindley, 1981). NP-EB is the sole to produce a closed-form estimate, i.e. the estimate obtained by replacing $p_{G}$ in (1) with the empirical distribution (2). NP-ML, N-PMD and P-EB require exact or approximate optimization algorithms: i) NP-ML and N-PMD estimates are obtained from Jana et al. (2024, Algorithm 1); ii) the P-EB estimate is obtained from Deely and Lindley (1981, Equation 4.4), with the parameters of the Gamma hyperprior estimated by maximum (marginal) likelihood. The empirical performance of the above estimates is evaluated by the root mean squared error (RMSE) and the mean absolute deviation (MAD).
3.1 Synthetic data
We set $n=50,100,200,500$, and generate $(Y_{1},\ldots,Y_{n})$ under the Poisson compound model, with $(\theta_{1},\ldots,\theta_{n})$ i.i.d from a Gamma prior $G$ with scale $2$ and shape $4$. For NP-QB, we set Newton’s algorithm as follows: i) $\theta$ takes value in $\Theta=\{0.01,0.02,\ldots,99.99,100\}$; ii) the number $d$ of mixture components is the cardinality of $\Theta$, i.e. $d=10,000$; iii) the initial guess $g_{0}$ is the Uniform distribution on $\Theta$; iv) the learning rate is $\alpha_{n}=(1+n)^{-1}$. Results are displayed in Figure 1 and Table 1.
Figure 1: Data $Y_{i}$’s against the true $\theta_{i}$’s (black), and estimates of the $\theta_{i}$’s: NP-EB (green), NP-ML (cyan), NP-MD (magenta), P-EB (red) and NP-QB (blue).
Table 1: Values of RMSE and MAD for the NP-EB, NP-ML, NP-MD, P-EB and NP-QB estimates.
NP-EB NP-ML NP-MD P-EB NP-QB
RMSE   $n=100$ 6.184 2.230 2.253 2.110 2.235
RMSE   $n=500$ 3.737 2.238 2.228 2.244 2.236
MAD     $n=100$ 4.100 1.722 1.731 1.632 1.719
MAD     $n=500$ 2.459 1.773 1.768 1.782 1.774
[…]
3.2 Real data
We consider a dataset of tweets with at least 50 retweets on Twitter in the period October 7 to November 6, 2011 (30 days). The dataset is available at https://snap.stanford.edu/seismic/, and it consists of $166,783$ tweets that generate $34,617,704$ retweets. The data provides information on the tweet ID, the tweet post time, the retweet post time and the number of followers of the poster/retweeter. For each tweet, we focus on the number of retweets generated within $30$ seconds from the tweet post time; in principle, the faster a retweet is, the more likely it was generated by a chatbot artificial intelligence rather than a human. The count data of fast retweets arrive at high-velocity, thus motivating the need of an online fast processing. From the first two rows of Table 2, $40,259$ tweets have 00 retweets in $30$ seconds, $28,339$ tweets have $1$ retweet in $30$ seconds, etc. Interest is in estimating the number of retweets expected for a tweet. Table 2 displays the estimates obtained through the application of NP-EB, NP-ML, NP-MD, P-FB and NP-QB.
Table 2: Counts $n_{y}$ of the number of tweets that have $y$ retweets within $30$ seconds from the tweet post time, for $y=0,1,\ldots,7$, and corresponding estimates of the number of retweets expected.
$y$ 0 1 2 3 4 5 6 7
$n_{y}$ 40,259 28,339 21,581 16,479 12,130 9,238 7,193 5,464
NP-EB 0.704 1.523 2.291 2.944 3.808 4.672 5.317 6.324
NP-ML 0.391 2.134 2.931 3.697 4.459 5.329 6.334 7.424
NP-MD 0.750 2.198 3.142 4.149 4.986 5.694 6.390 7.163
P-EB 0.445 1.335 2.226 3.116 4.007 4.897 5.788 6.678
NP-QB 0.607 0.818 1.632 3.076 4.007 4.519 5.023 5.707
We refer to Supplementary Material D.2 for a version of Table 2 with more values of $y$, which shows that NP-QB competes with NP-ML, NP-MD and P-EB, and all these methods outperform NP-EB. The Supplementary Material D.2 includes another application to same Twitter dataset, but with respect to the number of followers generated by fast retweets. See also Supplementary Material D.3 for an application to the benchmark auto accident dataset (Efron and Hastie, 2021, Table 6.1).
4 Discussion
The Supplementary Material E contains an extension of our sequential $g$-modeling to the multidimensional Poisson compound decision problem (Brown et al., 1985; Johnstone, 1986). Further extensions may be considered with respect to compound decision problems beyond the Poisson setting, with applications to high-dimensional problem, e.g. exponential families (Zhang, 2003). Sequential $f$-modeling is also an interesting direction for future research (Brown et al., 2013).
Appendix
B Preliminaries
B.1 Notation
[…]
B.2 Almost sure conditional convergence.
Given a probability space $(\Omega,\mathcal{F},\mathds{P})$ and a Polish space $\mathbb{X}$ with its Borel sigma-algebra $\mathcal{X}$, a kernel on $\mathbb{X}$ is a function $K:\mathcal{X}\times\Omega$ satisfying: i) for every $\omega\in\Omega$, $K(\cdot,\omega)$ is a probability measure on $\mathcal{X}$; ii) for each $B\in\mathcal{X}$, the function $K(B,\cdot)$ is measurable with respect to $\mathcal{F}$. A kernel $K$ on $\mathbb{R}$ is called a Gaussian kernel if $K(\cdot,\omega)$ is a Gaussian distribution with mean $\mu(\omega)$ and variance $\sigma^{2}(\omega)$, where $\mu$ and $\sigma^{2}$ are random variables defined on $(\Omega,\mathcal{F},\mathds{P})$. We denote the Gaussian kernel by $\mathcal{N}(\mu,\sigma^{2})$ and interpret the Gaussian distribution with zero variance as the degenerate law centered on the mean. A kernel $K$ on the Polish space $C(\mathbb{R})$, with the topology of uniform convergence on compact sets, is called a Gaussian process kernel if $K(\cdot,\omega)$ is the distribution of a Gaussian process in $C(\mathbb{R})$, with mean function $\mu(\omega)\in C(\mathbb{R})$ and covariance function $R(\omega)\in C(\mathbb{R}\times\mathbb{R})$. If $\mu(\omega)=0$ for all $\omega$, then the Gaussian process kernel is said to be centered. We extend the definition of almost sure conditional convergence for real sequences (Crimaldi, 2009), to sequences taking values in Polish spaces.
Definition B.1.
Let $(X_{n})$ be a sequence of random variables adapted to a filtration $(\mathcal{G}_{n})$ and taking values in a Polish space $\mathbb{X}$ with its Borel sigma-algebra $\mathcal{X}$. For every $n\geq 0$, let $K_{n}$ denote a regular version of the conditional distribution of $X_{n+1}$, given $\mathcal{G}_{n}$. If there exists a kernel $K$ such that the sequence $(K_{n}(\cdot,\omega))_{n}$ converges weakly to $K(\cdot,\omega)$ for almost every $\omega\in\Omega$, then we say that the sequence $(X_{n})$ converges to $K$ in the sense of almost-sure conditional convergence with respect to the filtration $(\mathcal{G}_{n})$.
The results are the main tool in the proof of Theorem 2.2.
Theorem B.2 (Crimaldi (2009), Theorem A.1).
For each $n\geq 1$, let $(M_{n,j})_{j\geq 1}$ be a real valued martingale with respect to the filtration $(\mathcal{F}_{n,j})_{j\geq 1}$, satisfying $M_{n,0}=0$, and converging in $L^{1}$ to a random variable $M_{n,\infty}$. Set
$X_{n,j}=M_{n,j}-M_{n,j-1}\mbox{ for }j\geq 1,\quad U_{n}=\sum_{j\geq 1}X_{n,j}^{2},\quad X_{n}^{*}=\sup_{j\geq 1}|X_{n,j}|.$
Assume that
  • (a)
    $(X_{n}^{*})_{n}$ is dominated in $L^{1}$ and converges to zero a.s.
  • (b)
    $(U_{n})_{n}$ converges a.s. to a non-negative random variable $U$.
Then the sequence $(M_{n,\infty})$ converges to the Gaussian kernel $\mathcal{N}(0,U)$ in the sense of almost-sure conditional convergence with respect to the filtration $(\mathcal{F}_{n,1})$.
Remark B.3.
Requesting that the almost sure limit of $(U_{n})$ is non-negative guarantees the well-definedness of the kernel $\mathcal{N}(0,U)$ for every $\omega\in\Omega$. Theorem A.1 in Crimaldi (2009) specifies that $U$ should be a “positive random variable”. Although not explicitly stated, this requirement should be understood as $U\geq 0$, as becomes evident upon a careful examination of the proof.
Theorem B.4 (Crimaldi et al. (2016), Lemma 4.1).
Let $(Z_{n})$ be a sequence of real valued random variables adapted to the filtration $(\mathcal{G}_{n})$ and such that $E(Z_{n+1}\mid\mathcal{G}_{n})\rightarrow Z$ a.s. for some random variable $Z$. Moreover, let $(a_{n})$ and $(b_{n})$ be sequences of real numbers such that
$b_{n}\uparrow+\infty,\quad\sum_{k=1}^{\infty}\frac{E(Z_{k}^{2})}{a_{k}^{2}b_{k}^{2}}<+\infty.$
Then we have:
  • (a)
    If $\frac{1}{b_{n}}\sum_{k=1}^{n}\frac{1}{a_{k}}\rightarrow c$ for some constant $c$, then $\frac{1}{b_{n}}\sum_{k=1}^{n}\frac{Z_{k}}{a_{k}}\rightarrow cZ$ a.s.
  • (b)
    If $b_{n}\sum_{k\geq n}\frac{1}{a_{k}b_{k}^{2}}\rightarrow c$ for some constant $c$, then ${b_{n}}\sum_{k\geq n}\frac{Z_{k}}{a_{k}b_{k}^{2}}\rightarrow cZ$ a.s.
B.3 Delta method for almost-sure conditional convergence
In this section, we extend the delta method to almost sure conditional convergence. Let $\mathbf{h}$ be a function defined on a convex set $\Delta\in\mathcal{B}(\mathbb{R}^{d})$ with values in $\mathbb{R}^{p}$. The function $\mathbf{h}$ is said to be uniformly differentiable on $\Delta$ if there exists, for every $\mathbf{g}^{*}\in\Delta$ a linear function $\mathbf{D}\mathbf{h}_{\mathbf{g}^{*}}:\Delta\rightarrow\mathbb{R}^{p}$ such that the following condition holds: for every $\epsilon>0$ there exists $\delta$ such that
$\frac{||\mathbf{h}(\mathbf{g})-\mathbf{h}(\mathbf{g}^{*})-\mathbf{D}\mathbf{h}_{\mathbf{g}^{*}}(\mathbf{g}-\mathbf{g}^{*})||}{||\mathbf{g}-\mathbf{g}^{*}||}<\epsilon\quad\mbox{ for every }\quad\mathbf{g}\neq\mathbf{g}^{*},||\mathbf{g}-\mathbf{g}^{*}||<\delta.$ (B.1)
Theorem B.5.
Let $(\mathbf{g}_{n})$ be a sequence of random vectors taking values in a convex set $\Delta\in\mathcal{B}(\mathbb{R}^{d})$ with values in $\mathbb{R}^{p}$, adapted to the filtration $(\mathcal{G}_{n})$, and let $\tilde{\mathbf{g}}$ be a random vector taking values in $\Delta$. Assume that there exist an increasing and diverging sequence $(r_{n})$ of positive numbers and a kernel $K(\cdot,\omega)$ tight with respect to $\omega$, $P$-a.s., such that $r_{n}(\mathbf{g}_{n}-\tilde{\mathbf{g}})$ converges in the sense of almost sure conditional convergence to $K(\cdot,\omega)$. If $\mathbf{h}$ is a uniformly differentiable function on $\Delta$, then $r_{n}(\mathbf{h}(\mathbf{g}_{n})-\mathbf{h}(\tilde{\mathbf{g}}))$ converges in the sense of almost sure conditional convergence to the kernel $K(\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}^{-1}(\cdot),\omega)$.
The proof is split into several lemmas.
Lemma B.6.
Under the assumptions of Theorem B.5, for every $\epsilon>0$
$P_{n}\left(\frac{||\mathbf{h}(\mathbf{g}_{n})-\mathbf{h}(\tilde{\mathbf{g}})-\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}(\mathbf{g}_{n}-\tilde{\mathbf{g}})||}{||\mathbf{g}_{n}-\tilde{\mathbf{g}}||}>\epsilon\right)\rightarrow 0\quad P-a.s.$
as $n\rightarrow\infty$, where $P_{n}(\cdot)=P(\cdot\mid\mathcal{G}_{n})$.
Proof.
Fix $\epsilon>0$ and let $\delta$ be such that (B.1) holds. Then
$P_{n}\left(\frac{||\mathbf{h}(\mathbf{g}_{n})-\mathbf{h}(\tilde{\mathbf{g}})-D\mathbf{h}_{\tilde{\mathbf{g}}}(\mathbf{g}_{n}-\tilde{\mathbf{g}})||}{||\mathbf{g}_{n}-\tilde{\mathbf{g}}||}>\epsilon\right)$ $\leq P_{n}(||\mathbf{g}_{n}-\tilde{\mathbf{g}}||\geq\delta)\leq P_{n}(r_{n}||\mathbf{g}_{n}-\tilde{\mathbf{g}}||\geq r_{n}\delta).$
Fix $\eta$, and let $M$ be such that $K([-M,M]^{d},\omega)>1-\eta$ for almost every $\omega$. Then, there exists $n_{0}=n_{0}(\omega)$ such that, for $\omega$ in a set of probability one, and $n>n_{0}(\omega)$,
$P_{n}(r_{n}||\mathbf{g}_{n}-\tilde{\mathbf{g}}||\geq M)(\omega)<2\eta.$
It follows that
$\limsup_{n}P_{n}\left(\frac{||\mathbf{h}(\mathbf{g}_{n})-\mathbf{h}(\tilde{\mathbf{g}})-\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}(\mathbf{g}_{n}-\tilde{\mathbf{g}})||}{||\mathbf{g}_{n}-\tilde{\mathbf{g}}||}>\epsilon\right)<2\eta\quad\quad P-a.s.$
The thesis follows from the arbitrariness of $\eta$. ∎
Lemma B.7.
Under the assumptions of Theorem B.5, for every $\epsilon>0$,
$P_{n}\left(||r_{n}(\mathbf{h}(\mathbf{g}_{n})-\mathbf{h}(\tilde{\mathbf{g}}))-r_{n}\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}(\mathbf{g}_{n}-\tilde{\mathbf{g}})||>\epsilon\right)\rightarrow 0\quad P-a.s.$
as $n\rightarrow\infty$.
Proof.
It can be proved as in Lemma B.6 that, for every $\eta>0$, there exists $n_{0}=n_{0}(\omega)$ such that, for $\omega$ in a set of probability one and $n\geq n_{0}(\omega)$,
$P_{n}(r_{n}||\mathbf{g}_{n}-\tilde{\mathbf{g}}||\geq M)(\omega)<\eta.$
Thus, for $n>n_{0}(\omega)$, it holds, $P$-a.s. that
$P_{n}\left(||r_{n}(\mathbf{h}(\mathbf{g}_{n})-\mathbf{h}(\tilde{\mathbf{g}}))-r_{n}\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}(\mathbf{g}_{n}-\tilde{\mathbf{g}})||>\epsilon\right)(\omega)$
$\quad\leq P_{n}\left(\frac{||\mathbf{h}(\mathbf{g}_{n})-\mathbf{h}(\tilde{\mathbf{g}})-\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}(\mathbf{g}_{n}-\tilde{\mathbf{g}})||}{||\mathbf{g}_{n}-\tilde{\mathbf{g}}||}\;r_{n}||\mathbf{g}_{n}-\tilde{\mathbf{g}}||>\epsilon\right)(\omega)$
$\quad\leq P_{n}(r_{n}||\mathbf{g}_{n}-\tilde{\mathbf{g}}||\geq M)(\omega)+P_{n}\left(M\frac{||\mathbf{h}(\mathbf{g}_{n})-\mathbf{h}(\tilde{\mathbf{g}})-\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}(\mathbf{g}_{n}-\tilde{\mathbf{g}})||}{||\mathbf{g}_{n}-\tilde{\mathbf{g}}||}>\epsilon\right)(\omega)$
$\quad\leq\eta+P_{n}\left(\frac{||\mathbf{h}(\mathbf{g}_{n})-\mathbf{h}(\tilde{\mathbf{g}})-\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}(\mathbf{g}_{n}-\tilde{\mathbf{g}})||}{||\mathbf{g}_{n}-\tilde{\mathbf{g}}||}>\frac{\epsilon}{M}\right)(\omega).$
The thesis follows from Lemma B.6 and the arbitrariness of $\eta$.
of Theorem B.5.
Let $N^{c}$ be an event such that $P(N^{c})=1$, $\{K(\cdot,\omega):\omega\in N^{c}\}$ is tight, $P_{n}(r_{n}(\mathbf{g}_{n}-\tilde{\mathbf{g}})\in\cdot)(\omega)\rightarrow K(\cdot,\omega)$ and $P_{n}(||r_{n}(\mathbf{h}(\mathbf{g}_{n})-\mathbf{h}(\tilde{g}))-r_{n}\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}(\mathbf{g}_{n}-\tilde{\mathbf{g}})||>\epsilon)(\omega)\rightarrow 0$ for every $\epsilon>0$. Fix $\omega\in N^{c}$ and let $A\in\mathcal{B}(\mathbb{R}^{p})$ be such that $K(\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}^{-1}(\partial A),\omega)=0$. For every $\epsilon>0$, let $A^{\epsilon}=\{\mathbf{x}\in\mathbb{R}^{p}:{\rm d}(\mathbf{x},A)\leq\epsilon\}$. Then
$\limsup_{n}P_{n}(r_{n}(\mathbf{h}(\mathbf{g}_{n})-\mathbf{h}(\tilde{\mathbf{g}}))\in A)(\omega)$
$\leq\limsup_{n}P_{n}(||r_{n}(\mathbf{h}(\mathbf{g}_{n})-\mathbf{h}(\tilde{\mathbf{g}}))-r_{n}\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}(\mathbf{g}_{n}-\tilde{\mathbf{g}})||>\epsilon)(\omega)$
$+\limsup_{n}P_{n}(\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}(r_{n}(\mathbf{g}_{n}-\tilde{\mathbf{g}}))\in A^{\epsilon})(\omega)$
$\leq K(\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}^{-1}(A^{\epsilon}),\omega).$
Analogously,
$\limsup_{n}P_{n}(r_{n}(\mathbf{h}(\mathbf{g}_{n})-\mathbf{h}(\tilde{\mathbf{g}}))\in A^{c})(\omega)\leq K(\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}^{-1}((A^{c})^{\epsilon}),\omega).$
Thus,
$\liminf_{n}P_{n}(r_{n}(\mathbf{h}(\mathbf{g}_{n})-\mathbf{h}(\tilde{\mathbf{g}}))\in A)(\omega)\geq K(\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}^{-1}(A_{\epsilon}),\omega),$
where $A_{\epsilon}$ is the complement of $(A^{c})^{\epsilon}$. Thus, for every $\epsilon>0$,
$K(\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}^{-1}(A_{\epsilon}),\omega)$ $\leq\liminf_{n}P_{n}(r_{n}(\mathbf{h}(\mathbf{g}_{n})-\mathbf{h}(\tilde{\mathbf{g}}))\in A)$
$\leq\limsup_{n}P_{n}(r_{n}(\mathbf{h}(\mathbf{g}_{n})-\mathbf{h}(\tilde{\mathbf{g}}))\in A)(\omega)$
$\leq K(\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}^{-1}(A^{\epsilon}),\omega).$
Since $K(\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}^{-1}(\partial A,\omega)=0$, then $K(\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}^{-1}(A^{\epsilon}),\omega)-K(\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}^{-1}(A_{\epsilon}),\omega)\rightarrow 0$, as $\epsilon\rightarrow 0$. Thus $\lim_{n}P_{n}(r_{n}(\mathbf{h}(\mathbf{g}_{n})-\mathbf{h}(\tilde{\mathbf{g}}))\in A)(\omega)=K(\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}^{-1}(A),\omega).$
C Proofs for Section 2
C.1 Proof of Theorem 2.1
For every $n\geq 0$ and every $j=1,\dots,d$,
$E(g_{n+1}(\vartheta_{j})\mid\mathcal{G}_{n})=(1-\alpha_{n+1})g_{n}(\vartheta_{j})+\alpha_{n+1}\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{j})g_{n}(\vartheta_{j})}{\sum_{i=1}^{d}k(y\mid\vartheta_{i})g_{n}(\vartheta_{i})}p_{g_{n}}(y)=g_{n}(\vartheta_{j}).$
Thus, the sequence $(g_{n}(\vartheta_{j}))$ is a bounded martingale with respect to the filtration $(\mathcal{G}_{n})$. Denoting by $\tilde{g}(\vartheta_{j})$ its $\mathds{P}$-a.s. limit, we have that $\tilde{g}$ is a probability mass function on $\Theta_{d}$, $g_{n}(\vartheta_{j})=E(\tilde{g}(\vartheta_{j})\mid\mathcal{G}_{n})$ for every $n\geq 0$ and $j=1,\dots,d$, and, for every $y\in\mathbb{N}_{0}$, $p_{g_{n}}(y)$ converges $\mathds{P}$-a.s. to $p_{\tilde{g}}(y)$. Since $\tilde{g}$ is a probability mass function on $\Theta_{d}$, then $p_{\tilde{g}}$ is a probability mass function on $\mathbb{N}_{0}$. Moreover, for every $y\in\mathbb{N}_{0}$,
$E(p_{g_{n+1}}(y)\mid\mathcal{G}_{n})=E(\sum_{j=1}^{d}k(y\mid\vartheta_{j})g_{n+1}(\vartheta_{j})\mid\mathcal{G}_{n})=p_{g_{n}}(y).$
Thus, $(p_{g_{n}}(y))$ is also a martingale with respect to $\mathcal{G}_{n}$, and $p_{g_{n}}(y)=E(p_{\tilde{g}}(y)\mid\mathcal{G}_{n})$ for every $y\in\mathbb{N}_{0}$ and $n\geq 0$. It follows that, for every $y\in\mathbb{N}_{0}$, every $n\geq 0$ and every $k\geq 1$,
$\mathds{P}(Y_{n+k}=y\mid\mathcal{G}_{n})=E(\mathds{P}(Y_{n+k}=y\mid\mathcal{G}_{n+k-1}\mid\mathcal{G}_{n})=E(p_{g_{n+k-1}}(y)\mid\mathcal{G}_{n})=p_{g_{n}}(y)=E(p_{\tilde{g}}(y)\mid\mathcal{G}_{n}),$
[…]
C.2 Proof of Theorem 2.2
For every $z\in\mathbb{N}_{0}$ let $\tilde{g}(\vartheta_{j}\mid z)=\frac{k(z\mid\vartheta_{j})\tilde{g}(\vartheta_{j})}{p_{\tilde{g}}(z)}$. Let us define the random matrices $\mathbf{V}_{\tilde{g}}$ and $\mathbf{V}_{n}$ ($n\geq 0$) with $(i,j)$ element
$\mathbf{V}_{\tilde{g},i,j}=\sum_{z\in\mathbb{N}_{0}}\tilde{g}(\vartheta_{i}\mid z)\tilde{g}(\vartheta_{j}\mid z)p_{\tilde{g}}(z)-\tilde{g}(\vartheta_{i})\tilde{g}(\vartheta_{j}).$ (C.2)
$\mathbf{V}_{n,i,j}=\sum_{z\in\mathbb{N}_{0}}g_{n}(\vartheta_{i}\mid z)g_{n}(\vartheta_{j}\mid z)p_{g_{n}}(z)-g_{n}(\vartheta_{i})g_{n}(\vartheta_{j}),$ (C.3)
for $i,j=1,\dots,d-1$.
Lemma C.1.
Let $\mathbf{V}_{\tilde{g}}$ and $\mathbf{V}_{n}$ be defined as in (C.2) and (C.3), respectively. Then, $\mathds{P}$-a.s., $(\mathbf{V}_{n})$ is a sequence of uniformly bounded, positive definite random matrices, converging to $\mathbf{V}_{\tilde{g}}$ as $n\rightarrow\infty$.
Proof.
Since, for every $i,j$
$|\mathbf{V}_{n,i,j}|\leq 1,$
then the sequence $(\mathbf{V}_{n})$ is uniformly bounded. Moreover for every vector $\mathbf{u}\in\mathbb{R}^{d-1}$,
$\mathbf{u}^{T}\mathbf{V}_{n}\mathbf{u}$ $=\sum_{i,j=1}^{d-1}\sum_{z\in\mathbb{N}_{0}}\mathbf{u}_{i}(g_{n}(\vartheta_{i}\mid z)-g_{n}(\vartheta_{i}))\mathbf{u}_{j}(g_{n}(\vartheta_{j}\mid z)-g(\vartheta_{j}))p_{g_{n}}(z)$
$=\sum_{z\in\mathbb{N}_{0}}\left(\sum_{i=1}^{d-1}\mathbf{u}_{i}(g_{n}(\vartheta_{i}\mid z)-g_{n}(\vartheta_{i}))\right)^{2}p_{g_{n}}(z)\geq 0,$
$\mathds{P}$-a.s. Hence, $\mathds{P}$-a.s., $\mathbf{V}_{n}$ is positive semi-definite. We now prove that it is $\mathds{P}$-a.s. positive definite. By contradiction, suppose that there exists a random vector $\mathbf{u}$ such that for $\omega$ in a set $N^{c}$ with positive probability $\mathbf{u}(\omega)\neq\mathbf{0}$ and $\mathbf{u}^{T}(\omega)\mathbf{V}_{n}(\omega)\mathbf{u}(\omega)=0$. Fix $\omega\in N^{c}$ and, without loss of generality, assume that $p_{g_{n}(\omega)}(z)>0$ for every $z\in\mathbb{N}_{0}$. Then, we can write (omitting $\omega$ for simplicity)
$\sum_{i=1}^{d-1}\mathbf{u}_{i}(g_{n}(\vartheta_{i}\mid z)-g_{n}(\vartheta_{i}))=0,$
for every $z\in\mathbb{N}_{0}$. It follows that for every $z\in\mathbb{N}_{0}$,
$\sum_{i=1}^{d-1}\mathbf{u}_{i}e^{-\vartheta_{i}}\vartheta_{i}^{z}g_{n}(\vartheta_{i})=\sum_{i=1}^{d}e^{-\vartheta_{i}}\vartheta_{i}^{z}g_{n}(\vartheta_{i})\sum_{i=1}^{d-1}\mathbf{u}_{i}g_{n}(\vartheta_{i})$
Define $u(\vartheta_{i})=\mathbf{u}_{i}$ for $i=1,\dots,d-1$ and $u(\vartheta_{d})=0$. Then
$\sum_{i=1}^{d}u(\vartheta_{i})e^{-\vartheta_{i}}\vartheta_{i}^{z}g_{n}(\vartheta_{i})=\sum_{i=1}^{d}e^{-\vartheta_{i}}\vartheta_{i}^{z}g_{n}(\vartheta_{i})\sum_{i=1}^{d}u(\vartheta_{i})g_{n}(\vartheta_{i}),$
which implies that, for every $k=0,1,\dots$,
$\sum_{i=1}^{d}u(\vartheta_{i})e^{k\vartheta_{i}}g_{n}(\vartheta_{i})=\sum_{i=1}^{d}e^{k\vartheta_{i}}g_{n}(\vartheta_{i})\sum_{i=1}^{d}u(\vartheta_{i})g_{n}(\vartheta_{i}).$ (C.4)
To demonstrate that this leads to a contradiction, consider a random variable $\tilde{\theta}$ with a probability mass function $g_{n}$ defined on $\Theta_{d}$. Since $g_{n}(\vartheta_{i})>0$ for every $i$, the random variables $(e^{k\tilde{\theta}})_{k\geq 0}$ are linearly independent. Furthermore, $u(\tilde{\theta})$ cannot be constant because $u(\vartheta_{d})=0$ and $\mathbf{u}$ is not the zero vector. Therefore, equation (C.4) can not hold.
We now prove that $\mathbf{V}_{n}$ converge to $\mathbf{V}_{\tilde{g}}$, $P$-a.s. Since the random matrices $\mathbf{V}_{n}$ and $\mathbf{V}_{\tilde{g}}$ have equal finite dimensions, then the convergence follows if we prove that
$\mathbf{V}_{n,i,j}\rightarrow\mathbf{V}_{\tilde{g},i,j}\quad P-a.s.,$
for every $i$ and $j$. Since $\mathbf{V}_{n,i,j}$ is a continuous function of $g_{n}$ and since $g_{n}$ converges to $\tilde{g}$, $P$-a.s, then $\mathbf{V}_{n,i,j}$ converges to $\mathbf{V}_{\tilde{g},i,j}$, $P$-a.s. ∎
Theorem C.2.
If $(b_{n})$ is a sequence of strictly positive numbers with
$b_{n}\uparrow+\infty,\quad\sum_{n\geq 1}\alpha_{n}^{4}b_{n}^{2}<+\infty,\quad b_{n}\sup_{k\geq n}\alpha_{k}^{2}\rightarrow 0,\quad b_{n}\sum_{k\geq n}\alpha_{k}^{2}\rightarrow 1,$ (C.5)
then, as $n\rightarrow+\infty$, $b_{n}^{1/2}(\mathbf{g}_{n}-\tilde{\mathbf{g}})$ converges, in the sense of almost-sure conditional convergence with respect to the filtration $(\mathcal{G}_{n})$, to a centered Gaussian kernel with random covariance matrix $\mathbf{V}_{\tilde{g}}$ as in (C.2).
Proof.
The proof is based on Theorem B.2. Fix a vector $\mathbf{u}$ in $\mathbb{R}^{d-1}$ such that $||\mathbf{u}||=1$. For each $n\geq 0$, define
$\mathcal{F}_{n,0}=\mathcal{G}_{n},\quad\mathcal{F}_{n,j}=\mathcal{G}_{n+j-1}\mbox{ for }j\geq 1,$
$M_{n,0}=0,\quad M_{n,j}=b_{n}^{1/2}(\langle g_{n},\mathbf{u}\rangle-\langle g_{n+j-1},\mathbf{u}\rangle)\mbox{ for }j\geq 1.$
For every $n\geq 0$, the sequence $(M_{n,j})_{j\geq 0}$ is a bounded martingale with respect to the filtration $(\mathcal{F}_{n,j})_{j\geq 0}$, converging in $L^{1}$ to
$M_{n,\infty}=b_{n}^{1/2}\langle\mathbf{g}_{n}-\tilde{\mathbf{g}},\mathbf{u}\rangle.$
Let
$X_{n,j}=M_{n,j}-M_{n,j-1}=b_{n}^{1/2}\langle\mathbf{g}_{n+j-1}-\mathbf{g}_{n+j},\mathbf{u}\rangle,$
$X_{n}^{*}=\sup_{j\geq 1}|X_{n,j}|=b_{n}^{1/2}\sup_{j\geq 1}|\langle\mathbf{g}_{n+j-1}-\mathbf{g}_{n+j},\mathbf{u}\rangle|,$
and
$U_{n}=\sum_{j=0}^{\infty}X_{n,j}^{2}=b_{n}\sum_{k\geq n}(\langle\mathbf{g}_{k}-\mathbf{g}_{k-1},\mathbf{u}\rangle)^{2}.$
The thesis follows from Theorem B.2, if we can verify that
(a) the sequence $(X^{*}_{n})$ is dominated in $L^{1}$ and converges to zero $\mathds{P}$-a.s. as $n\rightarrow+\infty$;
(b) $U_{n}\rightarrow\mathbf{u}^{T}\mathbf{V}_{\tilde{g}}\mathbf{u}$, $P$-a.s. as $n\rightarrow+\infty$.
(a) Notice that $\langle\mathbf{g}_{k+1}-\mathbf{g}_{k},\mathbf{u}\rangle=\alpha_{k}\langle\mathbf{g}_{k-1}(\cdot\mid Y_{k})-\mathbf{g}_{k-1},\mathbf{u}\rangle$ is bounded uniformly with respect to $k$. Thus, $X^{*}_{n}\leq b_{n}^{1/2}\sup_{k\geq n}\alpha_{k}|\langle g_{k-1}(\cdot\mid Y_{k})-g_{k-1},\mathbf{u}\rangle|$ is uniformly bounded. Furthermore, it converges to zero $P$-a.s., since $b_{n}\sup_{k\geq n}\alpha_{k}^{2}\rightarrow 0$ by assumption.
(b) The proof is based on Theorem B.4 (b). Let $Z_{n}=\alpha_{n}^{-2}(\langle g_{n}-g_{n-1},\mathbf{u}\rangle)^{2}$ and $a_{k}=(b_{k}\alpha_{k})^{-2}$. We can write that
$E_{n-1}(Z_{n})$ $=E_{n-1}\left(\alpha_{n}^{-2}(\langle g_{n}-g_{n-1},\mathbf{u}\rangle)^{2}\right)$
$=\sum_{z\in\mathbb{N}_{0}}\sum_{i,j=1}^{d-1}u_{i}u_{j}\left(g_{n}(\vartheta_{i}\mid z)g_{n}(\vartheta_{j}\mid z)-g_{n}(\vartheta_{i})g_{n}(\vartheta_{j})\right)p_{g_{n}}(z)=\mathbf{u}^{T}V_{n}\mathbf{u}.$
Thus, by Lemma C.1,
$E_{n-1}\left(Z_{n}\right)\rightarrow\mathbf{u}^{T}V_{\tilde{g}}\mathbf{u},\quad\mathds{P}-a.s.$
Moreover, $b_{n}\uparrow+\infty$ by assumption, and
$\sum_{k=1}^{\infty}\frac{E(Z_{k}^{2})}{a_{k}^{2}b_{k}^{2}}=\sum_{k=1}^{\infty}\alpha_{k}^{4}b_{k}^{2}E\left(\left[\langle\mathbf{g}_{k-1}(\cdot\mid Y_{k})-\mathbf{g}_{k-1},\mathbf{u}\rangle\right]^{4}\right)<+\infty,$
where $(\mathbf{g}(\cdot\mid Y_{k}))_{i}=g(\vartheta_{i}\mid Y_{k})$. Since, by assumption, $b_{n}\sum_{k\geq n}a_{k}^{-1}b_{k}^{-2}\rightarrow 1$, then by Theorem B.4 (b),
$b_{n}\sum_{k\geq n}(\langle g_{k}-g_{k-1},\mathbf{u}\rangle)^{2}=b_{n}\sum_{k\geq n}\frac{Z_{k}}{a_{k}b_{k}^{2}}\rightarrow\mathbf{u}^{T}V_{\tilde{g}}\mathbf{u},\quad\mathds{P}-a.s.$
It follows by Theorem B.2 that $M_{n,\infty}=b_{n}^{1/2}(\langle g_{n}-\tilde{g},\mathbf{u}\rangle)$ converges to $\mathcal{N}(0,\mathbf{u}^{T}V_{\tilde{g}}\mathbf{u})$ in the sense of almost-sure conditional convergence. Since this is true for every $\mathbf{u}$ with $||\mathbf{u}||=1$, then $b_{n}^{1/2}(g_{n}-\tilde{g})$ converges in the sense of almost sure conditional convergence to $\mathcal{N}(0,V_{\tilde{g}})$.
Now suppose that $\alpha_{n}=(\alpha+n)^{-\gamma}$. With $b_{n}=(2\gamma-1)n^{2\gamma-1}$, the conditions (F.20) hold. Indeed, $b_{n}\uparrow+\infty$,
$\sum_{n=1}^{\infty}\alpha_{n}^{4}b_{n}^{2}=\sum_{n=1}^{\infty}\left(\frac{1}{\alpha+n}\right)^{4\gamma}(2\gamma-1)^{2}\alpha^{2\gamma}n^{4\gamma-2}<+\infty,$
$b_{n}\sup_{k\geq n}\alpha_{k}^{2}=(2\gamma-1)n^{2\gamma-1}\left(\frac{1}{\alpha+n}\right)^{2\gamma}\rightarrow 0,$
and
$b_{n}\sum_{k\geq n}{\alpha_{k}^{2}}=(2\gamma-1)n^{2\gamma-1}\sum_{k\geq n}\left(\frac{1}{\alpha+k}\right)^{2\gamma}\rightarrow 1.$
of Theorem 2.2.
The proof is based on Theorem C.2 and Theorem B.5. Let us write
$\hat{\theta}_{g}(y)=h_{y}(\mathbf{g}),$
where
$h_{y}(\mathbf{g})=(y+1)\frac{p_{g}(y+1)}{p_{g}(y)}=(y+1)\frac{\sum_{i=1}^{d-1}k_{y+1}(\vartheta_{i})\mathbf{g}_{i}+k_{y+1}(\vartheta_{d})(1-\sum_{i=1}^{d-1}\mathbf{g}_{i})}{\sum_{i=1}^{d-1}k_{y}(\vartheta_{i})\mathbf{g}_{i}+k_{y}(\vartheta_{i})(1-\sum_{i=1}^{d-1}\mathbf{g}_{i})}$
is seen as a function of the vector $\mathbf{g}\in\Delta^{d}$. The function $h_{y}$ is differentiable with respect to $\mathbf{g}$, and, for $i=1,\dots,d-1$,
$\frac{\partial h_{y}(\mathbf{g})}{\partial\mathbf{g}_{i}}$ $=(y+1)\frac{k_{y+1}(\vartheta_{i})-k_{y+1}(\vartheta_{d})}{\sum_{i=1}^{d-1}k_{y}(\vartheta_{i})g(\vartheta_{i})+k_{y}(\vartheta_{d})(1-\sum_{i=1}^{d-1}g(\vartheta_{i}))}$
$-(y+1)\frac{(\sum_{i=1}^{d-1}k_{y+1}(\vartheta_{i})g(\vartheta_{i})+k_{y+1}(\vartheta_{d})(1-\sum_{i=1}^{d-1}g(\vartheta_{i})))(k_{y}(\vartheta_{i})-k_{y}(\vartheta_{d}))}{(\sum_{i=1}^{d-1}k_{y}(\vartheta_{i})g(\vartheta_{i})+k_{y}(\vartheta_{d})(1-\sum_{i=1}^{d-1}g(\vartheta_{i})))^{2}}$
$=\frac{y+1}{p_{g}(y)}\left[k_{y+1}(\vartheta_{i})-k_{y+1}(\vartheta_{d})-\frac{p_{g}(y+1)}{p_{g}(y)}(k_{y}(\vartheta_{i})-k_{y}(\vartheta_{d}))\right]$
$=\hat{\theta}_{g}(y)\left[\frac{k_{y+1}(\vartheta_{i})-k_{y+1}(\vartheta_{d})}{p_{g}(y+1)}-\frac{k_{y}(\vartheta_{i})-k_{y}(\vartheta_{d})}{p_{g}(y)}\right].$
Let us denote by $\nabla_{g}h_{y}(\mathbf{g})$ the row vector of partial derivatives of $h_{y}$. Since $\Theta_{d}$ is finite, then the partial derivatives are bounded, uniformly with respect to $\mathbf{g}\in\Delta^{d}$. Thus, $h_{y}$ is uniformly differentiable. By Theorem C.2, $b_{n}^{1/2}(\mathbf{g}_{n}-\tilde{\mathbf{g}})$ converges in the sense of almost sure conditional convergence to a centered normal kernel $\mathcal{N}(0,\mathbf{V}_{\tilde{\mathbf{g}}})$, with random covariance matrix $\mathbf{V}_{\tilde{\mathbf{g}}}$ as in (C.2). By Theorem B.5, $b_{n}^{1/2}(h_{y}(\mathbf{g}_{n})-h_{y}(\tilde{\mathbf{g}}))$ converges in the sense of almost sure conditional convergence, to a Gaussian kernel, with covariance matrix
$\nabla_{\tilde{\mathbf{g}}}h_{y}(\tilde{\mathbf{g}})\mathbf{V}_{\tilde{g}}\nabla_{\tilde{g}}h_{\tilde{g}}^{T}$
$=\hat{\theta}_{\tilde{g}}(y)^{2}\sum_{i,j=1}^{d-1}\left[\frac{k(y+1\mid\vartheta_{i})-k(y+1\mid\vartheta_{d})}{p_{\tilde{g}}(y+1)}-\frac{k(y\mid\vartheta_{i})-k(y\mid\vartheta_{d})}{p_{\tilde{g}}(y)}\right]$
$\quad\left[\frac{k(y+1\mid\vartheta_{j})]-k(y+1|\vartheta_{d})}{p_{\tilde{g}}(y+1)}-\frac{k(y\mid\vartheta_{j})-k(y\mid\vartheta_{d})}{p_{\tilde{g}}(y)}\right]$
$\left[\sum_{z\in\mathbb{N}_{0}}\tilde{g}(\vartheta_{i}\mid z)\tilde{g}(\vartheta_{j}\mid z)p_{\tilde{g}}(z)-\tilde{g}(\vartheta_{i})\tilde{g}(\vartheta_{j})\right]$
$=\hat{\theta}_{\tilde{g}}(y)^{2}\sum_{z\in\mathbb{N}_{0}}p_{\tilde{g}}(z)\left[\sum_{i=1}^{d-1}\tilde{g}(\vartheta_{i}\mid z)(\frac{\tilde{g}(\vartheta_{i}\mid y+1)}{\tilde{g}(\vartheta_{i})}-\frac{\tilde{g}(\vartheta_{d}\mid y+1)}{\tilde{g}(\vartheta_{d})}-\frac{\tilde{g}(\vartheta_{i}\mid y)}{\tilde{g}(\vartheta_{i})}+\frac{\tilde{g}(\vartheta_{d}\mid y)}{\tilde{g}(\vartheta_{d})})\right]^{2}$
$-\hat{\theta}_{\tilde{g}}(y)^{2}\left[\sum_{i=1}^{d-1}\tilde{g}(\vartheta_{i})(\frac{\tilde{g}(\vartheta_{i}\mid y+1)}{\tilde{g}(\vartheta_{i})}-\frac{\tilde{g}(\vartheta_{d}\mid y+1)}{\tilde{g}(\vartheta_{d})}-\frac{\tilde{g}(\vartheta_{i}\mid y)}{\tilde{g}(\vartheta_{i})}+\frac{\tilde{g}(\vartheta_{d}\mid y)}{\tilde{g}(\vartheta_{d})})\right]^{2}$
$=\hat{\theta}_{\tilde{g}}(y)^{2}\sum_{z\in\mathbb{N}_{0}}p_{\tilde{g}}(z)\left[\sum_{i=1}^{d-1}\tilde{g}(\vartheta_{i}\mid z)(\frac{\tilde{g}(\vartheta_{i}\mid y+1)-\tilde{g}(\vartheta_{i}\mid y)}{\tilde{g}(\vartheta_{i})})-(1-\tilde{g}(\vartheta_{d}\mid z))\frac{\tilde{g}(\vartheta_{d}\mid y+1)-\tilde{g}(\vartheta_{d}\mid y)}{\tilde{g}(\vartheta_{d})}\right]^{2}$
$-\hat{\theta}_{\tilde{g}}(y)^{2}\left[\sum_{i=1}^{d-1}\tilde{g}(\vartheta_{i})\frac{\tilde{g}(\vartheta_{i}\mid y+1)-\tilde{g}(\vartheta_{i}\mid y)}{\tilde{g}(\vartheta_{i})}-(1-\tilde{g}(\vartheta_{d}))\frac{\tilde{g}(\vartheta_{d}\mid y+1)-\tilde{g}(\vartheta_{d}\mid y)}{\tilde{g}(\vartheta_{d})}\right]^{2}$
$=\hat{\theta}_{\tilde{g}}(y)^{2}\sum_{z\in\mathbb{N}_{0}}p_{\tilde{g}}(z)\left[\sum_{i=1}^{d}\frac{\tilde{g}(\vartheta_{i}\mid z)}{\tilde{g}(\vartheta_{i})}(\tilde{g}(\vartheta_{i}\mid y+1)-\tilde{g}(\vartheta_{i}\mid y))\right]^{2}$
$-\hat{\theta}_{\tilde{g}}(y)^{2}\left[\sum_{i=1}^{d}\tilde{g}(\vartheta_{i}\mid y+1)-\tilde{g}(\vartheta_{i}\mid y)\right]^{2}$
$=\hat{\theta}_{\tilde{g}}(y)^{2}\sum_{z\in\mathbb{N}_{0}}p_{\tilde{g}}(z)\left[\sum_{i=1}^{d}\frac{\tilde{g}(\vartheta_{i}\mid z)}{\tilde{g}(\vartheta_{i})}(\tilde{g}(\vartheta_{i}\mid y+1)-\tilde{g}(\vartheta_{i}\mid y))\right]^{2}.$
C.3 Proof of Theorem 2.4
The proof of Theorem 2.4 is based on the following lemma C.3 and on the rate of convergence of $g_{n}$ to $g$, under i.i.d. observations. This last result is known in the literature (Martin and Tokdar, 2009), but we think that the proof is not complete. We provide for completeness the proof in Lemma C.5.
Lemma C.3.
Under the assumptions of Theorem 2.4, there exists a constant $C=C(d,\vartheta_{1},\vartheta_{d})$ such that for every $n\geq 1$
${\rm Regret}(g_{n},g^{*})\leq C||\mathbf{g}_{n}-\mathbf{g}^{*}||^{2},$
where $\mathbf{g}_{n}$ and $\mathbf{g}^{*}$ are the representations of $g_{n}$ and $g^{*}$ in $\mathbb{R}^{d}$.
Proof.
We can write that
$\sum_{y\in\mathbb{N}_{0}}(\hat{\theta}_{g_{n}}(y)-\hat{\theta}_{g^{*}}(y))^{2}p_{g^{*}}(y)$
$=\sum_{y\in\mathbb{N}_{0}}\left(\frac{\sum_{i=1}^{d}\vartheta_{i}^{y+1}e^{-\vartheta_{i}}g_{n}(\vartheta_{i})}{\sum_{i=1}^{d}\vartheta_{i}^{y}e^{-\vartheta_{i}}g_{n}(\vartheta_{i})}-\frac{\sum_{i=1}^{d}\vartheta_{i}^{y+1}e^{-\vartheta_{i}}g^{*}(\vartheta_{i})}{\sum_{i=1}^{d}\vartheta_{i}^{y}e^{-\vartheta_{i}}g^{*}(\vartheta_{i})}\right)^{2}\frac{1}{y!}\sum_{i=1}^{d}\vartheta_{i}^{y}e^{-\vartheta_{i}}g^{*}(\vartheta_{i})$
$\leq 2\sum_{y\in\mathbb{N}_{0}}\frac{1}{y!}\frac{\left(\sum_{i=1}^{d}\vartheta_{i}^{y+1}e^{-\vartheta_{i}}\left(g_{n}(\vartheta_{i})-g^{*}(\vartheta_{i})\right)\right)^{2}}{\sum_{i=1}^{d}\vartheta_{i}^{y}e^{-\vartheta_{i}}g^{*}(\vartheta_{i})}$
$+2\sum_{y\in\mathbb{N}_{0}}\frac{1}{y!}\frac{\left(\sum_{i=1}^{d}\vartheta_{i}^{y+1}e^{-\vartheta_{i}}g_{n}(\vartheta_{i})\right)^{2}}{\left(\sum_{i=1}^{d}\vartheta_{i}^{y}e^{-\vartheta_{i}}g_{n}(\vartheta_{i})\right)^{2}\sum_{i=1}^{d}\vartheta_{i}^{y}e^{-\vartheta_{i}}g^{*}(\vartheta_{i})}\left(\sum_{i=1}^{d}\vartheta_{i}^{y}e^{-\vartheta_{i}}(g_{n}(\vartheta_{i})-g^{*}(\vartheta_{i})\right)^{2}$
$\leq 2\sum_{y\in\mathbb{N}_{0}}\frac{1}{y!}\frac{\sum_{i=1}^{d}\vartheta_{i}^{2y+2}e^{-2\vartheta_{i}}||g_{n}-g^{*}||^{2}}{\sum_{i=1}^{d}\vartheta_{i}^{y}e^{-\vartheta_{i}}g^{*}(\vartheta_{i})}+2\sum_{y\in\mathbb{N}_{0}}\frac{1}{y!}\frac{\vartheta_{d}^{2}\sum_{i=1}^{d}\vartheta_{i}^{2y}e^{-2\vartheta_{i}}||g_{n}-g^{*}||^{2}}{\sum_{i=1}^{d}\vartheta_{i}^{y}e^{-\vartheta_{i}}g^{*}(\vartheta_{i})}$
$\leq 4\vartheta_{d}^{2}||\mathbf{g}_{n}-\mathbf{g}^{*}||^{2}\sum_{y\in\mathbb{N}_{0}}\frac{1}{y!}\frac{\sum_{i=1}^{d}\vartheta_{i}^{2y}e^{-2\vartheta_{i}}}{\sum_{i=1}^{d}\vartheta_{i}^{y}e^{-\vartheta_{i}}g^{*}(\vartheta_{i})}$
$\leq C||\mathbf{g}_{n}-\mathbf{g}^{*}||^{2},$
with $C=4d\vartheta_{d}^{2}e^{\vartheta_{d}^{2}/\vartheta_{1}-2\vartheta_{1}+\vartheta_{d}}$. Thus,
${\rm Regret}(g_{n},g^{*})\leq C||\mathbf{g}_{n}-\mathbf{g}^{*}||^{2}.$
Lemma C.4.
Under the assumptions of Theorem 2.4, as $n\rightarrow\infty$ the stochastic approximation in $\Delta^{d}$:
$\mathbf{g}_{n+1}=\mathbf{g}_{n}+\alpha_{n+1}\mathbf{h}(\mathbf{g}_{n})+\alpha_{n+1}{\bm{\epsilon}}_{n+1},$ (C.6)
with $\alpha_{n}>0$, $\sum_{n=1}^{\infty}\alpha_{n}=\infty$, $\sum_{n=1}^{\infty}\alpha_{n}^{2}<\infty$,
$\mathbf{h}(\mathbf{g})=\mathbf{g}\circ\left(\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}}{p_{\mathbf{g}}(y)}p_{\mathbf{g}^{*}}(y)-\bf 1\right),$
${\bm{\epsilon}}_{n+1}=\mathbf{g}_{n}\circ\left(\frac{\mathbf{k}_{Y_{n+1}}}{p_{\mathbf{g}_{n}}(Y_{n+1})}-\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}}{p_{\mathbf{g}_{n}}(y)}p_{\mathbf{g}^{*}}(y)\right),$
and initial state $\mathbf{g}_{0}\in\mathring{\Delta}^{d}$, converges to $\mathbf{g}^{*}$, $\mathds{P}^{*}$-a.s.
Proof.
Let $J=\partial\Delta^{d}\cup\{g^{*}\}$. We first prove that ${\rm d}(\mathbf{g}_{n},J)\rightarrow 0$, $\mathds{P}^{*}$-a.s., as $n\rightarrow\infty$, where ${\rm d}$ denotes the Euclidean distance in $\mathbb{R}^{d-1}$. The result follows from Theorem 2.2.1′′ in Chen (2005), if we can prove that the following conditions hold:
  • A1
    $\alpha_{n}>0$, $\alpha_{n}\rightarrow 0$, as $n\rightarrow\infty$ and $\sum_{n=1}^{\infty}\alpha_{n}=\infty$;
  • A2
    $J$ is closed, and there exists a is continuously differentiable function $\mathscr{l}:\Delta^{d}\rightarrow\mathbb{R}$ such that $\sup_{\delta_{1}\leq{\rm d}(\mathbf{g},J)\leq\delta_{2}}\mathbf{h}(\mathbf{g})^{T}\nabla_{\mathbf{g}}\mathscr{l}<0$ for every $\delta_{2}>\delta_{1}>0$:
  • A3
    $\lim_{T\rightarrow\infty}\limsup_{n\rightarrow\infty}\frac{1}{T}||\sum_{i=n}^{m(n,T)}\alpha_{i+1}\epsilon_{i+1}||=0,$ $\mathds{P}^{*}$-a.s., where $m(n,T)=\max\{m:\sum_{i=n}^{m}\alpha_{i+1}\leq T\}$;
  • A4
    $\mathbf{h}$ is measurable and locally bounded.
Let $\mathscr{l}$ be defined as
$\mathscr{l}(\mathbf{g})=\sum_{y\in\mathbb{N}_{0}}\log\left(\frac{p_{\mathbf{g}^{*}}(y)}{p_{\mathbf{g}}(y)}\right)p_{\mathbf{g}^{*}}(y).$
A1 holds by assumption; A3 holds since $({\bm{\epsilon}}_{n})$ is a martingale difference, and A4 is obvious.
We now prove A2. The set $J$ is closed and $\mathscr{l}$ is continuously differentiable. Moreover, for $\mathbf{g}\in\mathring{(\Delta^{d})}$,
$\mathbf{h}(\mathbf{g})^{T}\nabla_{\mathbf{g}}\mathscr{l}$ $=-\left(\mathbf{g}\circ\left(\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}}{p_{\mathbf{g}}(y)}p_{\mathbf{g}^{*}}(y)-\bf 1\right)\right)^{T}\left(\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}-\mathbf{k}_{y}(\vartheta_{d})\mathbf{1}}{p_{\mathbf{g}}(y)}p_{\mathbf{g}^{*}}(y)\right)$
$=-\left(\sum_{i=1}^{d}\mathbf{g}_{i}\left(\sum_{y\in\mathbb{N}_{0}}k_{y}(\vartheta_{i})\frac{p_{\mathbf{g}^{*}}(y)}{p_{\mathbf{g}}(y)}-1\right)^{2}\right)$
$\leq-\left(\sum_{i=1}^{d}\mathbf{g}_{i}\left(\sum_{y\in\mathbb{N}_{0}}k_{y}(\vartheta_{i})\frac{p_{\mathbf{g}^{*}}(y)}{p_{\mathbf{g}}(y)}-1\right)\right)^{2}=0,$
where, in the inequality, we have used Jensen inequality. We now prove that the inequality is strict, unless $\mathbf{g}=\mathbf{g}^{*}$. By contradiction, assume that there exist $\mathbf{g}^{\prime}$ and $\mathbf{g}^{\prime\prime}$ with $\mathbf{g}^{\prime}\neq\mathbf{g}^{\prime\prime}$ such that $\sum_{y\in\mathbb{N}_{0}}k_{y}(\vartheta_{i})\frac{p_{\mathbf{g}^{*}}(y)}{p_{\mathbf{g}}(y)}=1$ for every $i=1,\dots,d$ both for $\mathbf{g}=\mathbf{g}^{\prime}$ and for $\mathbf{g}=\mathbf{g}^{\prime\prime}$. Then, for every $a\in(0,1)$
$1=\sum_{i=1}^{d}(a\mathbf{g}^{\prime}_{i})+(1-a)\mathbf{g}^{\prime\prime}_{i})\sum_{y\in\mathbb{N}_{0}}k_{y}(\vartheta_{i})\frac{p_{g^{*}}(y)}{ap_{\mathbf{g}^{\prime}}(y)+(1-a)p_{\mathbf{g}^{\prime\prime}}(y)}$
$<\sum_{i=1}^{d}(a\mathbf{g}^{\prime}_{i}+(1-a)\mathbf{g}^{\prime\prime}_{i})\left(a\sum_{y\in\mathbb{N}_{0}}k_{y}(\vartheta_{i})\frac{p_{\mathbf{g}^{*}}(y)}{p_{\mathbf{g}^{\prime}}(y)}+(1-a)\sum_{y\in\mathbb{N}_{0}}k_{y}(\vartheta_{i})\frac{p_{\mathbf{g}^{*}}(y)}{p_{\mathbf{g}^{\prime\prime}}(y)}\right)$
$=1,$
which leads to a contradiction. Hence A2 holds, and ${\rm d}(\mathbf{g}_{n},J)\rightarrow 0$ $\mathds{P}^{*}$-a.s., as $n\rightarrow\infty$.
To conclude, we need to show that $\mathds{P}^{*}(\liminf_{n}{\rm d}(\mathbf{g}_{n},\partial\Delta^{d})=0)=0$. Since ${\rm d}(\mathbf{g}^{*},\partial\Delta^{d})>0$ and ${\rm d}(\mathbf{g}_{n},\{\mathbf{g}^{*}\}\cup\partial\Delta^{d})\rightarrow 0$ $\mathds{P}-a.s.$, then $\mathds{P}^{*}(\liminf_{n}{\rm d}(\mathbf{g}_{n},\partial\Delta^{d})=0)>0$ if and only if there exists $M$ with $\mathds{P}^{*}(M)>0$ such that ${\rm d}(\mathbf{g}_{n}(\omega),\partial\Delta^{d})\rightarrow 0$ for every $\omega\in M$, as $n\rightarrow\infty$. Let us make this assumption as a contradiction hypothesis. Define the function $\mathcal{L}(\mathbf{g}_{n})=\sum_{i=1}^{d}\log(g^{*}(\vartheta_{i})/g_{n}(\vartheta_{i}))g^{*}(\vartheta_{i})$. By the contradiction hypothesis, $\mathcal{L}(\mathbf{g}_{n})(\omega)\rightarrow\infty$ as $n\rightarrow\infty$, for every $\omega\in M$. Then $E^{*}({\mathcal{L}}(\mathbf{g}_{n}))\rightarrow\infty$ as $n\rightarrow+\infty$, where $E^{*}$ denote expectation with respect to $\mathds{P}^{*}$. Let $\overline{\mathbf{g}}_{n}=\eta_{n}\mathbf{g}_{n}+(1-\eta_{n})\mathbf{g}_{n+1}$, with the random variable $\eta_{n}$ taking values in $[0,1]$, be such that
${\mathcal{L}}(\mathbf{g}_{n+1})$ $=\mathcal{L}(\mathbf{g}_{n})+\nabla_{\mathbf{g}}\mathcal{L}(\mathbf{g}_{n})^{T}(\mathbf{g}_{n+1}-\mathbf{g}_{n})$
$+(\mathbf{g}_{n+1}-\mathbf{g}_{n})^{T}\nabla^{2}_{\mathbf{g}}\mathcal{L}(\overline{\mathbf{g}}_{n})(\mathbf{g}_{n+1}-\mathbf{g}_{n})$
$=\mathcal{L}(\mathbf{g}_{n})-\sum_{i=1}^{d}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})+\sum_{i=1}^{d}\frac{\mathbf{g}^{*}_{i}}{(\overline{\mathbf{g}}_{n,i})^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}.$
Notice that
$E^{*}\left(\sum_{i=1}^{d}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i}\right)=\alpha_{n+1}E^{*}\left(\sum_{i=1}^{d}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}(\mathbf{h}_{i}(\mathbf{g}_{n})+{\bm{\epsilon}}_{n+1,i})\right),$
with
$E^{*}\left(\sum_{i=1}^{d}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}{\bm{\epsilon}}_{n+1,i}\right)=0,$
and
$\liminf_{n}E^{*}\left(\sum_{i=1}^{d}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}\mathbf{h}_{i}(\mathbf{g}_{n})\right)=\liminf_{n}E^{*}\left(\sum_{i=1}^{d}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}\left(\sum_{y\in\mathbb{N}_{0}}\frac{k_{y}(\vartheta_{i})\mathbf{g}_{n,i}p_{\mathbf{g}^{*}}(y)}{p_{\mathbf{g}_{n}}(y)}-\mathbf{g}_{n,i}\right)\right)$
$=\liminf_{n}E^{*}\left(\sum_{y\in\mathbb{N}_{0}}\frac{p_{\mathbf{g}^{*}}(y)^{2}}{p_{\mathbf{g}_{n}}(y)}-1\right)$
$>\liminf_{n}E^{*}\left(\left(\sum_{y\in\mathbb{N}_{0}}\frac{p_{\mathbf{g}_{n}}(y)}{p_{\mathbf{g}^{*}}(y)}p_{\mathbf{g}^{*}_{i}}(y)\right)^{-1}-1\right)=0,$
where the strict inequality holds since, under the contradiction assumption, $\mathds{P}(p_{\mathbf{g}_{n}}\not\rightarrow p_{\mathbf{g}^{*}})>0$. Thus there exists a constant $c>0$ such that
$E^{*}\left(-\sum_{i=1}^{d}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})\right)<-c\alpha_{n+1}$
for $n$ large enough. On the other hand, denoting by $E^{*}_{n}$ the expectation under $\mathds{P}^{*}$, conditional to $\mathcal{G}_{n}$, we can write that
$E^{*}\left(\sum_{i=1}^{d}\frac{\mathbf{g}^{*}_{i}}{(\overline{\mathbf{g}}_{n,i})^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}\right)$
$\leq E^{*}\left(\sum_{i=1}^{d}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}\right)+E^{*}\left(\sum_{i=1}^{d}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n+1,i}^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}\right)$
$\leq\alpha_{n+1}^{2}E^{*}\left(\sum_{i=1}^{d}\mathbf{g}^{*}_{i}\left(\frac{k_{Y_{n+1}}(\vartheta_{i})}{p_{\mathbf{g}_{n}}(Y_{n+1})}-1\right)^{2}\right)+\alpha_{n+1}^{2}E^{*}\left(\sum_{i=1}^{d}\mathbf{g}^{*}_{i}\frac{\left(\frac{k_{Y_{n+1}}(\vartheta_{i})}{p_{\mathbf{g}_{n}}(Y_{n+1})}-1\right)^{2}}{\left(1+\alpha_{n+1}\left(\frac{k_{Y_{n+1}}(\vartheta_{i})}{p_{\mathbf{g}_{n}}(Y_{n+1})}-1\right)\right)^{2}}\right)$
$\leq\alpha_{n+1}^{2}E^{*}\left(1+E^{*}_{n}\left(\sum_{i=1}^{d}\left(\frac{k_{Y_{n+1}}(\vartheta_{i})}{\sum_{i=1}^{d}k_{Y_{n+1}}(\vartheta_{i})\mathbf{g}_{n,i}}\right)^{2}\right)\right)$
$\quad\quad+\frac{\alpha_{n+1}^{2}}{(1-\alpha_{n+1})^{2}}E^{*}\left(1+E^{*}_{n}\left(\sum_{i=1}^{d}\left(\frac{k_{Y_{n+1}}(\vartheta_{i})}{\sum_{i=1}^{d}k_{Y_{n+1}}(\vartheta_{i})\mathbf{g}_{n,i}}\right)^{2}\right)\right)$
$\leq 2\frac{\alpha_{n+1}^{2}}{(1-\alpha_{n+1})^{2}}\left(1+\sum_{y\in\mathbb{N}_{0}}\sum_{i,j,k}\frac{k_{y}(\vartheta_{i})^{2}}{(\sum_{\vartheta_{k}}k_{y}(\vartheta_{k})g_{n}(\vartheta_{k}))^{2}}k_{y}(\vartheta_{j})g^{*}(\vartheta_{j})\right)$
$\leq 2\frac{\alpha_{n+1}^{2}}{(1-\alpha_{n+1})^{2}}\left(1+\sum_{y\in\mathbb{N}_{0}}\sum_{i,j,k}\frac{k_{y}(\vartheta_{i})^{2}}{k_{y}(\vartheta_{k})^{2}}k_{y}(\vartheta_{j})g^{*}(\vartheta_{j})g_{n}(\vartheta_{k})\right)$
$\leq 2\frac{\alpha_{n+1}^{2}}{(1-\alpha_{n+1})^{2}}\left(1+d\sum_{y\in\mathbb{N}_{0}}\ max_{i,j,k}\frac{k_{y}(\vartheta_{i})^{2}}{k_{y}(\vartheta_{k})^{2}}k_{y}(\vartheta_{j})\right)$
$=o(\alpha_{n}),$
where the last equality holds since
$\sum_{y\in\mathbb{N}_{0}}\frac{k_{y}(\vartheta_{i})^{2}}{k_{y}(\vartheta_{j})^{2}}k_{y}(\vartheta_{k})=\sum_{y\in\mathbb{N}_{0}}e^{2\vartheta_{j}-2\vartheta_{i}-\vartheta_{k}}\left(\frac{\vartheta_{i}^{2}\vartheta_{k}}{\vartheta_{j}^{2}}\right)^{y}\frac{1}{y!}<\infty.$
It follows that,
$E^{*}\left(\mathcal{L}(\mathbf{g}_{n+1})\right)\leq E^{*}\left(\mathcal{L}(\mathbf{g}_{n})\right),$
for $n$ large enough, which contradicts $E(\mathcal{L}(\mathbf{g}_{n}))\rightarrow\infty$. ∎
Lemma C.5.
Let $\mathbf{g}_{n}$ and $\mathbf{g}^{*}$ be the representations of $g_{n}$ ($n\geq 0$) and $g^{*}$ in $\mathbb{R}^{d}$. Under the assumptions of Theorem 2.4, for every $\delta<2-1/\gamma$
$||\mathbf{g}_{n}-\mathbf{g}^{*}||^{2}=o(n^{-\delta})\quad\mathds{P}^{*}-a.s.$
Proof.
In this proof, $\mathbf{g}_{n}$ and $\mathbf{g}^{*}$ are considered as random vectors in $\{\mathbf{v}\in\mathbb{R}^{d}:||\mathbf{v}||\leq 1\}$. By Lemma F.4, $\mathbf{g}_{n}$, seen as vectors in $\Delta^{d}$ converge $P^{*}$-a.s. to $\mathbf{g}\in\Delta^{d}$. This implies that $\mathbf{g}_{n}$ converges to $\mathbf{g}^{*}$ also when seen as vectors in $\mathbb{R}^{d}$. Moreover, $(\mathbf{g}_{n})$ satisfies the stochastic approximation in $\mathbb{R}^{d}$:
$\mathbf{g}_{n+1}$ $=\mathbf{g}_{n}+\alpha_{n+1}\mathbf{g}_{n}\circ\left(\frac{\mathbf{k}_{Y_{n+1}}}{p_{\mathbf{g}_{n}}(Y_{n+1})}-\bf 1\right)$
$=\mathbf{g}_{n}+\alpha_{n+1}\mathbf{h}(\mathbf{g}_{n})+\alpha_{n+1}{\bm{\epsilon}}_{n+1},$
with initial value $\mathbf{g}_{0}$, where
$\mathbf{h}(\mathbf{g})=\mathbf{g}\circ\left(\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}}{p_{\mathbf{g}}(y)}p_{\mathbf{g}^{*}}(y)-\bf 1\right),$
and
${\bm{\epsilon}}_{n+1}=\mathbf{g}\circ\left(\frac{\mathbf{k}_{Y_{n+1}}}{p_{\mathbf{g}}(Y_{n+1})}p_{\mathbf{g}^{*}}(Y_{n+1})-\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}}{p_{\mathbf{g}}(y)}p_{\mathbf{g}^{*}}(y)\right).$
Based on Lemma C.4, $\mathbf{g}_{n}$ converges to $\mathbf{g}^{*}$, $\mathds{P}^{*}$-a.s. The thesis follows from Theorem 3.1.1 in Chen (2005) if we can prove that the following conditions hold:
  • B1
    $\alpha_{n}\rightarrow 0$, $\sum_{n=1}^{\infty}\alpha_{n}=\infty$, $\alpha_{n+1}^{-1}-\alpha_{n}^{-1}\rightarrow 0$;
  • B2
    $\sum_{n=1}^{\infty}\alpha_{n}^{1-\delta/2}\epsilon_{n}$ converges, $\mathds{P}^{*}$-a.s.;
  • B3
    $\mathbf{h}$ is measurable and locally bounded, and is differentiable at $\mathbf{g}^{*}$. Let $\mathbf{H}$ be a matrix such that, as $\mathbf{g}\rightarrow\mathbf{g}^{*}$,
    $\mathbf{h}(\mathbf{g})=\mathbf{H}(\mathbf{g}-\mathbf{g}^{*})+\mathbf{r}(g),\quad\mathbf{r}(g^{*})=\mathbf{0},\quad||\mathbf{r}(\mathbf{g})||=o(||\mathbf{g}-\mathbf{g}^{*}||).$
    All the eigenvalues of the matrix $\mathbf{H}$ have negative real parts.
Remark C.6.
Theorem 3.1.1 in Chen (2005) also requests the following condition: B4 There exists a continuously differentiable function $\mathscr{l}:\Delta^{d}\rightarrow\mathbb{R}$ such that $\sup_{\delta_{1}\leq{\rm d}(\mathbf{g},\mathbf{g}^{*})\leq\delta_{2}}\mathbf{h}(\mathbf{g})^{T}\nabla_{\mathbf{g}}\mathscr{l}<0$ for every $\delta_{2}>\delta_{1}>0$. However, a close inspection of the proof of Theorem 3.1.1 in Chen shows that B4 is employed only to ensure convergence of $\mathbf{g}_{n}$ towards $\mathbf{g}^{*}$. In our case, this property has been proved in Lemma C.4.
B1 is obvious. To prove B2, notice that $(\sum_{k=1}^{n}\alpha_{k}^{1-\delta/2}\epsilon_{k})_{n\geq 1}$ is a martingale. Moreover, there exists a constant $C$ such that, for every $i=1,\dots,d$,
$\sup_{n}E^{*}((\sum_{k=1}^{n}\alpha_{k}^{1-\delta/2}{\bm{\epsilon}}_{k,i})^{2})$ $=\sup_{n}\sum_{k=1}^{n}\alpha_{k}^{2-\delta}E^{*}({\bm{\epsilon}}_{k,i})^{2})$
$\leq C\sum_{k=1}^{\infty}\left(\frac{1}{\alpha+k}\right)^{\gamma(2-\delta)}<\infty,$
since $\gamma(2-\delta)>1$ and $({\bm{\epsilon}}_{n,i})$ are uniformly bounded. We now prove B3. The function $\mathbf{h}$ is continuously differentiable at $\mathbf{g}^{*}$ and
$\mathbf{H}_{i,j}=-\mathbf{g}^{*}_{i}\sum_{y\in\mathbb{N}_{0}}\frac{k_{y}(\vartheta_{i})k_{y}(\vartheta_{j})}{p_{g^{*}}(y)}.$
The matrix $\mathbf{H}$ can be seen as the product of the diagonal matrix ${\rm diag}(\mathbf{g}^{*})$, which is positive definite, times $-\mathbf{M}$, with $\mathbf{M}_{i,j}=\sum_{y\in\mathbb{N}_{0}}\frac{k_{y}(\vartheta_{i})k_{y}(\vartheta_{j})}{p_{g^{*}}(y)}$. We now prove that $\mathbf{M}$ is positive definite. Given a vector $\mathbf{u}\in\mathbb{R}^{d}$ with $||\mathbf{u}||=1$, we can write that
$\mathbf{u}^{T}\mathbf{M}\mathbf{u}=$ $\sum_{i,j=1}^{d}\mathbf{u}_{i}\mathbf{u}_{j}\sum_{y\in\mathbb{N}_{0}}\frac{k_{y}(\vartheta_{i})k_{y}(\vartheta_{j})}{p_{g^{*}}(y)}=\sum_{y\in\mathbb{N}_{0}}\frac{1}{p_{g^{*}}(y)}\left(\sum_{i=1}^{d}\mathbf{u}_{i}k_{y}(\vartheta_{i})\right)^{2}\geq 0,$
with equality if and only if $\sum_{i=1}^{d}\mathbf{u}_{i}k_{y}(\vartheta_{i})=0$ for every $y\in\mathbb{N}_{0}$. The only solution to the last equation in $\mathbf{u}$ is the zero vector, which is inconsistent with $||\mathbf{u}||=1$. Since the function $\mathbf{u}^{T}\mathbf{M}\mathbf{u}$ is continuous on the compact set $||\mathbf{u}||=1$, then the minimum is attained, and by the above reasoning it is strictly positive. It follows that $\mathbf{M}$ is positive definite. ∎
of Theorem 2.4.
Denoting by $(\mathbf{g}_{n})$ and $\mathbf{g}^{*}$ the representations of $(g_{n})$ and $g^{*}$ in $\mathbb{R}^{d}$, we can write, by Lemma C.3 and Lemma C.5,
${\rm Regret}(g_{n},g^{*})\leq C||\mathbf{g}_{n}-\mathbf{g}^{*}||^{2}=o(n^{-\delta}).$
D Numerical illustrations
D.1 Synthetic data
We set $n=50,100,200,500$, and generate $(Y_{1},\ldots,Y_{n})$ under the Poisson compound model, with $(\theta_{1},\ldots,\theta_{n})$ i.i.d from a Uniform prior $G$ on $[0,3]$. For the NP-QB method, we set Newton’s algorithm as follows: i) the parameter $\theta$ takes value in $\Theta=\{0.01,0.02,\ldots,99.99,100\}$; ii) the number $d$ of mixture components is the cardinality of $\Theta$, i.e. $d=10,000$; iii) the initial guess $g_{0}$ is the Uniform distribution on $\Theta$; iv) the learning rate is $\alpha_{n}=(1+n)^{-1}$. Results are displayed in Figure 2 and Table 3.
[…]
Figure 2: Data $Y_{i}$’s against the $\theta_{i}$’s (black circle), and estimates of the $\theta_{i}$’s: NP-EB (green line), NP-ML (cyan line), NP-MD (magenta line), P-EB (red line) and NP-QB (blue line).
Table 3: Values of RMSE and MAD for the NP-EB, NP-ML, NP-MD, P-EB and NP-QB estimates.
NP-EB NP-ML NP-MD P-EB NP-QB
RMSE   $n=100$ 1.01 0.75 1.31 0.75 0.75
RMSE   $n=500$ 0.95 0.75 0.74 0.75 0.78
MAD     $n=100$ 0.77 0.60 1.05 0.61 0.60
MAD     $n=500$ 0.65 0.59 0.59 0.60 0.61
Figure 3: Data $Y_{i}$’s against the $\theta_{i}$’s (black circle), and estimates of the $\theta_{i}$’s: NP-EB (green line), NP-ML (cyan line), NP-MD (magenta line), P-EB (red line) and NP-QB (blue line).
Table 4: Values of RMSE and MAD for the NP-EB, NP-ML, NP-MD, P-EB and NP-QB estimates.
NP-EB NP-ML NP-MD P-EB NP-QB
RMSE   $n=100$ 5.83 1.99 1.71 1.83 1.75
RMSE   $n=500$ 2.63 1.89 1.86 1.93 1.86
MAD     $n=100$ 3.40 1.52 1.32 1.38 1.34
MAD     $n=500$ 1.75 1.42 1.42 1.47 1.43
D.2 Real data: Tweeter dataset
We consider a dataset of tweets with at least 50 retweets on Twitter in the period October 7 to November 6, 2011 (30 days). The dataset is available at the link https://snap.stanford.edu/seismic/, and it consists of $166,783$ tweets that generate $34,617,704$ retweets. More specifically, the data provides information on the tweet ID, the tweet post time, the retweet post time and the number of followers of the poster/retweeter.
For each tweet, we focus on the number of retweets generated within $30$ seconds from the tweet post time. From the first two rows of Table 5, $40,259$ tweets have 00 retweets in $30$ seconds, $28,339$ tweets have $1$ retweet in $30$ seconds, etc. Table 5 and Figure 4 display the estimates of the number of retweets expected for a tweet.
For each tweet, we focus on the number of followers generated by retweets within $30$ seconds from the tweet post time, namely the total number of followers of fast retweets of a tweet, with the proviso that 00 retweets corresponds to 00 followers generated. From the first two rows of Table 6, $40,287$ tweets have generated 00 followers in $30$ seconds, $67$ tweets have generated $1$ follower in $30$ seconds, etc. Table 6 and Figure 5 display the estimates of the number of followers expected to be generated by the retweets of a tweet.
Table 5: Counts $n_{y}$ of the number of tweets that have $y$ retweets within $30$ seconds from the tweet post time, for $y=0,1,\ldots,77$, and corresponding estimates of the number of retweets expected.
$y$ 0 1 2 3 4 5 6 7
$n_{y}$ 40,259 28,339 21,581 16,479 12,130 9,238 7,193 5,464
NP-EB 0.70 1.52 2.29 2.94 3.81 4.67 5.32 6.32
NP-ML 0.39 2.13 2.93 3.70 4.46 5.33 6.33 7.42
NP-MD 0.75 2.20 3.14 4.15 4.99 5.69 6.39 7.16
P-EB 0.45 1.34 2.23 3.12 4.01 4.90 5.79 6.68
NP-QB 0.61 0.82 1.63 3.08 4.01 4.52 5.02 5.71
$y$ 8 11 14 17 20 23 26 29
$n_{y}$ 4,319 2,291 1,352 817 475 293 188 129
NP-EB 7.18 10.10 11.77 15.33 18.66 21.13 22.40 28.37
NP-ML 8.54 11.79 15.01 18.30 20.89 22.96 25.87 30.47
NP-MD 8.04 10.87 13.02 14.83 16.68 18.82 22.50 29.21
P-EB 7.57 10.24 12.91 15.58 18.25 20.93 23.60 26.27
NP-QB 6.67 10.37 12.69 14.64 17.55 21.38 24.78 27.53
$y$ 32 35 38 41 44 47 50 53
$n_{y}$ 89 69 60 61 36 36 32 21
NP-EB 30.78 32.35 30.55 26.85 48.75 45.33 43.03 41.14
NP-ML 34.88 37.75 39.81 41.73 43.79 46.18 49.13 52.72
NP-MD 35.48 39.08 41.63 44.05 46.58 49.21 51.93 54.70
P-EB 28.94 31.61 34.28 36.96 39.63 42.30 44.97 47.64
NP-QB 30.26 33.32 36.59 39.71 42.50 45.05 47.54 50.09
$y$ 56 59 62 65 68 71 74 77
$n_{y}$ 24 16 14 13 14 11 2 5
NP-EB 42.75 48.75 45.00 91.39 14.79 19.64 225.00 62.40
NP-ML 56.58 60.08 62.88 65.06 66.84 68.38 69.77 71.01
NP-MD 57.47 60.21 62.84 65.31 67.54 69.49 71.15 72.53
P-EB 50.31 52.99 55.66 58.33 61.00 63.67 66.35 69.02
NP-QB 52.74 55.40 57.94 60.26 62.33 64.18 65.87 67.45
Figure 4: Estimates of the number of retweets expected for a tweet with $y$ retweets (in 30 seconds): NP-EB (green line), NP-ML (cyan line), NP-MD (magenta line), P-EB (red line) and NP-QB (blue line).
Table 6: Counts $n_{y}$ of the number of tweets that have generated $y$ followers within $30$ seconds from the tweet post time, for $y=0,1,\ldots,450$, and corresponding estimates of the number of followers expected.
$y$ 0 30 60 90 120 150 180 210
$n_{y}$ 40,287 67 87 83 98 82 84 63
NP-EB 0 31.46 55.39 104.16 101.25 147.32 178.85 247.84
NP-ML 0 31.67 61.10 91.01 120.99 149.46 180.19 210.79
NP-MD 0.78 29.93 60.62 93.26 121.19 148.30 181.52 213.79
P-EB 0.56 30.410 60.26 90.12 119.97 149.82 179.67 209.52
NP-QB 0.54 31.41 60.90 91.19 120.68 150.89 180.31 210.85
$y$ 240 270 300 330 360 390 420 450
$n_{y}$ 85 76 65 81 81 63 75 71
NP-EB 275.02 274.57 342.68 314.65 352.09 378.59 370.48 330.31
NP-ML 240.81 270.90 300.96 330.97 361.00 391.05 421.04 450.46
NP-MD 244.11 270.94 300.95 330.98 360.98 390.96 420.81 449.93
P-EB 239.38 269.23 299.08 328.93 358.79 388.64 418.49 448.34
NP-QB 240.40 271.03 300.57 330.71 360.28 390.42 418.99 442.98
Figure 5: Estimates of the number of followers expected to be generated by a tweet whose retweets have $y$ followers (in 30 seconds): NP-EB (green line), NP-ML (cyan line), NP-MD (magenta line), P-EB (red line) and NP-QB (blue line).
D.3 Real data: auto accident dataset
We consider the auto accident data of Efron and Hastie (2021, Table 6.1). This is a benchmark dataset in empirical Bayes. These are one year of claims data for a European automobile insurance company on $9461$ automobile insurance policy holders. From the first two rows of Table 7, $7840$ of the $9461$ policy holders made no claims during the year, $1317$ made a single claim, $239$ made two claims each, etc., continuing to the one person who made seven claims. Interest is in estimating the number of claims expected in a succeeding year by each automobile insurance policy holder. Table 7 displays the estimates of the number of claims expected.
Table 7: Counts $n_{y}$ of number of claims $y$ made in a single year, for $y=0,1,\ldots,7$, and corresponding estimates of the number of claims expected.
Claims $y$ 0 1 2 3 4 5 6 7
Counts $n_{y}$ 7,840 1,317 239 42 14 4 4 1
NP-EB 0.17 0.36 0.53 1.33 1.43 6.00 1.75 0
NP-ML 0.33 1.39 2.31 3.39 4.30 4.93 5.362 5.69
NP-MD 0.15 0.41 0.82 1.00 1.17 1.79 3.24 4.46
P-FB 0.16 0.42 0.68 0.94 1.20 1.47 1.75 2.03
NP-QB 0.12 0.25 0.47 0.91 1.85 3.17 4.33 5.22
E An extension to the multidimensional Poisson compound decision problem
The results in Section 2.1 extend, with minor modifications, to the $k$-dimension coordinate-wise independent Poisson model. We assume that, $\mathbf{Y}=(\mathbf{Y}_{1},\dots,\mathbf{Y}_{k})$ is a $k$-dimensional random vector with independent coordinates $\mathbf{Y}_{i}\sim{\rm Poisson}(\bm{\theta}_{i})$, and denote
$k(\mathbf{y}\mid\bm{\theta})=\prod_{j=1}^{k}e^{-\bm{\theta}_{j}}\frac{\bm{\theta}_{j}^{\mathbf{y}_{j}}}{\mathbf{y}_{j}!},\quad\mathbf{y}=(\mathbf{y}_{1},\dots,\mathbf{y}_{k})\in\mathbb{N}_{0}^{k},\bm{\theta}_{j}\in\Theta_{d}\;(j=1,\dots,k),$ (E.7)
with $\Theta_{d}=\{\vartheta_{1},\dots,\vartheta_{d}\}$. Denote $\bm{\theta}=(\bm{\theta}_{1},\dots,\bm{\theta}_{k})\in\Theta_{d}^{k}$ and $\Theta_{d}^{k}=\{{\bm{\vartheta}}_{1},\dots,{\bm{\vartheta}}_{D}\}$, $D=d^{k}$. If $\bm{\theta}$ has probability mass function $g$ on $\Theta_{d}^{k}$, then the marginal probability mass function of $\mathbf{Y}=(\mathbf{Y}_{1},\dots,\mathbf{Y}_{k})$ is
$p_{g}(\mathbf{y})=\sum_{i=1}^{D}k(\mathbf{y}\mid{\bm{\vartheta}}_{i})g({\bm{\vartheta}}_{i}).$ (E.8)
Given a sequence $(\mathbf{Y}_{n})_{n\geq 1}$ of random vectors with independent coordinates $\mathbf{Y}_{n,j}\sim{\rm Poisson}(\bm{\theta}_{n,j})$, we consider the problem of estimating the unknown parameters $\bm{\theta}_{n,j}$.
If we knew that the random vectors $(\bm{\theta}_{n})_{n\geq 1}$ were independent and identically distributed with a known probability mass function $g$, we could estimate $\bm{\theta}_{n,j}$, given that $\mathbf{Y}_{n}=\mathbf{y}$, with
$\hat{\bm{\theta}}_{g,j}(\mathbf{y})=E_{g}[\bm{\theta}_{n,j}|\mathbf{Y}_{n}=\mathbf{y}]=\frac{1}{p_{g}(\mathbf{y})}\sum_{i=1}^{D}e^{-{\bm{\vartheta}}_{i,j}}\frac{{\bm{\vartheta}}_{i,j}^{\mathbf{y}_{j}+1}}{\mathbf{y}_{j}!}\prod_{j^{\prime}\neq j}e^{-{\bm{\vartheta}}_{i,j^{\prime}}}\frac{{\bm{\vartheta}}_{i,j^{\prime}}^{\mathbf{y}_{j^{\prime}}}}{\mathbf{y}_{j^{\prime}}!}=(\mathbf{y}_{j}+1)\frac{p_{g}(\mathbf{y}+\mathbf{e}_{j})}{p_{g}(\mathbf{y})},$ (E.9)
where $\mathbf{e}_{j}$ is a $k$-dimensional vector with $\mathbf{e}_{j,i}=0$ for $i\neq j$ and $\mathbf{e}_{j,j}=1$.
Being $g$ unknown, we can estimate it recursively as
$g_{n+1}({\bm{\vartheta}}_{i})=(1-\alpha_{n+1})g_{n}({\bm{\vartheta}}_{i})+\alpha_{n+1}\frac{k(\mathbf{Y}_{n+1}\mid{\bm{\vartheta}}_{i})g_{n}({\bm{\vartheta}}_{i})}{\sum_{i^{\prime}=1}^{D}k(\mathbf{Y}_{n+1}\mid{\bm{\vartheta}}_{i^{\prime}})g_{n}({\bm{\vartheta}}_{i^{\prime}})}\qquad i=1,\dots,D,$ (E.10)
with $g_{0}({\bm{\vartheta}}_{i})>0$ for every $i=1,\dots,D$, and adopt for $\bm{\theta}_{g,j}(\mathbf{y})$ the plug-in estimator
$\hat{\bm{\theta}}_{g_{n},j}(\mathbf{y})=(\mathbf{y}_{j}+1)\frac{p_{g_{n}}(\mathbf{y}+\mathbf{e}_{j})}{p_{g_{n}}(\mathbf{y})}\qquad\mathbf{y}\in\mathbb{N}_{0}^{k},$ (E.11)
where
$p_{g_{n}}(\mathbf{y})=\sum_{i=1}^{D}k(\mathbf{y}\mid{\bm{\vartheta}}_{i})g_{n}({\bm{\vartheta}}_{i}).$ (E.12)
We now model the sequence of observations $(\mathbf{Y}_{n})$, assuming that, given the information available at time $n$, $\mathbf{Y}_{n+1}$ follows the probability mass function $p_{g_{n}}$. In other words, denoting by $(\mathcal{G}_{n})_{n\geq 0}$ the natural filtration of $(\mathbf{Y}_{n})_{n\geq 1}$, we assume that $\mathbf{Y}_{n+1}\mid\mathcal{G}_{n}\sim p_{g_{n}}$. Under this assumption, the sequence $(g_{n})_{n\geq 0}$ is a martingale with respect to $(\mathcal{G}_{n})_{n\geq 0}$, converging $\mathds{P}$ a.s. as $n\rightarrow\infty$ to a random probability mass function $\tilde{g}$ on $\Theta_{d}^{k}$. In particular, for every $\mathbf{y}\in\mathbb{N}_{0}^{k}$, as $n\rightarrow+\infty$
$\hat{\bm{\theta}}_{g_{n},j}(\mathbf{y})\rightarrow\hat{\bm{\theta}}_{\tilde{g},j}(\mathbf{y})=(\mathbf{y}_{j}+1)\frac{p_{\tilde{g}}(\mathbf{y}+\mathbf{e}_{j})}{p_{\tilde{g}}(\mathbf{y})}\qquad\mathds{P}\,\text{-}\,a.s.$ (E.13)
The next result extends Theorem 2.2 to the multidimensional case.
Theorem E.1.
If $(b_{n})_{n\geq 1}$ are strictly positive numbers with
$b_{n}\uparrow+\infty,\quad\sum_{n\geq 1}\alpha_{n}^{4}b_{n}^{2}<+\infty,\quad b_{n}\sup_{k\geq n}\alpha_{k}^{2}\rightarrow 0,\quad b_{n}\sum_{k\geq n}\alpha_{k}^{2}\rightarrow 1,$
then, for every $\mathbf{y}\in\mathbb{N}_{0}^{k}$, $b_{n}^{1/2}(\hat{\bm{\theta}}_{g_{n}}(\mathbf{y})-\hat{\bm{\theta}}_{\tilde{g}}(\mathbf{y}))$ converges, in the sense of almost-sure conditional convergence, as $n\rightarrow+\infty$, to a centered (i.e., zero-mean) Gaussian kernel with the random covariance matrix $\mathbf{W}(\mathbf{y})$, whose $(j,j^{\prime})$ element is given by
$\mathbf{W}_{\tilde{g},j,j^{\prime}}(\mathbf{y})=\hat{\bm{\theta}}_{\tilde{g},j}(\mathbf{y})\hat{\bm{\theta}}_{\tilde{g},j^{\prime}}(\mathbf{y})$ (E.14)
$E_{Z\sim p_{\tilde{g}}}\Bigl{[}\Bigl{(}\sum_{i=1}^{D}\frac{\tilde{g}({\bm{\vartheta}}_{i}\mid Z)}{\tilde{g}({\bm{\vartheta}}_{i})}\bigl{(}\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}+\mathbf{e}_{j})-\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y})\bigr{)}\Bigr{)}\Bigl{(}\sum_{i=1}^{D}\frac{\tilde{g}({\bm{\vartheta}}_{i}\mid Z)}{\tilde{g}({\bm{\vartheta}}_{i})}\bigl{(}\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}+\mathbf{e}_{j^{\prime}})-\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y})\bigr{)}\Bigr{)}\Bigr{]},$ (E.15)
See Supplementary Material F for the proof of Theorem 2.2. For $\alpha_{n}=(\alpha+n)^{-\gamma}$ with $\gamma\in(1/2,1]$, the theorem holds with $b_{n}=(2\gamma-1)/n^{2\gamma-1}$. In particular, with $\alpha_{n}=(1+n)^{-1}$, $b_{n}=n$. Theorem 2.2 provides a tool to obtain asymptotic credible regions for $\hat{\bm{\theta}}_{g}(\mathbf{y})$, for $y\in\mathbb{N}_{0}^{k}$.
We now consider the Bayesian regret
$\text{Regret}(g_{n},g^{*})=\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}||\hat{\bm{\theta}}_{g_{n}}(\mathbf{y})-\hat{\bm{\theta}}_{g^{*}}(\mathbf{y})||^{2}p_{g^{*}}(\mathbf{y}).$ (E.16)
under a “true” Poisson compound model for $(\mathbf{Y}_{n})_{n\geq 1}$, under which, for any $n\geq 1$: i) the distribution of $\mathbf{Y}_{n}$ given $\bm{\theta}^{\ast}_{n}$ is a product of Poisson distributions with parameters $\theta^{\ast}_{n,j}$; ii) the $\bm{\theta}^{\ast}_{n}$’s are independent and identically distributed from the “true” prior distribution, say the oracle prior $G^{\ast}$ on $\Theta_{d}^{k}$, with probability mass function $g^{\ast}$ on $\Theta_{d}^{k}$. The asymptotic behavior of regret for large $n$ in the multidimensional case follows the same rate of convergence as in the univariate case.
Theorem E.2.
Let $g_{n}$ be as in (E.10) with $\alpha_{n}=(\alpha+n)^{-\gamma}$ for $\alpha>0$ and $\gamma\in(1/2,1)$, and let $g^{\ast}$ be the probability mass function of the oracle prior $G^{\ast}$. Denote by $\mathds{P}^{*}$ a probability measure under which the “true” Poisson compound model for $(\mathbf{Y}_{n})_{n\geq 1}$ holds true. Then, for every $\delta<2-1/\gamma$, it holds that
$\text{Regret}(g_{n},g^{*})=o(n^{-\delta})\quad\mathds{P}^{\ast}\,\text{-}\,a.s.$
See Supplementary Material F for the proof of Theorem E.2.
F Proofs for Section E
F.1 Notation
We denote by $\Delta^{D}=\{\mathbf{v}=(\mathbf{v}_{1},\dots,\mathbf{v}_{D-1}):\mathbf{v}_{i}\geq 0,\sum_{i=1}^{D-1}\mathbf{v}_{i}\leq 1\}$ the $D$-dimensional simplex, and treat probability mass functions on $\Theta_{d}^{k}$ as vectors in $\Delta^{D}$ or in $\mathbb{R}^{D}$, by setting $\mathbf{g}_{i}=g({\bm{\vartheta}}_{i})$. Analogously, we write $\mathbf{k}_{\mathbf{y},i}=k_{\mathbf{y}}({\bm{\vartheta}}_{i})=k(\mathbf{y}\mid{\bm{\vartheta}}_{i})$ for $i=1,\dots,D$. The probability mass function $p_{g}$ is denoted as $p_{\mathbf{g}}$, when $g$ is considered as a vector. For every $g$, we denote
$g({\bm{\vartheta}}_{i}\mid\mathbf{y})=\frac{k(\mathbf{y}\mid{\bm{\vartheta}}_{i})g({\bm{\vartheta}}_{i})}{p_{g}(\mathbf{y})},$
for $\mathbf{y}\in\mathbb{N}_{0}^{k}$, $i=1,\dots,D$.
F.2 Proof of Theorem E.1
Let us define the $(D-1)\times(D-1)$ random matrices $\mathbf{V}_{\tilde{g}}$ and $\mathbf{V}_{n}$, with $(i,i^{\prime})$ elements (for $i,i^{\prime}=1,\dots,D-1$) given by
$\mathbf{V}_{\tilde{g},i,i^{\prime}}=\sum_{\mathbf{z}\in\mathbb{N}_{0}^{k}}\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{z})\tilde{g}({\bm{\vartheta}}_{i^{\prime}}\mid\mathbf{z})p_{\tilde{g}}(\mathbf{z})-\tilde{g}({\bm{\vartheta}}_{i})\tilde{g}({\bm{\vartheta}}_{i^{\prime}}),$ (F.17)
and
$\mathbf{V}_{n,i,i^{\prime}}=\sum_{\mathbf{z}\in\mathbb{N}_{0}^{k}}g_{n}({\bm{\vartheta}}_{i}\mid\mathbf{z})g_{n}({\bm{\vartheta}}_{i^{\prime}}\mid\mathbf{z})p_{g_{n}}(\mathbf{z})-g_{n}({\bm{\vartheta}}_{i})g_{n}({\bm{\vartheta}}_{i^{\prime}}).$ (F.18)
Lemma F.1.
Let $\mathbf{V}_{\tilde{g}}$ and $\mathbf{V}_{n}$ be defined as in (F.17) and (F.18). Then, $\mathds{P}$-a.s, $(V_{n})$ is a sequence of uniformly bounded, positive definite random matrices, converging to $\mathbf{V}_{\tilde{g}}$ as $n\rightarrow\infty$.
Proof.
Since, for every $i,{i^{\prime}}$
$|\mathbf{V}_{n,i,{i^{\prime}}}|\leq 1,$
then the sequence $(\mathbf{V}_{n})$ is uniformly bounded. Furthermore, for every random vector $\mathbf{u}\in\mathbb{R}^{D-1}$ with $||\mathbf{u}||=1$,
$\mathbf{u}^{T}\mathbf{V}_{n}\mathbf{u}$ $=\sum_{i,{i^{\prime}}=1}^{D-1}\sum_{\mathbf{z}\in\mathbb{N}_{0}^{k}}\mathbf{u}_{i}(g_{n}({\bm{\vartheta}}_{i}\mid\mathbf{z})-g_{n}({\bm{\vartheta}}_{i}))\mathbf{u}_{{i^{\prime}}}(g_{n}({\bm{\vartheta}}_{i^{\prime}}\mid\mathbf{z})-g({\bm{\vartheta}}_{i^{\prime}}))p_{g_{n}}(\mathbf{z})$
$=\sum_{\mathbf{z}\in\mathbb{N}_{0}^{k}}\left(\sum_{i=1}^{D-1}\mathbf{u}_{i}(g_{n}({\bm{\vartheta}}_{i}\mid\mathbf{z})-g_{n}({\bm{\vartheta}}_{i}))\right)^{2}p_{g_{n}}(\mathbf{z})\geq 0,$
$\mathds{P}$-a.s. Hence, $\mathds{P}$-a.s., $\mathbf{V}_{n}$ is positive semi-definite. We now prove that it is $\mathds{P}$-a.s. positive definite. By contradiction, suppose that there exists a random vector $\mathbf{u}$ such that for $\omega$ in a set $N^{c}$ with positive probability $\mathbf{u}(\omega)\neq\mathbf{0}$ and $\mathbf{u}^{T}(\omega)\mathbf{V}_{n}(\omega)\mathbf{u}(\omega)=0$. Fix $\omega\in N^{c}$ and, without loss of generality, assume that $p_{g_{n}(\omega)}(\mathbf{z})>0$ for every $\mathbf{z}\in\mathbb{N}_{0}^{k}$. Then, we can write (omitting $\omega$ for simplicity)
$\sum_{i=1}^{D-1}\mathbf{u}_{i}(g_{n}({\bm{\vartheta}}_{i}\mid\mathbf{z})-g_{n}({\bm{\vartheta}}_{i}))=0,$
for every $\mathbf{z}\in\mathbb{N}_{0}^{k}$. Define $u({\bm{\vartheta}}_{i})=\mathbf{u}_{i}$ for $i=1,\dots,D-1$ and $u({\bm{\vartheta}}_{D})=0$. Then, for every $\mathbf{z}=(\mathbf{z}_{1},\dots,\mathbf{z}_{k})\in\mathbb{N}_{0}^{k}$,
$\sum_{i=1}^{D}u({\bm{\vartheta}}_{i})\prod_{j=1}^{k}e^{-{\bm{\vartheta}}_{i,j}}{\bm{\vartheta}}_{i,j}^{\mathbf{z}_{j}}g_{n}({\bm{\vartheta}}_{i})=\sum_{i=1}^{D}\prod_{j=1}^{k}e^{-{\bm{\vartheta}}_{i,j}}{\bm{\vartheta}}_{i,j}^{\mathbf{z}_{j}}g_{n}({\bm{\vartheta}}_{i})\sum_{i=1}^{D}u({\bm{\vartheta}}_{i})g_{n}({\bm{\vartheta}}_{i}),$
which implies that, for every $\mathbf{x}\in\mathbb{N}_{0}^{k}$,
$\sum_{i=1}^{D}u({\bm{\vartheta}}_{i})e^{\langle\mathbf{x},{\bm{\vartheta}}_{i}\rangle}g_{n}({\bm{\vartheta}}_{i})=\sum_{i=1}^{D}e^{\langle\mathbf{x},{\bm{\vartheta}}_{i}\rangle}g_{n}({\bm{\vartheta}}_{i})\sum_{i=1}^{D}u({\bm{\vartheta}}_{i})g_{n}({\bm{\vartheta}}_{i}).$ (F.19)
To demonstrate that this leads to a contradiction, consider a random vector $\tilde{\bm{\theta}}$ with a probability mass function $g_{n}$ defined on $\Theta_{d}^{k}$. Since $g_{n}({\bm{\vartheta}}_{i})>0$ for every $i$, the random variables $(e^{\langle\mathbf{x},\tilde{\bm{\theta}}_{n}\rangle})_{\mathbf{x}\in\mathbb{N}_{0}^{k}}$ are linearly independent. Furthermore, $u(\tilde{\bm{\theta}})$ cannot be constant because $u({\bm{\vartheta}}_{D})=0$ and $\mathbf{u}$ is not the zero vector. Therefore, equation (F.19) can not hold.
We now prove that $\mathbf{V}_{n}$ converge to $\mathbf{V}$, $P$-a.s. Since the matrices $\mathbf{V}_{n}$ and $\mathbf{V}_{\tilde{g}}$ have equal finite dimensions, then the convergence holds if and only if for every $i$ and $i^{\prime}$
$\mathbf{V}_{n,i,i^{\prime}}\rightarrow\mathbf{V}_{\tilde{g},i,i^{\prime}}\quad\mathds{P}-a.s.,$
which is true since $g_{n}$ converges to $\tilde{g}$ and $\mathbf{V}_{n,i,i^{\prime}}$ is a continuous function of $g_{n}$. ∎
Theorem F.2.
Let $(\mathbf{g}_{n})$ and $\tilde{\mathbf{g}}$ be the representation of $(g_{n})$ and of $\tilde{g}$ in $\Delta^{D}$, respectively. If $(b_{n})$ is a sequence of strictly positive numbers with
$b_{n}\uparrow+\infty,\quad\sum_{n\geq 1}\alpha_{n}^{4}b_{n}^{2}<+\infty,\quad b_{n}\sup_{m\geq n}\alpha_{m}^{2}\rightarrow 0,\quad b_{n}\sum_{m\geq n}\alpha_{m}^{2}\rightarrow 1,$ (F.20)
then, as $n\rightarrow+\infty$, $b_{n}^{1/2}(\mathbf{g}_{n}-\tilde{\mathbf{g}})$ converges, in the sense of almost-sure conditional convergence, to a centered Gaussian kernel with random covariance matrix $\mathbf{V}_{\tilde{g}}$ with $(i,i^{\prime})$ element $\mathbf{V}_{\tilde{g},i,i^{\prime}}$ as in (F.17).
Proof.
The proof is based on Theorem B.2. Fix a vector $\mathbf{u}$ in $\mathbb{R}^{D-1}$ such that $||\mathbf{u}||=1$. For each $n\geq 0$, define
$\mathcal{F}_{n,0}=\mathcal{G}_{n},\quad\mathcal{F}_{n,j}=\mathcal{G}_{n+j-1}\mbox{ for }j\geq 1,$
where $(\mathcal{G}_{n})_{n\geq 0}$ is the natural filtration of $(\mathbf{Y}_{n})_{n\geq 1}$,
$M_{n,0}=0,\quad M_{n,j}=b_{n}^{1/2}(\langle\mathbf{g}_{n},\mathbf{u}\rangle-\langle\mathbf{g}_{n+j-1},\mathbf{u}\rangle)\mbox{ for }j\geq 1.$
For every $n\geq 0$, the sequence $(M_{n,j})_{j\geq 0}$ is a bounded martingale with respect to the filtration $(\mathcal{F}_{n,j})_{j\geq 0}$, converging in $L^{1}$ to
$M_{n,\infty}=b_{n}^{1/2}\langle\mathbf{g}_{n}-\tilde{\mathbf{g}},\mathbf{u}\rangle.$
Let
$X_{n,j}=M_{n,j}-M_{n,j-1}=b_{n}^{1/2}\langle\mathbf{g}_{n+j-1}-\mathbf{g}_{n+j},\mathbf{u}\rangle,$
$X_{n}^{*}=\sup_{j\geq 1}|X_{n,j}|=b_{n}^{1/2}\sup_{j\geq 1}|\langle\mathbf{g}_{n+j-1}-\mathbf{g}_{n+j},\mathbf{u}\rangle|,$
and
$U_{n}=\sum_{j=0}^{\infty}X_{n,j}^{2}=b_{n}\sum_{m\geq n}(\langle\mathbf{g}_{m}-\mathbf{g}_{m-1},\mathbf{u}\rangle)^{2}.$
The thesis follows from Theorem B.2, if we can verify that
  • (a)
    the sequence $(X^{*}_{n})$ is dominated in $L^{1}$ and converges to zero $\mathds{P}$-a.s. as $n\rightarrow+\infty$.
  • (b)
    $U_{n}\rightarrow\mathbf{u}^{T}\mathbf{V}_{\tilde{g}}\mathbf{u}$, $P$-a.s. as $n\rightarrow+\infty$.
(a) Notice that $\langle\mathbf{g}_{m+1}-\mathbf{g}_{m},\mathbf{u}\rangle=\alpha_{m}\sum_{i=1}^{D-1}(g_{m-1}({\bm{\vartheta}}_{i}\mid\mathbf{Y}_{m})-g_{m-1}({\bm{\vartheta}}_{i}))\mathbf{u}_{i}$ is bounded. Thus, $X^{*}_{n}\leq b_{n}^{1/2}\sup_{m\geq n}\alpha_{m}|\sum_{i=1}^{D-1}(\langle g_{m-1}({\bm{\vartheta}}_{i}\mid\mathbf{Y}_{m})-g_{m-1}({\bm{\vartheta}}_{i})\mathbf{u}_{i}|$ is uniformly bounded and converges to zero $\mathds{P}$-a.s., since $b_{n}\sup_{m\geq n}\alpha_{m}^{2}\rightarrow 0$ by assumption.
(b) The proof is based on Theorem B.4 (b). Let $Z_{n}=\alpha_{n}^{-2}(\langle\mathbf{g}_{n}-\mathbf{g}_{n-1},\mathbf{u}\rangle)^{2},$ and $a_{n}=(b_{n}\alpha_{n})^{-2}$. We can write that
$E_{n-1}(Z_{n})$ $=E_{n-1}\left(\alpha_{n}^{-2}(\langle\mathbf{g}_{n}-\mathbf{g}_{n-1},\mathbf{u}\rangle)^{2}\right)$
$=\sum_{\mathbf{z}\in\mathbb{N}_{0}^{k}}\sum_{i,{i^{\prime}}=1}^{D-1}\mathbf{u}_{i}\mathbf{u}_{i^{\prime}}\left(g_{n}({\bm{\vartheta}}_{i}\mid\mathbf{z})g_{n}({\bm{\vartheta}}_{i^{\prime}}\mid\mathbf{z})-g_{n}({\bm{\vartheta}}_{i})g_{n}({\bm{\vartheta}}_{i^{\prime}})\right)p_{g_{n}}(\mathbf{z})=\mathbf{u}^{T}\mathbf{V}_{n}\mathbf{u}.$
Thus, by Lemma F.1,
$E_{n-1}\left(Z_{n}\right)\rightarrow\mathbf{\mathbf{u}}^{T}\mathbf{V}_{\tilde{g}}\mathbf{\mathbf{u}},\quad\mathds{P}-a.s.$
Moreover, $b_{n}\uparrow+\infty$ by assumption, and
$\sum_{m=1}^{\infty}\frac{E(Z_{m}^{2})}{a_{m}^{2}b_{m}^{2}}=\sum_{m=1}^{\infty}\alpha_{m}^{4}b_{m}^{2}E\left(\left[\sum_{i=1}^{D-1}g_{m-1}({\bm{\vartheta}}_{i}\mid\mathbf{Y}_{m})-g_{m-1}({\bm{\vartheta}}_{i})\mathbf{u}_{i}\right]^{4}\right)<+\infty.$
Since, by assumption, $b_{n}\sum_{m\geq n}a_{m}^{-1}b_{m}^{-2}\rightarrow 1$, then by Theorem B.4 (b),
$b_{n}\sum_{m\geq n}(\langle\mathbf{g}_{m}-\mathbf{g}_{m-1},\mathbf{u}\rangle)^{2}=b_{n}\sum_{m\geq n}\frac{Z_{m}}{a_{m}b_{m}^{2}}\rightarrow\mathbf{u}^{T}\mathbf{V}_{\tilde{g}}\mathbf{u},\quad\mathds{P}-a.s.$
It follows by Theorem B.2 that, as $n\rightarrow\infty$, $M_{n,\infty}=b_{n}^{1/2}(\langle\mathbf{g}_{n}-\tilde{\mathbf{g}},\mathbf{u}\rangle)$ converges to $\mathcal{N}(0,\mathbf{u}^{T}\mathbf{V}_{\tilde{g}}\mathbf{u})$ in the sense of almost-sure conditional convergence. Since this is true for every $\mathbf{u}$ with $||\mathbf{u}||=1$, then $b_{n}^{1/2}(\mathbf{g}_{n}-\tilde{\mathbf{g}})$ converges in the sense of almost sure conditional convergence to $\mathcal{N}(0,\mathbf{V}_{\tilde{g}})$.
Now suppose that $\alpha_{n}=(\alpha+n)^{-\gamma}$. With $b_{n}=(2\gamma-1)n^{2\gamma-1}$, the conditions (F.20) hold. Indeed, $b_{n}\uparrow+\infty$,
$\sum_{n=1}^{\infty}\alpha_{n}^{4}b_{n}^{2}=\sum_{n=1}^{\infty}\left(\frac{1}{\alpha+n}\right)^{4\gamma}(2\gamma-1)^{2}n^{4\gamma-2}<+\infty,$
$b_{n}\sup_{m\geq n}\alpha_{m}^{2}=(2\gamma-1)n^{2\gamma-1}\left(\frac{1}{\alpha+n}\right)^{2\gamma}\rightarrow 0,$
and
$b_{n}\sum_{m\geq n}{\alpha_{m}^{2}}=(2\gamma-1)n^{2\gamma-1}\sum_{m\geq n}\left(\frac{1}{\alpha+m}\right)^{2\gamma}\rightarrow 1.$
of Theorem E.1.
The proof is based on Theorem F.2 and Theorem B.5. We treat $\hat{\bm{\theta}}_{g}(\mathbf{y})$ as a vector-valued function of $\mathbf{g}\in\Delta^{D}$, where, for $j=1,\dots,k$, the $j$th component is given by
$\hat{\theta}_{g,j}(\mathbf{y})=\mathbf{h}^{(j)}_{\mathbf{y}}(\mathbf{g})=(\mathbf{y}_{j}+1)\frac{p_{g}(\mathbf{y}+\mathbf{e}_{j})}{p_{g}(\mathbf{y})}=(\mathbf{y}_{j}+1)\frac{\sum_{i=1}^{D-1}k_{\mathbf{y}+\mathbf{e}_{j}}({\bm{\vartheta}}_{i})\mathbf{g}({\bm{\vartheta}}_{i})+k_{\mathbf{y}+\mathbf{e}_{j}}({\bm{\vartheta}}_{D})(1-\sum_{i=1}^{D-1}\mathbf{g}({\bm{\vartheta}}_{i}))}{\sum_{i=1}^{D-1}k_{\mathbf{y}}({\bm{\vartheta}}_{i})\mathbf{g}({\bm{\vartheta}}_{i})+k_{\mathbf{y}}({\bm{\vartheta}}_{D})(1-\sum_{i=1}^{D-1}\mathbf{g}({\bm{\vartheta}}_{i}))}.$
The function $\mathbf{h}_{\mathbf{y}}=[\mathbf{h}^{(1)}_{\mathbf{y}}(\mathbf{g}),\dots,\mathbf{h}^{(k)}_{\mathbf{y}}(\mathbf{g})]^{T}$ is differentiable with respect to $\mathbf{g}$, and, for $i=1,\dots,D-1$,
$\frac{\partial\mathbf{h}^{(j)}_{\mathbf{y}}(\mathbf{g})}{\partial\mathbf{g}_{i}}$ $=(\mathbf{y}_{j}+1)\frac{k_{\mathbf{y}+\mathbf{e}_{j}}({\bm{\vartheta}}_{i})-k_{\mathbf{y}+\mathbf{e}_{j}}({\bm{\vartheta}}_{D})}{\sum_{i=1}^{D-1}k_{\mathbf{y}}({\bm{\vartheta}}_{i})\mathbf{g}({\bm{\vartheta}}_{i})+k_{\mathbf{y}}[D](1-\sum_{i=1}^{D-1}\mathbf{g}({\bm{\vartheta}}_{i}))}$
$-(\mathbf{y}_{j}+1)\frac{(\sum_{i=1}^{D-1}k_{\mathbf{y}+\mathbf{e}_{j}}({\bm{\vartheta}}_{i})\mathbf{g}({\bm{\vartheta}}_{i})+k_{\mathbf{y}+\mathbf{e}_{j}}({\bm{\vartheta}}_{D})(1-\sum_{i=1}^{D-1}\mathbf{g}({\bm{\vartheta}}_{i})))(k_{\mathbf{y}}({\bm{\vartheta}}_{i})-k_{\mathbf{y}}({\bm{\vartheta}}_{D}))}{(\sum_{i=1}^{D-1}k_{\mathbf{y}}({\bm{\vartheta}}_{i})\mathbf{g}({\bm{\vartheta}}_{i})+k_{\mathbf{y}}({\bm{\vartheta}}_{D})(1-\sum_{i=1}^{D-1}\mathbf{g}({\bm{\vartheta}}_{i})))^{2}}$
$=\frac{\mathbf{y}_{j}+1}{p_{g}(\mathbf{y})}\left[k_{\mathbf{y}+\mathbf{e}_{j}}({\bm{\vartheta}}_{i})-k_{\mathbf{y}+\mathbf{e}_{j}}({\bm{\vartheta}}_{D})-\frac{p_{g}(\mathbf{y}+\mathbf{e}_{j})}{p_{g}(\mathbf{y})}(k_{\mathbf{y}}({\bm{\vartheta}}_{i})-k_{\mathbf{y}}({\bm{\vartheta}}_{D}))\right]$
$=\hat{\theta}_{g,j}(\mathbf{y})\left[\frac{k_{\mathbf{y}+\mathbf{e}_{j}}({\bm{\vartheta}}_{i})-k_{\mathbf{y}+\mathbf{e}_{j}}({\bm{\vartheta}}_{D})}{p_{g}(\mathbf{y}+\mathbf{e}_{j})}-\frac{k_{\mathbf{y}}({\bm{\vartheta}}_{i})-k_{\mathbf{y}}({\bm{\vartheta}}_{D})}{p_{g}(\mathbf{y})}\right].$
Let us denote by $\nabla_{\mathbf{g}}h^{(j)}_{\mathbf{y}}(g)$ the row vector of partial derivatives of $\mathbf{h}^{(j)}_{\mathbf{y}}$. Since $\Theta_{d}^{k}$ is finite, then the partial derivatives are bounded, uniformly with respect to $\mathbf{g}\in\Delta^{D}$. Thus, $\mathbf{h}^{(j)}_{\mathbf{y}}(\mathbf{g})$ is uniformly differentiable. By Theorem F.2, $b_{n}^{1/2}(\mathbf{g}_{n}-\tilde{\mathbf{g}})$ converges in the sense of almost sure conditional convergence to a centered normal kernel $\mathcal{N}(0,\mathbf{V}_{\tilde{g}})$, with random covariance matrix $\mathbf{V}_{\tilde{g}}$ as in (F.17). Denote by $D_{\tilde{\mathbf{g}}}\mathbf{h}_{\mathbf{y}}(\tilde{\mathbf{g}})$ the Jacobian matrix of $\mathbf{h}_{\mathbf{y}}$. By Theorem B.5, $b_{n}^{1/2}(\mathbf{h}_{\mathbf{y}}(\mathbf{g}_{n})-\mathbf{h}_{\mathbf{y}}(\tilde{\mathbf{g}}))$ converges in the sense of almost sure conditional convergence, to a Gaussian kernel with covariance matrix $(D_{\tilde{\mathbf{g}}}\mathbf{h}_{\mathbf{y}}(\tilde{\mathbf{g}}))\mathbf{V}_{\tilde{\mathbf{g}}}(D_{\tilde{\mathbf{g}}}\mathbf{h}_{\tilde{\mathbf{g}}})^{T}$. Its $(j,j^{\prime})$th element can be written as
$((D_{\tilde{\mathbf{g}}}\mathbf{h}_{\mathbf{y}}(\tilde{\mathbf{g}}))\mathbf{V}_{\tilde{\mathbf{g}}}(D_{\tilde{\mathbf{g}}}\mathbf{h}_{\tilde{\mathbf{g}}})^{T})_{j,j^{\prime}}=\hat{\theta}_{\tilde{\mathbf{g}},j}(\mathbf{y})\hat{\theta}_{\tilde{\mathbf{g}},j^{\prime}}(\mathbf{y})$
$\sum_{i,i^{\prime}=1}^{D-1}\left[\frac{k_{\mathbf{y}+\mathbf{e}_{j}}({\bm{\vartheta}}_{i})-k_{\mathbf{y}+\mathbf{e}_{j}}({\bm{\vartheta}}_{D})}{p_{\tilde{g}}(\mathbf{y}+\mathbf{e}_{j})}-\frac{k_{\mathbf{y}}({\bm{\vartheta}}_{i})-k_{\mathbf{y}}({\bm{\vartheta}}_{D})}{p_{\tilde{g}}(\mathbf{y})}\right]\left[\sum_{\mathbf{z}\in\mathbb{N}_{0}^{k}}\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{z})\tilde{g}({\bm{\vartheta}}_{i^{\prime}}\mid\mathbf{z})p_{\tilde{g}}(\mathbf{z})-\tilde{g}({\bm{\vartheta}}_{i})\tilde{g}({\bm{\vartheta}}_{i^{\prime}})\right]$
$\left[\frac{k_{\mathbf{y}+\mathbf{e}_{j^{\prime}}}({\bm{\vartheta}}_{i^{\prime}})-k_{\mathbf{y}+\mathbf{e}_{j^{\prime}}}({\bm{\vartheta}}_{D})}{p_{\tilde{g}}(\mathbf{y}+\mathbf{e}_{j^{\prime}})}-\frac{k_{\mathbf{y}}({\bm{\vartheta}}_{i^{\prime}})-k_{\mathbf{y}}({\bm{\vartheta}}_{D})}{p_{\tilde{g}}(\mathbf{y})}\right]$
$=\hat{\theta}_{\tilde{g},j}(\mathbf{y})\hat{\theta}_{\tilde{g},j^{\prime}}(\mathbf{y})\sum_{\mathbf{z}\in\mathbb{N}_{0}^{k}}p_{\tilde{g}}(\mathbf{z})\left[\sum_{i=1}^{D-1}\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{z})(\frac{\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}+\mathbf{e}_{j})}{\tilde{g}({\bm{\vartheta}}_{i})}-\frac{\tilde{g}({\bm{\vartheta}}_{D}\mid\mathbf{y}+\mathbf{e}_{j})}{\tilde{g}({\bm{\vartheta}}_{D})}-\frac{\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y})}{\tilde{g}({\bm{\vartheta}}_{i})}+\frac{\tilde{g}({\bm{\vartheta}}_{D}\mid\mathbf{y})}{\tilde{g}(\theta_{D})})\right]$
$\left[\sum_{i=1}^{D-1}\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{z})(\frac{\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}+\mathbf{e}_{j^{\prime}})}{\tilde{g}({\bm{\vartheta}}_{i})}-\frac{\tilde{g}({\bm{\vartheta}}_{D}\mid\mathbf{y}+\mathbf{e}_{j^{\prime}})}{\tilde{g}({\bm{\vartheta}}_{D})}-\frac{\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y})}{\tilde{g}({\bm{\vartheta}}_{i})}+\frac{\tilde{g}({\bm{\vartheta}}_{D}\mid\mathbf{y})}{\tilde{g}(\theta_{D})})\right]$
$-\hat{\theta}_{\tilde{g},j}(\mathbf{y})\hat{\theta}_{\tilde{g},j^{\prime}}(\mathbf{y})\left[\sum_{i=1}^{D-1}\tilde{g}({\bm{\vartheta}}_{i})(\frac{\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}+\mathbf{e}_{j})}{\tilde{g}({\bm{\vartheta}}_{i})}-\frac{\tilde{g}({\bm{\vartheta}}_{D}\mid\mathbf{y}+\mathbf{e}_{j})}{\tilde{g}({\bm{\vartheta}}_{D})}-\frac{\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y})}{\tilde{g}({\bm{\vartheta}}_{i})}+\frac{\tilde{g}({\bm{\vartheta}}_{D}\mid\mathbf{y})}{\tilde{g}({\bm{\vartheta}}_{D})})\right]$
$\left[\sum_{i=1}^{D-1}\tilde{g}({\bm{\vartheta}}_{i})(\frac{\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}+\mathbf{e}_{j^{\prime}})}{\tilde{g}({\bm{\vartheta}}_{i})}-\frac{\tilde{g}({\bm{\vartheta}}_{D}\mid\mathbf{y}+\mathbf{e}_{j^{\prime}})}{\tilde{g}({\bm{\vartheta}}_{D})}-\frac{\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y})}{\tilde{g}({\bm{\vartheta}}_{i})}+\frac{\tilde{g}({\bm{\vartheta}}_{D}\mid\mathbf{y})}{\tilde{g}({\bm{\vartheta}}_{D})})\right]$
$=\hat{\theta}_{\tilde{g},j}(\mathbf{y})\hat{\theta}_{\tilde{g},j^{\prime}}(\mathbf{y})\sum_{\mathbf{z}\in\mathbb{N}_{0}^{k}}p_{\tilde{g}}(\mathbf{z})\left[\sum_{i=1}^{D-1}\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{z})(\frac{\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}+\mathbf{e}_{j})-\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y})}{\tilde{g}({\bm{\vartheta}}_{i})})-(1-\tilde{g}({\bm{\vartheta}}_{D}\mid\mathbf{z}))\frac{\tilde{g}({\bm{\vartheta}}_{D}\mid\mathbf{y}+\mathbf{e}_{j})-\tilde{g}({\bm{\vartheta}}_{d}\mid\mathbf{y})}{\tilde{g}({\bm{\vartheta}}_{D})}\right]$
$\left[\sum_{i=1}^{D-1}\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{z})(\frac{\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}+\mathbf{e}_{j^{\prime}})-\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y})}{\tilde{g}({\bm{\vartheta}}_{i})})-(1-\tilde{g}({\bm{\vartheta}}_{D}\mid\mathbf{z}))\frac{\tilde{g}({\bm{\vartheta}}_{D}\mid\mathbf{y}+\mathbf{e}_{j^{\prime}})-\tilde{g}({\bm{\vartheta}}_{d}\mid\mathbf{y})}{\tilde{g}({\bm{\vartheta}}_{D})}\right]$
$-\hat{\theta}_{\tilde{g},j}(\mathbf{y})\hat{\theta}_{\tilde{g},j^{\prime}}(\mathbf{y})\left[\sum_{i=1}^{D-1}\tilde{g}({\bm{\vartheta}}_{i})\frac{\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}+\mathbf{e}_{j})-\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y})}{\tilde{g}({\bm{\vartheta}}_{i})}-(1-\tilde{g}({\bm{\vartheta}}_{D}))\frac{\tilde{g}({\bm{\vartheta}}_{D}\mid\mathbf{y}+\mathbf{e}_{j})-\tilde{g}({\bm{\vartheta}}_{D}\mid\mathbf{y})}{\tilde{g}({\bm{\vartheta}}_{D})}\right]$
$\left[\sum_{i=1}^{D-1}\tilde{g}({\bm{\vartheta}}_{i})\frac{\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}+\mathbf{e}_{j^{\prime}})-\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y})}{\tilde{g}({\bm{\vartheta}}_{i})}-(1-\tilde{g}({\bm{\vartheta}}_{D}))\frac{\tilde{g}({\bm{\vartheta}}_{D}\mid\mathbf{y}+\mathbf{e}_{j^{\prime}})-\tilde{g}({\bm{\vartheta}}_{D}\mid\mathbf{y})}{\tilde{g}({\bm{\vartheta}}_{D})}\right]$
$=\hat{\theta}_{\tilde{g},j}(\mathbf{y})\hat{\theta}_{\tilde{g},j^{\prime}}(\mathbf{y})\sum_{\mathbf{z}\in\mathbb{N}_{0}^{k}}p_{\tilde{g}}(\mathbf{z})\left[\sum_{i=1}^{D}\frac{\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{z})}{\tilde{g}({\bm{\vartheta}}_{i})}(\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}+\mathbf{e}_{j})-\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}))\right]\left[\sum_{i=1}^{D}\frac{\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{z})}{\tilde{g}({\bm{\vartheta}}_{i})}(\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}+\mathbf{e}_{j^{\prime}})-\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}))\right]$
$-\hat{\theta}_{\tilde{g},j}(\mathbf{y})\hat{\theta}_{\tilde{g},j^{\prime}}(\mathbf{y})\left[\sum_{i=1}^{D}\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}+\mathbf{e}_{j})-\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y})\right]\left[\sum_{i=1}^{D}\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}+\mathbf{e}_{j^{\prime}})-\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y})\right]$
$=\hat{\theta}_{\tilde{g},j}(\mathbf{y})\hat{\theta}_{\tilde{g},j^{\prime}}(\mathbf{y})\sum_{\mathbf{z}\in\mathbb{N}_{0}^{k}}p_{\tilde{g}}(\mathbf{z})\left[\sum_{i=1}^{D}\frac{\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{z})}{\tilde{g}({\bm{\vartheta}}_{i})}(\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}+\mathbf{e}_{j})-\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}))\right]\left[\sum_{i=1}^{D}\frac{\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{z})}{\tilde{g}({\bm{\vartheta}}_{i})}(\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}+\mathbf{e}_{j^{\prime}})-\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}))\right]$
F.3 Proof of Theorem E.2
The proof of Theorem E.2 is based on the following lemmas.
Lemma F.3.
Let $(\mathbf{g}_{n})$ and $\mathbf{g}^{*}$ be the representations of $(g_{n})$ and $g^{*}$ in $\mathbb{R}^{D}$, respectively. Under the assumptions of Theorem E.2, there exists a constant $C=C(d,k,\vartheta_{1},\vartheta_{d})$ such that for every $n\geq 1$
${\rm Regret}(g_{n},g^{*})\leq C||\mathbf{g}_{n}-\mathbf{g}^{*}||^{2}.$
Proof.
We can write that
${\rm Regret}(g_{n},g^{*})$
$=\sum_{l=1}^{k}\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}(\hat{\theta}_{g_{n},l}(\mathbf{y})-\hat{\theta}_{g^{*},l}(\mathbf{y}))^{2}p_{g^{*}}(\mathbf{y})$
$=\sum_{l=1}^{k}\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\left(\frac{\sum_{i=1}^{D}{{\bm{\vartheta}}_{i,l}}\prod_{j=1}^{k}e^{-{\bm{\vartheta}}_{i,j}}{\bm{\vartheta}}_{i,j}^{y_{j}}g_{n}({\bm{\vartheta}}_{i})}{\sum_{i=1}^{D}\prod_{j=1}^{k}{\bm{\vartheta}}_{i,j}^{y_{j}}e^{-{\bm{\vartheta}}_{i,j}}g_{n}({\bm{\vartheta}}_{i})}-\frac{\sum_{i=1}^{D}{\bm{\vartheta}}_{i,l}\prod_{j=1}^{k}e^{-{\bm{\vartheta}}_{i,j}}{\bm{\vartheta}}_{i,j}^{y_{j}}g^{*}(\bm{\theta})}{\sum_{i=1}^{D}\prod_{j=1}^{k}{\bm{\vartheta}}_{i,j}^{y_{j}}e^{-{\bm{\vartheta}}_{i,j}}g^{*}({\bm{\vartheta}}_{i})}\right)^{2}$
×$\quad\quad\times\frac{1}{\prod_{j=1}^{k}y_{j}!}\sum_{i=1}^{D}\prod_{j=1}^{k}{\bm{\vartheta}}_{i,j}^{y_{j}}e^{-{\bm{\vartheta}}_{i,j}}g^{*}({\bm{\vartheta}}_{i})$
$\leq 2\sum_{l=1}^{k}\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{1}{\prod_{j=1}^{k}y_{j}!}\frac{\left(\sum_{i=1}^{D}e^{-2{\bm{\vartheta}}_{i,l}}{\bm{\vartheta}}_{i,l}^{2y_{l}+1}\prod_{j\neq l}e^{-2{\bm{\vartheta}}_{i,j}}{\bm{\vartheta}}_{i,j}^{2y_{j}}\left(g_{n}({\bm{\vartheta}}_{i})-g^{*}({\bm{\vartheta}}_{i})\right)\right)^{2}}{\sum_{i=1}^{D}\prod_{j=1}^{k}{\bm{\vartheta}}_{i,j}^{y_{j}}e^{-{\bm{\vartheta}}_{i,j}}g^{*}({\bm{\vartheta}}_{i})}$
$+2\sum_{l=1}^{k}\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{1}{\prod_{j=1}^{k}y_{j}!}\frac{\left(\sum_{i=1}^{D}{\bm{\vartheta}}_{i,l}^{y_{l}+1}e^{-{\bm{\vartheta}}_{i,l}}\prod_{j\neq l}{\bm{\vartheta}}_{i,j}^{y_{j}}e^{-{\bm{\vartheta}}_{i,j}}g_{n}({\bm{\vartheta}}_{i})\right)^{2}}{\left(\sum_{i=1}^{D}\prod_{j=1}^{k}{\bm{\vartheta}}_{i,j}^{y_{j}}e^{-{\bm{\vartheta}}_{i,j}}g_{n}({\bm{\vartheta}}_{i})\right)^{2}\sum_{i=1}^{D}\prod_{j=1}^{k}{\bm{\vartheta}}_{i,j}^{y_{j}}e^{-{\bm{\vartheta}}_{i,j}}g^{*}({\bm{\vartheta}}_{i})}$
×$\quad\quad\times\left(\sum_{i=1}^{D}\prod_{j=1}^{k}{\bm{\vartheta}}_{i,j}^{y_{j}}e^{-{\bm{\vartheta}}_{i,j}}(g_{n}({\bm{\vartheta}}_{i})-g^{*}({\bm{\vartheta}}_{i})\right)^{2}$
$\leq 2\sum_{l=1}^{k}\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{1}{\prod_{j=!}^{k}y_{j}!}\frac{\sum_{i=1}^{D}{\bm{\vartheta}}_{i,l}^{2y_{l}+2}e^{-2{\bm{\vartheta}}_{i,l}}\prod_{j\neq l}{\bm{\vartheta}}_{i,j}^{2y_{j}}e^{-2{\bm{\vartheta}}_{i,j}}||\mathbf{g}_{n}-\mathbf{g}^{*}||^{2}}{\sum_{i=1}^{D}\prod_{j=1}^{k}{\bm{\vartheta}}_{i,j}^{y_{j}}e^{-{\bm{\vartheta}}_{i,j}}g^{*}({\bm{\vartheta}}_{i})}$
$\quad\quad+2\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{1}{\prod_{j=!}^{k}y_{j}!}\frac{{\bm{\vartheta}}_{id}^{2}\sum_{i=1}^{D}\prod_{j=1}^{k}{\bm{\vartheta}}_{i,j}^{2y_{j}}e^{-2{\bm{\vartheta}}_{i,j}}||\mathbf{g}_{n}-\mathbf{g}^{*}||^{2}}{\sum_{i=1}^{D}\prod_{j=1}^{k}{\bm{\vartheta}}_{i,j}^{y_{j}}e^{-{\bm{\vartheta}}_{i,j}}g^{*}({\bm{\vartheta}}_{i})}$
$\leq 4\vartheta_{d}^{2}||\mathbf{g}_{n}-\mathbf{g}^{*}||^{2}\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{1}{\prod_{j=1}^{k}y_{j}!}\frac{\sum_{i=1}^{D}\prod_{j=1}^{k}{\bm{\vartheta}}_{i,j}^{2y_{j}}e^{-2{\bm{\vartheta}}_{i,j}}}{\sum_{i=1}^{D}\prod_{j=1}^{k}{\bm{\vartheta}}_{i,j}^{y_{j}}e^{-{\bm{\vartheta}}_{i,j}}g^{*}({\bm{\vartheta}}_{i})}$
$\leq C||\mathbf{g}_{n}-\mathbf{g}^{*}||^{2},$
with $C=4D\vartheta_{d}^{2}exp(k(\vartheta_{d}^{2}/\vartheta_{1}-2\vartheta_{1}+\vartheta_{d})).$
Lemma F.4.
Under the assumptions of Theorem E.2, the stochastic approximation in $\Delta^{D}$
$\mathbf{g}_{n+1}=\mathbf{g}_{n}+\alpha_{n+1}\mathbf{h}(\mathbf{g}_{n})+\alpha_{n+1}{\bm{\epsilon}}_{n+1},$ (F.21)
with $\alpha_{n}>0$, $\sum_{n=1}^{\infty}\alpha_{n}=\infty$, $\sum_{n=1}^{\infty}\alpha_{n}^{2}<\infty$,
$\mathbf{h}(\mathbf{g})=\mathbf{g}\circ\left(\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{k_{\mathbf{y}}}{p_{\mathbf{g}}(\mathbf{y})}p_{\mathbf{g}^{*}}(\mathbf{y})-\bf 1\right),$
${\bm{\epsilon}}_{n+1}=\mathbf{g}_{n}\circ\left(\frac{k_{\mathbf{Y}_{n+1}}}{p_{\mathbf{g}_{n}}(\mathbf{Y}_{n+1})}-\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{k_{\mathbf{y}}}{p_{\hat{\mathbf{g}}_{n}}(\mathbf{y})}p_{\mathbf{g}^{*}}(\mathbf{y})\right),$
and initial state $\mathbf{g}_{0}\in\mathring{(\Delta^{D})}$, converges $\mathds{P}^{*}$-a.s. to $\mathbf{g}^{*}$ as $n\rightarrow\infty$.
Proof.
Let $J=\partial\Delta^{D}\cup\{\mathbf{g}^{*}\}$. We first prove that ${\rm d}(\mathbf{g}_{n},J)\rightarrow 0$, $\mathds{P}^{*}$-a.s., as $n\rightarrow\infty$, where ${\rm d}$ denotes the Euclidean distance in $\mathbb{R}^{D-1}$. The result follows from Theorem 2.2.1′′ in Chen (2005) if we can prove that
  • A1
    $\alpha_{n}>0$, $\alpha_{n}\rightarrow 0$, as $n\rightarrow\infty$ and $\sum_{n=1}^{\infty}\alpha_{n}=\infty$,
  • A2
    $J$ is closed and there exists a is continuously differentiable function $\mathscr{l}:\Delta^{D}\rightarrow\mathbb{R}$ such that $\sup_{\delta_{1}\leq{\rm d}(\mathbf{g},J)\leq\delta_{2}}\mathbf{h}(\mathbf{g})^{T}\nabla_{\mathbf{g}}\mathscr{l}<0$ for every $\delta_{2}>\delta_{1}>0$,
  • A3
    $\lim_{T\rightarrow\infty}\limsup_{n\rightarrow\infty}\frac{1}{T}||\sum_{i=n}^{m(n,T)}\alpha_{i+1}{\bm{\epsilon}}_{i+1}||=0,$ $\mathds{P}^{*}$-a.s., where $m(n,T)=\max\{m:\sum_{i=n}^{m}\alpha_{i+1}\leq T\}$,
  • A4
    $\mathbf{h}$ is measurable and locally bounded.
Let $\mathscr{l}$ be defined as
$\mathscr{l}(\mathbf{g})=\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\log\left(\frac{p_{\mathbf{g}^{*}}(\mathbf{y})}{p_{\mathbf{g}}(\mathbf{y})}\right)p_{\mathbf{g}^{*}}(\mathbf{y}).$
A1 holds by assumption; A3 holds since $({\bm{\epsilon}}_{n})$ is a martingale difference, and A4 is obvious. We now prove A2. The set $J$ is closed and $\mathscr{l}$ is continuously differentiable. Moreover, for $\mathbf{g}\in\mathring{(\Delta^{D})}$,
$\mathbf{h}(\mathbf{g})^{T}\nabla_{\mathbf{g}}\mathscr{l}$ $=-\left(\mathbf{g}\circ\left(\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{k_{\mathbf{y}}}{p_{\mathbf{g}}(\mathbf{y})}p_{\mathbf{g}^{*}}(\mathbf{y})-\bf 1\right)\right)^{T}\left(\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{k_{\mathbf{y}}-k_{\mathbf{y},D}\mathbf{1}}{p_{\mathbf{g}}(\mathbf{y})}p_{\mathbf{g}^{*}}(\mathbf{y})\right)$
$=-\left(\sum_{i=1}^{D}\mathbf{g}_{i}\left(\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}k_{\mathbf{y}}({\bm{\vartheta}}_{i})\frac{p_{\mathbf{g}^{*}}(\mathbf{y})}{p_{\mathbf{g}}(\mathbf{y})}-1\right)^{2}\right)$
$\leq-\left(\sum_{i=1}^{D}\mathbf{g}_{i}\left(\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}k_{\mathbf{y}}({\bm{\vartheta}}_{i})\frac{p_{\mathbf{g}^{*}}(\mathbf{y})}{p_{\mathbf{g}}(\mathbf{y})}-1\right)\right)^{2}=0,$
where we have used Jensen inequality. We now prove that the inequality is strict, unless $\mathbf{g}=\mathbf{g}^{*}$. By contradiction, assume that there exist $\mathbf{g}^{\prime}$ and $\mathbf{g}^{\prime\prime}$ in $\mathring{(\Delta^{D})}$ with $\mathbf{g}^{\prime}\neq\mathbf{g}^{\prime\prime}$ such that, when $\mathbf{g}=\mathbf{g}^{\prime}$ and when $\mathbf{g}=\mathbf{g}^{\prime\prime}$, $\sum_{i=1}^{D}\mathbf{g}_{i}(\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}k_{\mathbf{y}}({\bm{\vartheta}}_{i})\frac{p_{\mathbf{g}^{*}}(\mathbf{y})}{p_{\mathbf{g}}(\mathbf{y})}-1)^{2}=0$, thus $\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}k_{\mathbf{y}}({\bm{\vartheta}}_{i})\frac{p_{\mathbf{g}^{*}}(\mathbf{y})}{p_{\mathbf{g}}(\mathbf{y})}=1$ for every $i=1,\dots,D$. Then, for every $i=1,\dots,D$ for every $a\in(0,1)$
$1=\sum_{i=1}^{D}(a\mathbf{g}^{\prime}({\bm{\vartheta}}_{i})+(1-a)\mathbf{g}^{\prime\prime}({\bm{\vartheta}}_{i}))\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}k_{\mathbf{y}}({\bm{\vartheta}}_{i})\frac{p_{\mathbf{g}^{*}}(\mathbf{y})}{ap_{\mathbf{g}^{\prime}}(\mathbf{y})+(1-a)p_{\mathbf{g}^{\prime\prime}}(\mathbf{y})}$
$<\sum_{i=1}^{D}(a\mathbf{g}^{\prime}({\bm{\vartheta}}_{i})+(1-a)\mathbf{g}^{\prime\prime}({\bm{\vartheta}}_{i}))\left(a\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}k_{\mathbf{y}}({\bm{\vartheta}}_{i})\frac{p_{\mathbf{g}^{*}}(\mathbf{y})}{p_{\mathbf{g}^{\prime}}(\mathbf{y})}+(1-a)\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}k_{\mathbf{y}}({\bm{\vartheta}}_{i})\frac{p_{\mathbf{g}^{*}}(\mathbf{y})}{p_{\mathbf{g}^{\prime\prime}}(\mathbf{y})}\right)$
$=1,$
which leads to a contradiction. Hence A2 holds, and ${\rm d}(\mathbf{g}_{n},J)\rightarrow 0$ $\mathds{P}^{*}$-a.s., as $n\rightarrow\infty$.
To conclude, we need to show that $\mathds{P}^{*}(\liminf_{n}{\rm d}(\mathbf{g}_{n},\partial\Delta^{D})=0)=0$. Since ${\rm d}(\mathbf{g}^{*},\partial\Delta^{D})>0$ and ${\rm d}(\mathbf{g}_{n},\{\mathbf{g}^{*}\}\cup\partial\Delta^{D})\rightarrow 0$ $\mathds{P}-a.s.$, then $\mathds{P}^{*}(\liminf_{n}{\rm d}(\mathbf{g}_{n},\partial\Delta^{D})=0)>0$ if and only if there exists $M$ with $\mathds{P}^{*}(M)>0$ such that ${\rm d}(\mathbf{g}_{n}(\omega),\partial\Delta^{D})\rightarrow 0$ for every $\omega\in M$, as $n\rightarrow\infty$. Let us make this assumption as a contradiction hypothesis. Define the function $\mathcal{L}(\mathbf{g})=\sum_{i=1}^{D}\log(\mathbf{g}^{*}_{i}/\mathbf{g}_{i})\mathbf{g}^{*}_{i}$. By the contradiction hypothesis, $\mathcal{L}(\mathbf{g}_{n})(\omega)\rightarrow\infty$ as $n\rightarrow\infty$, for every $\omega\in M$. Then $E^{*}({\mathcal{L}}(\mathbf{g}_{n}))\rightarrow\infty$ as $n\rightarrow+\infty$, where $E^{*}$ denote expectation with respect to $\mathds{P}^{*}$. Let $\tilde{\mathbf{g}}_{n}=\eta_{n}\mathbf{g}_{n}+(1-\eta_{n})\mathbf{g}_{n+1}$, with the random variable $\eta_{n}$ taking values in $[0,1]$, be such that
${\mathcal{L}}(\mathbf{g}_{n+1})$ $=\mathcal{L}(\mathbf{g}_{n})+\nabla_{\mathbf{g}}\mathcal{L}(\mathbf{g}_{n})^{T}(\mathbf{g}_{n+1}-\mathbf{g}_{n})$
$+(\mathbf{g}_{n+1}-\mathbf{g}_{n})^{T}\nabla^{2}_{\mathbf{g}}\mathcal{L}(\tilde{\mathbf{g}}_{n})(\mathbf{g}_{n+1}-\mathbf{g}_{n}).$
Then,
${\mathcal{L}}(\mathbf{g}_{n+1})$ $=\mathcal{L}(\mathbf{g}_{n})-\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})+\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{(\tilde{\mathbf{g}}_{n,i})^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}.$
Notice that
$E^{*}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})\right)=\alpha_{n+1}E^{*}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}(\mathbf{h}_{i}(\mathbf{g}_{n})+{\bm{\epsilon}}_{n+1,i})\right),$
with
$E^{*}\left(\sum_{i=1}^{D}\frac{\mathbf{g}_{i}}{\mathbf{g}_{n,i}}{\bm{\epsilon}}_{n+1,i}\right)=0,$
and
$\liminf_{n}E^{*}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}\mathbf{h}_{i}(\mathbf{g}_{n})\right)=\liminf_{n}E^{*}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}\left(\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{k_{\mathbf{y}}({\bm{\vartheta}}_{i})\mathbf{g}_{n,i}p_{\mathbf{g}^{*}}(\mathbf{y})}{p_{\mathbf{g}_{n}}(\mathbf{y})}-\mathbf{g}_{n,i}\right)\right)$
$=\liminf_{n}E^{*}\left(\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{p_{\mathbf{g}^{*}}(\mathbf{y})^{2}}{p_{\mathbf{g}_{n}}(\mathbf{y})}-1\right)$
$>\liminf_{n}E^{*}\left(\left(\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{p_{\mathbf{g}_{n}}(\mathbf{y})}{p_{\mathbf{g}^{*}}(\mathbf{y})}p_{\mathbf{g}^{*}}(\mathbf{y})\right)^{-1}-1\right)=0,$
where the strict inequality holds since, under the contradiction assumption, $\mathds{P}^{*}(p_{\mathbf{g}_{n}}\not\rightarrow p_{\mathbf{g}^{*}})>0$. Thus there exists a constant $c>0$ such that
$E^{*}\left(-\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})\right)<-c\alpha_{n+1}$
for $n$ large enough. On the other hand, denoting by $E^{*}_{n}$ the conditional expectation given $\mathcal{G}_{n}$, we can write that
$E^{*}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{(\tilde{\mathbf{g}}_{n,i})^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}\right)$
$\leq E^{*}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}\right)+E^{*}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n+1,i}^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}\right)$
$\leq\alpha_{n+1}^{2}E^{*}\left(\sum_{i=1}^{D}\mathbf{g}^{*}_{i}\left(\frac{k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})}{p_{\mathbf{g}_{n}}(\mathbf{Y}_{n+1})}-1\right)^{2}\right)+\alpha_{n+1}^{2}E^{*}\left(\sum_{i=1}^{D}\mathbf{g}^{*}_{i}\frac{\left(\frac{k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})}{p_{\mathbf{g}_{n}}(\mathbf{Y}_{n+1})}-1\right)^{2}}{\left(1+\alpha_{n+1}\left(\frac{k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})}{p_{\mathbf{g}_{n}}(\mathbf{Y}_{n+1})}-1\right)\right)^{2}}\right)$
$\leq\alpha_{n+1}^{2}E^{*}\left(1+E^{*}_{n}\left(\sum_{i=1}^{D}\left(\frac{k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})}{\sum_{i=1}^{D}k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})\mathbf{g}_{n,i}}\right)^{2}\right)\right)$
$\quad\quad+\frac{\alpha_{n+1}^{2}}{(1-\alpha_{n+1})^{2}}E^{*}\left(1+E^{*}_{n}\left(\sum_{i=1}^{D}\left(\frac{k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})}{\sum_{i=1}^{D}k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})\mathbf{g}_{n,i}}\right)^{2}\right)\right)$
$\leq 2\frac{\alpha_{n+1}^{2}}{(1-\alpha_{n+1})^{2}}\left(1+\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\sum_{i,{i^{\prime}}=1}^{D}\frac{k_{\mathbf{y}}({\bm{\vartheta}}_{i})^{2}}{(\sum_{{i^{\prime\prime}}}k_{\mathbf{y}}({\bm{\vartheta}}_{i^{\prime\prime}})\mathbf{g}_{n}({\bm{\vartheta}}_{i^{\prime\prime}}))^{2}}k_{\mathbf{y}}({\bm{\vartheta}}_{i^{\prime}})\mathbf{g}^{*}({\bm{\vartheta}}_{i^{\prime}})\right)$
$\leq 2\frac{\alpha_{n+1}^{2}}{(1-\alpha_{n+1})^{2}}\left(1+\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\sum_{i,{i^{\prime}},{i^{\prime\prime}}=1}^{D}\frac{k_{\mathbf{y}}({\bm{\vartheta}}_{i})^{2}}{k_{\mathbf{y}}({\bm{\vartheta}}_{i^{\prime\prime}})^{2}}k_{\mathbf{y}}({\bm{\vartheta}}_{i^{\prime}})\mathbf{g}^{*}({\bm{\vartheta}}_{i^{\prime}})\mathbf{g}_{n}({\bm{\vartheta}}_{i^{\prime\prime}})\right)$
$\leq 2\frac{\alpha_{n+1}^{2}}{(1-\alpha_{n+1})^{2}}\left(1+D\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\max_{i,{i^{\prime}},{i^{\prime\prime}}}\frac{k_{\mathbf{y}}({\bm{\vartheta}}_{i})^{2}}{k_{\mathbf{y}}({\bm{\vartheta}}_{i^{\prime\prime}})^{2}}k_{\mathbf{y}}({\bm{\vartheta}}_{i^{\prime}})\right)$
$=o(\alpha_{n}),$
where the last equality holds since for every $i,{i^{\prime}},{i^{\prime\prime}}$
$\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{k_{\mathbf{y}}({\bm{\vartheta}}_{i})^{2}}{k_{\mathbf{y}}({\bm{\vartheta}}_{i^{\prime}})^{2}}k_{\mathbf{y}}({\bm{\vartheta}}_{i^{\prime\prime}})$ $=\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{\prod_{j=1}^{k}e^{-2{\bm{\vartheta}}_{i,j}}\vartheta_{i,j}^{2\mathbf{y}_{j}}/(\mathbf{y}_{j}!)^{2}}{\prod_{j=1}^{k}e^{-2{\bm{\vartheta}}_{i^{\prime},j}}\vartheta_{i^{\prime},j}^{2\mathbf{y}_{j}}/(\mathbf{y}_{j}!)^{2}}\prod_{j=1}^{k}e^{-{\bm{\vartheta}}_{i^{\prime\prime},j}}\vartheta_{i^{\prime\prime},j}^{\mathbf{y}_{j}}/\mathbf{y}_{j}!$
$=\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\prod_{j=1}^{k}e^{2{\bm{\vartheta}}_{i^{\prime},j}-2{\bm{\vartheta}}_{i,j}-{\bm{\vartheta}}_{i^{\prime\prime},j}}\left(\frac{{\bm{\vartheta}}_{i,j}^{2}{\bm{\vartheta}}_{i^{\prime\prime},j}}{{\bm{\vartheta}}_{i^{\prime},j}^{2}}\right)^{y}\frac{1}{y_{j}!}$
$=\prod_{j=1}^{k}\sum_{\mathbf{y}_{j}\in\mathbb{N}_{0}}e^{2{\bm{\vartheta}}_{i^{\prime},j}-2{\bm{\vartheta}}_{i,j}-{\bm{\vartheta}}_{i^{\prime\prime},j}}\left(\frac{{\bm{\vartheta}}_{i,j}^{2}{\bm{\vartheta}}_{i^{\prime\prime},j}}{{\bm{\vartheta}}_{i^{\prime},j}^{2}}\right)^{y}\frac{1}{y_{j}!}<\infty.$
It follows that,
$E^{*}\left(\mathcal{L}(\mathbf{g}_{n+1})\right)\leq E^{*}\left(\mathcal{L}(\mathbf{g}_{n})\right),$
for $n$ large enough, which contradicts $E^{*}(\mathcal{L}(\mathbf{g}_{n}))\rightarrow\infty$. ∎
Lemma F.5.
Under the assumptions of Theorem 2.4, for every $\delta<2-1/\gamma$,
$||\mathbf{g}_{n}-\mathbf{g}^{*}||^{2}=o(n^{-\delta})\quad\mathds{P}^{*}-a.s.$
Proof.
In this proof, $\mathbf{g}_{n}$ and $\mathbf{g}^{*}$ are considered as random vectors in $\{\mathbf{v}\in\mathbb{R}^{D}:||\mathbf{v}||\leq 1\}$. By Lemma F.4, $\mathbf{g}_{n}$, seen as vectors in $\Delta^{D}$ converge $P^{*}$-a.s. to $\mathbf{g}\in\Delta^{D}$. This implies that $\mathbf{g}_{n}$ converges to $\mathbf{g}^{*}$ also when seen as vectors in $\mathbb{R}^{D}$. Moreover, $(\mathbf{g}_{n})$ satisfies the stochastic approximation in $\mathbb{R}^{D}$:
$\mathbf{g}_{n+1}$ $=\mathbf{g}_{n}+\alpha_{n+1}\mathbf{h}(\mathbf{g}_{n})+\alpha_{n+1}{\bm{\epsilon}}_{n+1},$
with initial value $\mathbf{g}_{0}$, where
$\mathbf{h}(\mathbf{g})=\mathbf{g}\circ\left(\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{k_{\mathbf{y}}}{p_{\mathbf{g}}(\mathbf{y})}p_{\mathbf{g}^{*}}(\mathbf{y})-\bf 1\right),$
and
${\bm{\epsilon}}_{n+1}=\mathbf{g}\circ\left(\frac{k_{\mathbf{Y}_{n+1}}}{p_{\mathbf{g}}(\mathbf{Y}_{n+1})}p_{\mathbf{g}^{*}}(\mathbf{Y}_{n+1})-\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{k_{\mathbf{y}}}{p_{\mathbf{g}}(\mathbf{y})}p_{\mathbf{g}^{*}}(\mathbf{y})\right).$
The thesis follows from Theorem 3.1.1 in Chen (2005) if we can prove that
  • B1
    $\alpha_{n}\rightarrow 0$, $\sum_{n=1}^{\infty}\alpha_{n}=\infty$, $\alpha_{n+1}^{-1}-\alpha_{n}^{-1}\rightarrow 0$,
  • B2
    $\sum_{n=1}^{\infty}\alpha_{n}^{1-\delta/2}{\bm{\epsilon}}_{n}$ converges, $\mathds{P}^{*}$-a.s.
  • B3
    $\mathbf{h}$ is measurable and locally bounded, and is differentiable at $\mathbf{g}^{*}$. Let $H$ be a matrix such that, as $\mathbf{g}\rightarrow\mathbf{g}^{*}$,
    $\mathbf{h}(\mathbf{g})=\mathbf{H}(\mathbf{g}-\mathbf{g}^{*})+\mathbf{r}(\mathbf{g}),\quad\mathbf{r}(g^{*})=0,\quad\mathbf{r}(g)=o(||\mathbf{g}-\mathbf{g}^{*}||).$
    All the eigenvalues of the matrix $H$ have negative real parts.
Remark F.6.
Theorem 3.1.1 in Chen (2005) also requests the following condition: B4 There exists a is continuously differentiable function $\mathscr{l}:\{\mathbf{v}\in\mathbb{R}^{D}:||\mathbf{v}||\leq 1\}\rightarrow\mathbb{R}$ such that $\sup_{\delta_{1}\leq{\rm d}(\mathbf{g},\mathbf{g}^{*})\leq\delta_{2}}\mathbf{h}(\mathbf{g})^{T}\nabla_{\mathbf{g}}\mathscr{l}<0$ for every $\delta_{2}>\delta_{1}>0$. However, a close inspection of the proof of Theorem 3.1.1 in Chen (2005) shows that B4 is employed only to ensure convergence of $\mathbf{g}_{n}$ towards $\mathbf{g}^{*}$. In our case, this property has been proved in Lemma F.4.
B1 is obvious. To prove B2, notice that $(\sum_{k=1}^{n}\alpha_{k}^{1-\delta/2}{\bm{\epsilon}}_{k})_{n\geq 1}$ is a martingale. Moreover, there exoists a constant $C$ such that, for every $i=1,\dots,D$,
$\sup_{n}E^{*}((\sum_{k=1}^{n}\alpha_{k}^{1-\delta/2}{\bm{\epsilon}}_{k,i})^{2})$ $=\sup_{n}\sum_{k=1}^{n}\alpha_{k}^{2-\delta}E^{*}({\bm{\epsilon}}_{k,i}^{2})$
$\leq C\sum_{k=1}^{\infty}\left(\frac{\alpha}{\alpha+k}\right)^{\gamma(2-\delta)}<\infty,$
since $\gamma(2-\delta)>1$ and $({\bm{\epsilon}}_{n,i})$ are uniformly bounded. We now prove B3. The function $\mathbf{h}$ is continuously differentiable at $\mathbf{g}^{*}$ and
$\mathbf{H}_{i,j}=-\mathbf{g}^{*}_{i}\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{k_{\mathbf{y}}({\bm{\vartheta}}_{i})k_{\mathbf{y}}({\bm{\vartheta}}_{j})}{p_{\mathbf{g}^{*}}(\mathbf{y})}.$
This matrix can be seen as the product of the diagonal matrix ${\rm diag}(\mathbf{g}^{*})$, which is positive definite, times $-\mathbf{M}$, with $\mathbf{M}_{i,j}=\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{k_{\mathbf{y}}({\bm{\vartheta}}_{i})k_{\mathbf{y}}({\bm{\vartheta}}_{j})}{p_{\mathbf{g}^{*}}(\mathbf{y})}$. We prove that $M$ is positive definite. Given a vector $\mathbf{u}\in\mathbb{R}^{D}$ with $||\mathbf{u}||=1$, we can write that
$\mathbf{u}^{T}\mathbf{M}\mathbf{u}=$ $\sum_{i,j=1}^{D}\mathbf{u}_{i}\mathbf{u}_{j}\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{k_{\mathbf{y}}({\bm{\vartheta}}_{i})k_{\mathbf{y}}({\bm{\vartheta}}_{j})}{p_{\mathbf{g}^{*}}(\mathbf{y})}=\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{1}{p_{\mathbf{g}^{*}}(\mathbf{y})}\left(\sum_{i=1}^{D}\mathbf{u}_{i}k_{\mathbf{y}}({\bm{\vartheta}}_{i})\right)^{2}\geq 0,$
with equality if and only if $\sum_{i=1}^{D}\mathbf{u}_{i}k_{\mathbf{y}}({\bm{\vartheta}}_{i})=0$ for every $\mathbf{y}\in\mathbb{N}_{0}^{k}$. The only solution to the last equation in $\mathbf{u}$ is the zero vector, which is inconsistent with $||\mathbf{u}||=1$. Since the function $\mathbf{u}^{T}\mathbf{M}\mathbf{u}$ is continuous on the compact set $||\mathbf{u}||=1$, then the minimum is attained, and by the above reasoning it is strictly positive. It follows that $\mathbf{M}$ is positive definite. ∎
of Theorem E.2.
Denoting by $(\mathbf{g}_{n})$ and $\mathbf{g}^{*}$ the representations of $(g_{n})$ and $g^{*}$ in $\mathbb{R}^{D}$, we can write, by Lemma F.3 and Lemma F.5,
${\rm Regret}(g_{n},g^{*})\leq C||\mathbf{g}_{n}-\mathbf{g}^{*}||^{2}=o(n^{-\delta}).$

### v2

@@
 Abstract
-The Poisson compound decision problem is a classical problem in statistics, for which parametric and nonparametric empirical Bayes methodologies are available to estimate the Poisson’s means in static or batch domains.
-In this paper, we consider the Poisson compound decision problem in a streaming or online domain.
-By relying on a quasi-Bayesian approach, often referred to as Newton’s algorithm, we obtain sequential Poisson’s mean estimates that are of easy evaluation, computationally efficient and with a constant computational cost as data increase, which is desirable for streaming data.
-Large sample asymptotic properties of the proposed estimates are investigated, also providing frequentist guarantees in terms of a regret analysis.
-We validate empirically our methodology, both on synthetic and real data, comparing against the most popular alternatives.
-Keywords: Empirical Bayes; Newton’s algorithm; Poisson compound decision; quasi-Bayes; streaming data.
+The Poisson compound decision problem is a long-standing problem in statistics, where empirical Bayes methodologies are commonly used to estimate Poisson’s means in static or batch domains.
+In this paper, we study the Poisson compound decision problem in a streaming or online domain.
+Adopting a quasi-Bayesian approach, referred to as Newton’s algorithm, we obtain a sequential estimate that is easy to evaluate, computationally efficient, and maintain a constant per-observation computational cost as data accumulate.
+Asymptotic frequentist guarantees of this estimate are established, showing consistency and asymptotic optimality, where the latter is understood as vanishing excess Bayes risk or regret.
+We demonstrate the effectiveness of our methodology through empirical analysis on synthetic and real data, with comparisons to existing approaches.
+Keywords: Empirical Bayes; Newton’s algorithm; Poisson compound decision; quasi-Bayes; streaming data
 1 Introduction
 Given $n\geq 1$ observations modeled as independent Poisson random variables $Y_{1},\ldots,Y_{n}$ with means $\theta_{1},\ldots,\theta_{n}$, respectively, the Poisson compound decision problem refers to the estimation of $(\theta_{1},\ldots,\theta_{n})$ under the squared-error loss.
-Empirical Bayes is a general approach to handle compound decision problems as first suggested by Robbins (1951, 1956); see also Efron (2003) and Zhang (2003).
-Assuming the parameters $\theta_{i}$’s to be i.i.d as a prior $G$ on $\Theta\subseteq\mathbb{R}^{+}$, the best estimate of $\theta_{i}$ under the squared-error loss, i.e.
-the Bayes estimate, is the posterior mean, namely for $y\in\mathbb{N}_{0}$
+Empirical Bayes is a general approach to handle compound decision problems (Robbins, 1951, 1956; Zhang, 2003).
+Assuming the $\theta_{i}$’s to be i.i.d.
+as a prior $G$ on $\Theta\subseteq\mathbb{R}^{+}$, the best estimate of $\theta_{i}$ is the posterior mean or Bayes estimate, namely for $y\in\mathbb{N}_{0}$
 $\hat{\theta}_{G}(y)=E_{G}[\theta_{i}\mid Y_{i}=y]=\frac{\int_{\Theta}\theta\frac{\text{e}^{-\theta}\theta^{y}}{y!}G(\mathrm{d}\theta)}{\int_{\Theta}\frac{\text{e}^{-\theta}\theta^{y}}{y!}G(\mathrm{d}\theta)}=(y+1)\frac{p_{G}(y+1)}{p_{G}(y)},$ (1)
 where $p_{G}$ is the probability mass function of the $Y_{i}$’s.
-The empirical Bayes approach proposes to bypass the knowledge of $G$, which might be unavailable in practice, estimating either $p_{G}$ or ${G}$.
-Methodologies to estimate $p_{G}$ and ${G}$ are referred to as $f$-modeling and $g$-modeling, respectively (Efron, 2014).
-Under $f$-modeling, one obtains an estimate of $p_{G}$ from $(Y_{1},\ldots,Y_{n})$, and then apply (1).
-For example, Robbins (1956) proposed an estimate of $\hat{\theta}_{G}(y)$ by replacing $p_{G}$ in (1) with the empirical distribution
-$\hat{p}_{G}(y)=\sum_{i=1}^{n}I(Y_{i}=y)\qquad y\in\mathbb{N}_{0},$ (2)
-where $I$ denotes the indicator function.
-This is the celebrated nonparametric empirical Bayes estimate (Brown et al., 2013; Polyanskiy and Wu, 2021).
-Despite its simplicity and (minimax) optimality, Robbins’ method produces unstable estimates in practice; this occurs particularly when $\hat{p}_{G}$ is small or zero (Efron and Hastie, 2021).
-To alleviate the shortcomings of Robbins’ method, a number of modifications that enforce smoothness of $\hat{p}_{G}(y)$ have been proposed, including $g$-modeling.
-Under $g$-modeling, one obtains an estimate of $G$ from $(Y_{1},\ldots,Y_{n})$, and then apply (1).
-An example is the estimate of $\hat{\theta}_{G}(y)$ obtained by replacing $G$ in (1) with the nonparametric maximum likelihood estimate
-$\hat{G}=\operatorname*{arg\,max}_{G}\frac{1}{n}\sum_{i=1}^{n}\log p_{G}(Y_{i}),$ (3)
-where the maximization is taken over all distributions on $\mathbb{R}^{+}$ (Kiefer and Wolfowitz, 1956).
-Much of $g$-modeling involves the estimate (3) (Laird, 1978; Jiang and Zhang, 2009) and nonparametric minimum distance estimates (Jana et al., 2024).
-Parametric $g$-modeling has been much less developed (Efron and Morris, 1978; Efron, 2016).
-A fully Bayesian approach is possible by placing a hyperprior distribution on $G$; this is the Bayes empirical Bayes method (Deely and Lindley, 1981).
-The Poisson compound decision problem provides a flexible framework for count data, with a broad range of applications where static or batch processing have traditionally been sufficient, e.g.
-missing-species problems and microarrays (Efron, 2003), insurance analytics (Efron and Hastie, 2021, Chapter 6), deconvolution (Efron, 2016), data confidentiality (Zhang, 2005) and sports analytics (Jana et al., 2024).
-However, the emergence of streaming count data in some of these applications, as well as in social media analytics, network traffic monitoring, bio-surveillance and healthcare, and e-commerce web analytics, makes it natural and necessary to consider the Poisson compound decision problem in a streaming or online domain, where a timely analysis can enhance decision-making and responsiveness to new information.
-While it is indeed possible to adapt existing empirical Bayes methodologies to streaming data, the resulting estimates typically suffer from computational inefficiencies, especially when they involve optimizations and posterior computations.
-1.1 Our contribution
-We propose a principled sequential $g$-modeling methodology for the Poisson compound decision problem: starting from an initial guess $G_{0}$ for the prior $G$, for $n\geq 1$ we update $G_{n-1}$ at each new observation $Y_{n}$ following a procedure of Smith and Makov (1978), often referred to as Newton’s algorithm (Newton et al., 1998; Martin and Ghosh, 2008).
-This recursive procedure is however regarded not just as an algorithm, but as a quasi-Bayesian learning model (Fortini and Petrone, 2020), which provides a predictive construction for an asymptotically equivalent Bayesian model.
-The resulting quasi-Bayes $g$-modeling allows to estimate $\hat{\theta}_{G}(y)$, based on $(Y_{1},\dots,Y_{n})$, with the plug-in estimate $\hat{\theta}_{G_{n}}(y)$ obtained from (1) by replacing $G$ with $G_{n}$.
-Besides its ease of evaluation, the sequential estimate $\hat{\theta}_{G_{n}}(y)$ is computationally efficient and with constant computational cost as data increase, which is desirable for streaming data.
-By considering Newton’s algorithm as a learning model, it is possible to measure uncertainty of $\hat{\theta}_{G_{n}}(y)$; here, we obtain a Gaussian central limit theorem for $\hat{\theta}_{G_{n}}(y)$, $n\rightarrow+\infty$, leading to asymptotic credible intervals for $\hat{\theta}_{G}(y)$.
-We provide the estimate $\hat{\theta}_{G_{n}}(y)$ with asymptotic frequentist guarantees.
-Assuming the “true” mixing distribution $G$ to be known, say equal to the oracle prior $G^{\ast}$, and that the observations $Y_{n}$’s are i.i.d.
-from a Poisson mixture model with mixing distribution $G^{*}$, we study the large $n$ asymptotic behavior of the Bayesian regret of $\hat{\theta}_{G_{n}}(y)$, which is defined as the difference between the Bayes risk of $\hat{\theta}_{G_{n}}(y)$ and the Bayes risk of the plug-in estimate $\hat{\theta}_{G^{\ast}}(y)$ obtained from (1) by replacing $G$ with $G^{\ast}$.
-In particular, we characterize the rate at which the Bayesian regret goes to zero as $n\rightarrow+\infty$.
-An empirical validation of our sequential $g$-modeling is presented on synthetic data generated i.i.d from a “true” mixing distribution $G$, and on real data.
-The estimate $\hat{\theta}_{G_{n}}(y)$, at fixed $n\geq 1$, outperforms the nonparametric empirical Bayes method and competes with the nonparametric maximum likelihood and the minimum squared Hellinger distance methods.
-Further, $\hat{\theta}_{G_{n}}(y)$ competes with the parametric Bayes empirical Bayes method, assuming a Gamma hyperprior on $G$.
+The empirical Bayes approach proposes to bypass the knowledge of $G$, which might be unavailable in practice, estimating either the marginal $p_{G}$, a strategy known as $f$-modeling, or the prior ${G}$, a strategy known as $g$-modeling (Efron, 2014).
+1.1 Background and motivation
+A prominent example of $f$-modeling is Robbins’ method (Robbins, 1956), which estimates $p_{G}$ in (1) with the empirical distribution of the $Y_{i}$’s (Brown et al., 2013; Polyanskiy and Wu, 2021; Jana et al., 2023); see Efron (2014, Section 3 and Section 4).
+Although widely used in practice for its computational simplicity, $f$-modeling lacks robustness and numerical stability; in fact, it is easily derailed by “outliers” in the data, namely data points that appear only a few times, causing estimators to take exceptionally small or large values (Efron and Hastie, 2021; Shen and Wu, 2024).
+Examples of $g$-modeling include estimates of $G$ in (1) based on maximum likelihood and minimum distance methods (Jiang and Zhang, 2009; Jana et al., 2024); see Efron (2014, Section 5).
+A fully Bayesian approach is also possible by placing a hyperprior on $G$, a strategy known as Bayes empirical Bayes (Deely and Lindley, 1981; Efron, 2019).
+While $g$-modeling yields more accurate estimates than $f$-modeling, as well as better incorporating prior knowledge, it is computationally demanding, particularly in nonparametric and high-dimensional settings (Jana et al., 2023).
+The Poisson compound decision problem offers a flexible framework for modeling count data, with successful applications in settings where static or batch processing has traditionally sufficed, such as microarrays (Efron, 2003), insurance analytics (Efron and Hastie, 2021, Chapter 6), deconvolution (Efron, 2016), data confidentiality (Zhang, 2005) and sports analytics (Jana et al., 2024).
+However, the rise of streaming data in some of these fields, as well as in social media analytics, network traffic monitoring, bio-surveillance and healthcare, and e-commerce web analytics, calls for extending the Poisson compound decision problem to the online setting, where timely analysis is critical for responsive and informed decision-making.
+Although the existing empirical Bayes methods can, in principle, be adapted to streaming data, they often suffer from computational inefficiencies, particularly when involving optimization or posterior computations.
+1.2 Preview of our contributions
+We propose a principled sequential $g$-modeling methodology for the Poisson compound decision problem.
+Starting from an initial guess $G_{0}$ for $G$, for $n\geq 1$ our approach updates $G_{n-1}$ at each new observation $Y_{n}$ following a procedure of Smith and Makov (1978), referred to as Newton’s algorithm (Newton et al., 1998; Martin and Ghosh, 2008).
+This recursive procedure is regarded as a quasi-Bayesian learning model, yielding a predictive construction for a large-$n$ asymptotically equivalent Bayesian model (Fortini and Petrone, 2020).
+A quasi-Bayes empirical Bayes estimate of $\hat{\theta}_{G}(y)$ is then obtained from (1) by replacing $G$ with $G_{n}$.
+Besides being easy to evaluate, the estimate $\hat{\theta}_{G_{n}}(y)$ is computationally efficient, and maintain a constant per-observation computational cost as data accumulate.
+Considering Newton’s algorithm as a learning model allows to measure uncertainty for $\hat{\theta}_{G_{n}}(y)$, for which we build an asymptotic credible interval via the central limit theorem.
+Asymptotic frequentist guarantees are established for $\hat{\theta}_{G_{n}}(y)$, assuming that the observations $Y_{n}$’s are i.i.d.
+from a Poisson mixture model with a “true” mixing distribution, say the oracle prior $G^{\ast}$.
+We provide sufficient conditions on $G^{\ast}$ to ensure consistency and asymptotic optimality of $\hat{\theta}_{G_{n}}(y)$ as $n\rightarrow+\infty$ , where the latter is understood as vanishing excess Bayes risk or regret.
+Under additional assumptions on Newton’s algorithm, the rate at which the regret converges to zero is also derived.
+We validate our methodology empirically on synthetic and real data.
+For fixed $n\geq 1$, the estimate $\hat{\theta}_{G_{n}}(y)$ consistently outperforms Robbins’ method and performs competitively with the nonparametric maximum likelihood method and the minimum squared Hellinger distance method.
+It also performs on par with a parametric Bayes empirical Bayes method with a Gamma hyperprior.
 2 Quasi-Bayes empirical Bayes
-We consider a stream of observations $(Y_{n})_{n\geq 1}$, such that the $Y_{n}$’s have Poisson distributions with means $\theta_{n}$’s, respectively.
-To estimate sequentially the parameters $\theta_{n}$’s, we adopt a $g$-modeling methodology that, on a probability space $(\Omega,\mathcal{F},\mathds{P})$, specifies sequentially the distribution of $(Y_{n})_{n\geq 1}$ by giving the conditional distribution of $Y_{n+1}$ given $(Y_{1},\dots,Y_{n})$, for each $n\geq 0$.
-In particular, denoting by $(\mathcal{G}_{n})_{n\geq 1}$ the natural filtration of $(Y_{n})_{n\geq 1}$, for $n\geq 0$ each conditional distribution with respect to $\mathcal{G}_{n}$ is assumed to be a mixture of Poisson kernels $k(\cdot\mid\theta)$, with $\theta$ defined in a sufficiently fine and wide grid of points $\Theta_{d}=\{\vartheta_{1},\dots,\vartheta_{d}\}$, with $\vartheta_{j}<\vartheta_{j+1}$ ($j=1,\dots,d-1$), namely
+For a stream of observations $(Y_{n})_{n\geq 1}$, we propose a sequential $g$-modeling methodology that, on a probability space $(\Omega,\mathcal{F},\mathds{P})$, specifies the distribution of $(Y_{n})_{n\geq 1}$ through the conditional distribution of $Y_{n+1}$ given $(Y_{1},\dots,Y_{n})$, for each $n\geq 0$.
+As in the Bayesian framework, the model is conceived here as a learning model and not as a data-generating process (Fortini and Petrone, 2025).
+Letting $(\mathcal{G}_{n})_{n\geq 1}$ denote the natural filtration of $(Y_{n})_{n\geq 1}$, for each $n\geq 0$ the conditional distribution of $Y_{n+1}$ given $\mathcal{G}_{n}$ is modeled according to a mixture of the Poisson kernel $k(\cdot\mid\theta)$, with $\theta$ ranging over a sufficiently fine and wide grid $\Theta_{d}=\{\vartheta_{1},\dots,\vartheta_{d}\}$, with $\vartheta_{j}<\vartheta_{j+1}$ for $j=1,\dots,d-1$, i.e.,
 $p_{g_{n}}(y)=\mathds{P}(Y_{n+1}=y\mid\mathcal{G}_{n})=\sum_{j=1}^{d}k(y\mid\vartheta_{j})g_{n}(\vartheta_{j})\quad y\in\mathbb{N}_{0},$ (2)
@@
 $g_{n+1}(\vartheta_{j})=(1-\alpha_{n+1})g_{n}(\vartheta_{j})+\alpha_{n+1}\frac{k(Y_{n+1}\mid\vartheta_{j})g_{n}(\vartheta_{j})}{\sum_{i=1}^{d}k(Y_{n+1}\mid\vartheta_{i})g_{n}(\vartheta_{i})}\qquad j=1,\ldots,d,$ (3)
-given an initial guess $g_{0}$ and a sequence $(\alpha_{n})_{n\geq 1}$ of real numbers in $(0,1)$ such that $\sum_{n\geq 1}\alpha_{n}=+\infty$ and $\sum_{n\geq 1}\alpha^{2}_{n}<+\infty$.
-From (5), $g_{n}$ is updated as $Y_{n+1}$ becomes available, by taking a weighted average, with respect to $\alpha_{n}$, of $g_{n}$ itself and of its posterior distribution based on $Y_{n+1}$.
+given an initial guess $g_{0}$ and the $\alpha_{n}$’s in $(0,1)$ such that $\sum_{n\geq 1}\alpha_{n}=+\infty$ and $\sum_{n\geq 1}\alpha^{2}_{n}<+\infty$.
+From (3), the learning model updates $g_{n}$ as $Y_{n+1}$ becomes available, by taking a weighted average, with respect to $\alpha_{n}$, between $g_{n}$ and its posterior distribution based on $Y_{n+1}$.
 The sequence $(\alpha_{n})_{n\geq 1}$ is referred to as the learning rate, for which a standard choice is $\alpha_{n}=(\alpha+n)^{-\gamma}$ for $\alpha>0$ and $\gamma\in(1/2,1]$.
-The next result explains in which sense our sequential $g$-modeling is quasi-Bayesian.
+The following proposition outlines an effective procedure for constructing the grid $\Theta_{d}$ used in (2)-(3).
+Beyond serving as a practical tool for implementing the sequential $g$-modeling methodology, this construction also plays a central role in establishing asymptotic frequentist guarantees.
+We denote by $\text{KL}(p||q)$ the Kullback-Leibler divergence between the probability distributions $p$ and $q$.
+Proposition 2.1.
+Let the prior $G$ in (1) be such that $\sum_{y\in\mathbb{N}_{0}}y^{k}p_{G}(y)\leq m_{k}$ for some $k\in\mathbb{N}$ with $k\geq 2$ and $m_{k}\in\mathbb{R}$, and for a fixed $\eta>0$ define the equally spaced grid $\Theta_{\eta}=\{\vartheta_{i}=i\eta\text{ : }i=1,\ldots,d_{\eta}\}$ with $d_{\eta}=\inf\{n\in\mathbb{N}\text{ : }n>\eta^{-1}\text{ and }n^{1-k}\log(n\eta)m_{k}\leq\eta^{k}\}$.
+If $\mathcal{S}_{\eta}$ is the set of all the probability mass functions on $\Theta_{\eta}$, then $g_{\eta}=\text{argmin}_{f\in\mathcal{S}_{\eta}}\text{KL}(p_{G}||p_{f})$ is unique and $\text{KL}(p_{G}||p_{g_{\eta}})<2\eta$.
+See C.1 for the proof of Proposition 2.1.
+The probability mass function $g_{\eta}$ in Proposition 2.1 is referred to as the $d_{\eta}$-KL-discretization of a prior $G$ on $\Theta\subseteq\mathbb{R}^{+}$.
+This form of discretization is of independent interest in $g$-modeling approaches to the Poisson compound decision problem (Jana et al., 2024).
+2.1 Quasi-Bayes properties
+We show that learning model defined by (2)-(3) is quasi-Bayesian in the sense of Fortini and Petrone (2020), yielding a predictive construction for a large-$n$ asymptotically equivalent Bayesian model.
 Theorem 2.2.
 […]
-See Supplementary Material C.1 for the proof of Theorem 2.1.
-Under (4)-(5), for each $n\geq 0$, the future observations of $(Y_{n})_{n\geq 1}$ are conditionally identically distributed given $\mathcal{G}_{n}$.
+See C.2 for the proof of Theorem 2.2.
+Under the learning model (2)-(3), for each $n\geq 0$, the future observations of $(Y_{n})_{n\geq 1}$ are considered conditionally identically distributed given $\mathcal{G}_{n}$.
 From Berti et al.
-(2004), this implies that the $Y_{n}$’s are asymptotically i.i.d., given $\tilde{g}$, with law $p_{\tilde{g}}$, taking on the interpretation of a quasi-Bayesian framework (Fortini and Petrone, 2020).
-In particular, $p_{\tilde{g}}$ is the random distribution that would be assigned if we had access to an infinitely large sample, from which it is reasonable to consider $\tilde{g}$ as the parameter driving our sequential $g$-modeling.
-Differently from the quasi-Bayesian framework, the Bayesian framework relies on a prior distribution on $g$, such that the $Y_{n}$’s are i.i.d, given $g$.
-Under this framework, the computation cost to evaluate the posterior distribution of $g$ is typically cumbersome, and the posterior distribution must be recomputed for each new observation.
-The quasi-Bayesian framework offers a significantly lower computational cost.
-2.1 Sequential estimation
-Under the sequential $g$-modeling methodology of Section 2, the empirical Bayes estimate of $\hat{\theta}_{G}(y)$ in (1) that would be assigned if we had access to an infinitely large sample would be of the form
+(2004, Theorem 2.2), this implies that the $Y_{n}$’s are asymptotically i.i.d., given the random limit $\tilde{g}$, with law $p_{\tilde{g}}$, which fits within the quasi-Bayesian framework of Fortini and Petrone (2020).
+In this perspective, $p_{\tilde{g}}$ represents the random distribution that would be assigned if we had access to an infinitely large sample, which makes $\tilde{g}$ the parameter of interest.
+Accordingly, (1) becomes
 $\hat{\theta}_{\tilde{g}}(y)=(y+1)\frac{p_{\tilde{g}}(y+1)}{p_{\tilde{g}}(y)}\qquad y\in\mathbb{N}_{0}.$ (4)
-An estimate of $\hat{\theta}_{G}(y)$ based on the sample $(Y_{1},\ldots,Y_{n})$ is obtained from (6) by replacing $\tilde{g}$ with $g_{n}$, i.e.
-$\hat{\theta}_{g_{n}}(y)=(y+1)\frac{p_{g_{n}}(y+1)}{p_{g_{n}}(y)}\qquad y\in\mathbb{N}_{0}.$ (7)
-Although the mixing distribution is concentrated on $\Theta_{d}=\{\vartheta_{1},\dots,\vartheta_{d}\}$, we are not restricting the parameter $\theta$ or its estimate to take values in $\Theta_{d}$.
-Under (4)-(5), $g_{n}$ converges $\mathds{P}$ a.s.
-to $\tilde{g}$, as $n\rightarrow+\infty$.
-Then, it follows that for every $y\in\mathbb{N}_{0}$, $\hat{\theta}_{g_{n}}(y)$ converges $\mathds{P}$ a.s.
+Differently from the Bayesian framework, which posits a prior distribution on $g$ such that the $Y_{n}$’s are i.i.d.
+given $g$, the quasi-Bayesian approach avoids the computational cost of recomputing the posterior distribution with each new observation, thereby offering substantial computational efficiency.
+2.2 Sequential estimation and credible intervals
+Under the learning model (2)-(3), an estimate of $\hat{\theta}_{\tilde{g}}(y)$ is obtained by replacing $\tilde{g}$ in (4) with $g_{n}$, i.e.
+$\hat{\theta}_{g_{n}}(y)=(y+1)\frac{p_{g_{n}}(y+1)}{p_{g_{n}}(y)}\qquad y\in\mathbb{N}_{0},$ (5)
+which is referred to as quasi-Bayes empirical Bayes.
+As a corollary of Theorem 2.2, for every $y\in\mathbb{N}_{0}$ it holds that $\hat{\theta}_{g_{n}}(y)$ converges $\mathds{P}$ a.s.
 to $\hat{\theta}_{\tilde{g}}(y)$, as $n\rightarrow+\infty$.
-Next theorem gives an asymptotic Gaussian approximation of the posterior distribution of $\hat{\theta}_{\tilde{g}}(y)$ for $y\in\mathbb{N}_{0}$.
+The following central limit theorem allows to measure uncertainty for $\hat{\theta}_{G_{n}}(y)$, building an asymptotic credible interval for $\hat{\theta}_{\tilde{g}}(y)$.
 Theorem 2.3.
@@
 $W_{\tilde{g}}(y)=\hat{\theta}^{2}_{\tilde{g}}(y)E_{Z\sim p_{\tilde{g}}}\left[\left(\sum_{j=1}^{d}\frac{k(Z\mid\vartheta_{j})\tilde{g}(\vartheta_{j})}{p_{\tilde{g}}(Z)}\left(\frac{k(y+1\mid\vartheta_{j})}{p_{\tilde{g}}(y+1)}-\frac{k(y\mid\vartheta_{j})}{p_{\tilde{g}}(y)}\right)\right)^{2}\right].$ (6)
-See Supplementary Material C.2 for the proof of Theorem 2.2.
-For the standard choice $\alpha_{n}=(\alpha+n)^{-\gamma}$ for $\alpha>0$ and $\gamma\in(1/2,1]$ then $b_{n}=(2\gamma-1)n^{2\gamma-1}$.
-Let $W_{g_{n}}(y)$ be defined as in (8), with $g_{n}$ in the place of $\tilde{g}$.
-Since $g_{n}$ converges to $\tilde{g}$ ($\mathds{P}$-a.s.), then $W_{g_{n}}(y)$ converges to $W_{\tilde{g}}(y)$, $\mathds{P}$-a.s.
-Therefore, the conditional distribution of $\hat{\theta}_{\tilde{g}}(y)$ given $\mathcal{G}_{n}$ is approximately Gaussian with mean $\hat{\theta}_{g_{n}}(y)$ and variance $W_{g_{n}}(y)$.
-This provides asymptotic credible intervals for $\hat{\theta}_{\tilde{g}}(y)$, for $y\in\mathbb{N}_{0}$.
-Remark 2.3.
-A sequential $g$-modeling methodology may be also considered from a frequentist perspective, by assuming i.i.d.
-observations $(Y_{1},\ldots,Y_{n})$ and applying Newton’s algorithm as a consistent estimate of a “true” $g$ (Tokdar et al., 2009; Martin and Tokdar, 2009).
-This approach would provide the same estimate $\hat{\theta}_{g_{n}}(y)$ of $\hat{\theta}_{G}(y)$ as in (7), and an upper bound for the rate of convergence (see Section 2.2).
-However, to the best of our knowledge, going beyond point estimation, for example by building asymptotic confidence intervals, is not feasible based on asymptotic results available for Newton’s algorithm.
-Indeed, these are based on stochastic approximation (Martin and Ghosh, 2008), but central limit theorems in this field does not apply to Newton’s algorithm.
+See C.3 for the proof of Theorem 2.3.
+If $W_{g_{n}}(y)$ is defined as in (6), with $g_{n}$ in the place of $\tilde{g}$, then for every $y\in\mathbb{N}_{0}$ it holds that $W_{g_{n}}(y)$ converges $\mathds{P}$-a.s.
+to $W_{\tilde{g}}(y)$, as $n\rightarrow+\infty$.
+Therefore, the conditional distribution of $\hat{\theta}_{\tilde{g}}(y)$ given $\mathcal{G}_{n}$ is approximately a Gaussian distribution with mean $\hat{\theta}_{g_{n}}(y)$ and variance $W_{g_{n}}(y)$.
+This result provides an asymptotic credible interval for $\hat{\theta}_{\tilde{g}}(y)$, for $y\in\mathbb{N}_{0}$.
+A sequential $g$-modeling methodology may be also considered from a frequentist perspective, where $(Y_{n})_{n\geq 1}$ are i.i.d.
+from a Poisson mixture model with a “true” mixing distribution $g^{\ast}$, and Newton’s algorithm is employed as a consistent estimate of $g^{\ast}$ (Tokdar et al., 2009; Martin and Tokdar, 2009).
+While this perspective leads to the same estimate $\hat{\theta}_{g_{n}}(y)$ as (5), constructing asymptotic confidence intervals remains out of reach, as the stochastic approximation framework underlying Newton’s method (Martin and Ghosh, 2008) does not apply to Gaussian central limit theorems.
 2.3 Frequentist guarantees
-The “true” Poisson compound model for $(Y_{n})_{n\geq 1}$ is the model under which, for any $n\geq 1$: i) $Y_{n}$ given $\theta^{\ast}_{n}$ is Poisson distributed with parameter $\theta^{\ast}_{n}$; ii) the $\theta^{\ast}_{n}$’s are independent and identically distributed from the “true” prior distribution, say the oracle prior $G^{\ast}$ with probability mass function $g^{\ast}$ on $\Theta_{d}$.
-If $G^{\ast}$ was known, then we could estimate $\hat{\theta}_{G}(y)$ by $\hat{\theta}_{g^{*}}(y)$, namely the plug-in estimate obtained from (1) by replacing $g$ with $g^{\ast}$.
-Under the “true” model, the Bayes risk of $\hat{\theta}_{g^{*}}(y)$ is
-$\mathcal{R}(g^{*},\hat{\theta}_{g^{*}})=E^{*}(\hat{\theta}_{g^{*}}(Y_{n})-\theta_{n}^{*})^{2}=\sum_{j=1}^{d}\sum_{y\in\mathbb{N}_{0}}^{\infty}(\hat{\theta}_{g^{*}}(y)-\vartheta_{j})^{2}k(y\mid\vartheta_{j})g^{*}(\vartheta_{j}),$
-with $E^{*}$ denoting the expected value with respect to the true model.
-Not knowing the oracle $g^{*}$, we employ the estimate $\hat{\theta}_{g_{n}}(y)$ in (7).
-In particular, under the “true” model, the Bayes risk of $\hat{\theta}_{g_{n}}(y)$ is
-$\mathcal{R}(g^{*},\hat{\theta}_{g_{n}})=E^{*}(\hat{\theta}_{g_{n}}(Y_{n})-\theta^{*}_{n})^{2}=\mathcal{R}(g^{*},\hat{\theta}_{g^{*}})+E^{*}(\hat{\theta}_{g_{n}}(Y_{n})-\hat{\theta}_{g^{*}}(Y_{n}))^{2}.$
-The Bayesian regret (of not knowing $g^{*}$ and using $g_{n}$ instead) is the difference between $\mathcal{R}(g^{*},\hat{\theta}_{g_{n}})$ and $\mathcal{R}(g^{*},\hat{\theta}_{g^{*}})$, i.e.
-$\text{Regret}(g_{n},g^{*})=E^{*}((\hat{\theta}_{g_{n}}(Y_{n})-\hat{\theta}_{g^{*}}(Y_{n}))^{2}=\sum_{j=1}^{d}\sum_{y\in\mathbb{N}_{0}}(\hat{\theta}_{g_{n}}(y)-\hat{\theta}_{g^{*}}(y))^{2}k(y\mid\vartheta_{j})g^{*}(\vartheta_{j}).$ (9)
-See Efron (2019) for details on (9). The next theorem provides the large $n$ asymptotic behaviour of the regret.
+The “true” Poisson compound model assumes that the observations $Y_{n}$’s are i.i.d.
+from a Poisson mixture model with a “true” mixing distribution, say the oracle prior $G^{\ast}$ on $\Theta\subseteq\mathbb{R}^{+}$.
+If $G^{\ast}$ was known, then we could estimate $\hat{\theta}_{G}(y)$ in (1) with the oracle $\hat{\theta}_{G^{*}}(y)$.
+Instead, not knowing the oracle prior $G^{*}$, we employ the estimate $\hat{\theta}_{g_{n}}(y)$ in (5).
+Here, Proposition 2.1 is leveraged in order to establish asymptotic frequentist guarantees of $\hat{\theta}_{g_{n}}(y)$ under the “true” model for $(Y_{n})_{n\geq 1}$, demonstrating consistency and asymptotic optimality in terms of vanishing excess Bayes risk or regret.
+Let $G^{\ast}$ be an oracle prior on $\Theta\subseteq\mathbb{R}^{+}$ that satisfies the assumptions of Proposition 2.1, and, for a fixed $\eta>0$, let $g^{\ast}_{\eta}$ be its $d_{\eta}$-KL-discretization.
+We define the regret of using $\hat{\theta}_{g_{n}}(y)$ instead of $\hat{\theta}_{g_{\eta}^{*}}(y)$ as
+$\text{Regret}(g_{n},g_{\eta}^{*})=E_{G^{*}}((\hat{\theta}_{g_{n}}(Y_{n})-\hat{\theta}_{g_{\eta}^{*}}(Y_{n}))^{2}=\int_{\Theta}\sum_{y\in\mathbb{N}_{0}}(\hat{\theta}_{g_{n}}(y)-\hat{\theta}_{g_{\eta}^{*}}(y))^{2}k(y\mid\theta)G^{\ast}(\mathrm{d}\theta)$ (7)
+with $E_{G^{*}}$ being the expected value with respect to the “true” model; see Efron (2019, Section 3) for details on (7).
+We denote by $\mathds{P}_{G^{*}}$ the probability measure under which the “true” model for $(Y_{n})_{n\geq 1}$ holds true.
 Theorem 2.4.
-Let $g_{n}$ be as in (5) with $\alpha_{n}=(\alpha+n)^{-\gamma}$ for $\alpha>0$ and $\gamma\in(1/2,1)$, and let $g^{\ast}$ be the probability mass function of the oracle prior $G^{\ast}$.
-Denote by $\mathds{P}^{*}$ a probability measure under which the “true” Poisson compound model for $(Y_{n})_{n\geq 1}$ holds true.
-Then, for every $\delta<2-1/\gamma$, it holds that
-$\text{Regret}(g_{n},g^{*})=o(n^{-\delta})\quad\mathds{P}^{\ast}\,\text{-}\,a.s.$
-See Supplementary Material C.3 for the proof of Theorem 2.4.
-The regret analysis of Theorem 2.4 is provided under the standard choice $\alpha_{n}=(1/(\alpha+n))^{\gamma}$, though other learning rates may be considered.
+Let $G^{\ast}$ be an oracle prior on $\Theta\subseteq\mathbb{R}^{+}$ such that $\sum_{y\in\mathbb{N}_{0}}y^{k}p_{G^{\ast}}(y)\leq m_{k}$ for some $k\in\mathbb{N}$ with $k\geq 2$ and $m_{k}\in\mathbb{R}$, and for a fixed $\eta>0$ let $g^{\ast}_{\eta}$ be the $d_{\eta}$-KL-discretization of $G^{\ast}$ on $\Theta_{\eta}$.
+For each $n\geq 1$, let $g_{n}(\vartheta_{j})$ be as in (3) for $\vartheta_{j}\in\Theta_{\eta}$, with $g_{0}(\vartheta_{j})>0$ for every $j=1,\dots,d_{\eta}$, and with $\alpha_{n}=(\alpha+n)^{-\gamma}$ for $\alpha>0$ and $\gamma\in(1/2,1]$.
+If $\sum_{y\in\mathbb{N}_{0}}d_{\eta}^{2y}p_{G^{\ast}}(y)<+\infty$ then, for every $y\in\mathbb{N}_{0}$,
+$\hat{\theta}_{g_{n}}(y)\rightarrow\hat{\theta}_{g^{\ast}_{\eta}}(y)\quad\mathds{P}_{G^{*}}\,\text{-}\,a.s.$ (8)
+Furthermore, if $\gamma<1$ then for every $\delta<2-1/\gamma$
+$\text{Regret}(g_{n},g_{\eta}^{\ast})=o(n^{-\delta})\quad\mathds{P}_{G^{*}}\,\text{-}\,a.s.$ (9)
+See C.4 for the proof of Theorem 2.4.
+Under the “true” model for the observations $Y_{n}$’s, with the oracle prior $G^{\ast}$ belonging to the class of light-tail distributions, Theorem 2.4 establishes that $\hat{\theta}_{g_{n}}(y)$ is a consistent (8) and asymptotically optimal (9) estimate of $\hat{\theta}_{g^{\ast}_{\eta}}(y)$.
+Although Theorem 2.4 is stated with respect to the standard choice $\alpha_{n}=(\alpha+n)^{-\gamma}$, alternative learning rates may also be considered.
 3 Numerical illustrations
-We validate empirically our sequential $g$-modeling on synthetic and real data.
-We compare the quasi-Bayes (NP-QB) estimate $\hat{\theta}_{g_{n}}$ of $\hat{\theta}_{G}$, at fixed $n\geq 1$, with the estimates obtained by the following methods: i) nonparametric empirical Bayes (NP-EB) (Robbins, 1956); ii) nonparametric maximum likelihood (NP-ML) (Jiang and Zhang, 2009); iii) nonparametric minimum squared Hellinger distance (NP-MD) (Jana et al., 2024); parametric Bayes empirical Bayes (P-EB) with a Gamma hyperprior on $G$ (Deely and Lindley, 1981).
-NP-EB is the sole to produce a closed-form estimate, i.e.
-the estimate obtained by replacing $p_{G}$ in (1) with the empirical distribution (2).
-NP-ML, N-PMD and P-EB require exact or approximate optimization algorithms: i) NP-ML and N-PMD estimates are obtained from Jana et al.
-(2024, Algorithm 1); ii) the P-EB estimate is obtained from Deely and Lindley (1981, Equation 4.4), with the parameters of the Gamma hyperprior estimated by maximum (marginal) likelihood.
-The empirical performance of the above estimates is evaluated by the root mean squared error (RMSE) and the mean absolute deviation (MAD).
+We validate our methodology empirically, by comparing the quasi-Bayes empirical Bayes (QB-EB) estimate $\hat{\theta}_{g_{n}}$, at fixed $n\geq 1$, with estimates from: i) nonparametric empirical Bayes (NP-EB) (Robbins, 1956); ii) nonparametric maximum likelihood (NP-ML) (Jiang and Zhang, 2009); iii) nonparametric minimum squared Hellinger distance (NP-MD) (Jana et al., 2024); parametric Bayes empirical Bayes (P-EB) with a Gamma hyperprior on $G$ (Deely and Lindley, 1981).
+NP-EB uniquely yields a closed-form estimate by replacing $p_{G}$ in (1) with the empirical distribution, while NP-ML, NP-MD and P-EB require exact or approximate optimization algorithms.
+NP-ML and NP-MD estimates follow the vertex direction algorithm Jana et al.
+(2024, Algorithm 1), which, as QB-EB, relies on a discretization of $G$ over a finite grid.
+P-EB estimates use Deely and Lindley (1981, Equation 4.4), with hyperparameters of $G$ estimated by maximum (marginal) likelihood.
+Performance is assessed via root mean squared error (RMSE) and mean absolute deviation (MAD).
 3.1 Synthetic data
-We set $n=50,100,200,500$, and generate $(Y_{1},\ldots,Y_{n})$ under the Poisson compound model, with $(\theta_{1},\ldots,\theta_{n})$ i.i.d from a Gamma prior $G$ with scale $2$ and shape $4$.
-For NP-QB, we set Newton’s algorithm as follows: i) $\theta$ takes value in $\Theta=\{0.01,0.02,\ldots,99.99,100\}$; ii) the number $d$ of mixture components is the cardinality of $\Theta$, i.e.
-$d=10,000$; iii) the initial guess $g_{0}$ is the Uniform distribution on $\Theta$; iv) the learning rate is $\alpha_{n}=(1+n)^{-1}$.
-Results are displayed in Figure 1 and Table 1.
-Figure 1: Data $Y_{i}$’s against the true $\theta_{i}$’s (black), and estimates of the $\theta_{i}$’s: NP-EB (green), NP-ML (cyan), NP-MD (magenta), P-EB (red) and NP-QB (blue).
-Table 1: Values of RMSE and MAD for the NP-EB, NP-ML, NP-MD, P-EB and NP-QB estimates.
-NP-EB NP-ML NP-MD P-EB NP-QB
-RMSE   $n=100$ 6.184 2.230 2.253 2.110 2.235
-RMSE   $n=500$ 3.737 2.238 2.228 2.244 2.236
-MAD     $n=100$ 4.100 1.722 1.731 1.632 1.719
-MAD     $n=500$ 2.459 1.773 1.768 1.782 1.774
+[…]
+Figure 1: Data $Y_{i}$’s against the $\theta_{i}$’s (black circle), and estimates of the $\theta_{i}$’s: NP-EB (green), NP-ML (cyan), NP-MD (magenta), P-EB (red) and QB-EB (blue).
+Table 1: Values of RMSE and MAD for NP-EB, NP-ML, NP-MD, P-EB and QB-EB estimates.
+NP-EB NP-ML NP-MD P-EB QB-EB
+RMSE   $n=100$ 2.516 1.695 1.648 1.478 1.675
+RMSE   $n=500$ 2.108 1.341 1.515 1.353 1.348
+MAD     $n=100$ 1.807 1.334 1.3 1.179 1.335
+MAD     $n=500$ 1.4 1.068 1.197 1.073 1.071
 […]
 3.2 Real data
-We consider a dataset of tweets with at least 50 retweets on Twitter in the period October 7 to November 6, 2011 (30 days).
-The dataset is available at https://snap.stanford.edu/seismic/, and it consists of $166,783$ tweets that generate $34,617,704$ retweets.
-The data provides information on the tweet ID, the tweet post time, the retweet post time and the number of followers of the poster/retweeter.
-For each tweet, we focus on the number of retweets generated within $30$ seconds from the tweet post time; in principle, the faster a retweet is, the more likely it was generated by a chatbot artificial intelligence rather than a human.
-The count data of fast retweets arrive at high-velocity, thus motivating the need of an online fast processing.
-From the first two rows of Table 2, $40,259$ tweets have 00 retweets in $30$ seconds, $28,339$ tweets have $1$ retweet in $30$ seconds, etc.
-Interest is in estimating the number of retweets expected for a tweet.
-Table 2 displays the estimates obtained through the application of NP-EB, NP-ML, NP-MD, P-FB and NP-QB.
-Table 2: Counts $n_{y}$ of the number of tweets that have $y$ retweets within $30$ seconds from the tweet post time, for $y=0,1,\ldots,7$, and corresponding estimates of the number of retweets expected.
+We analyze a dataset of tweets that received at least $50$ retweets on Twitter between October 7 and November 6, 2011, available at https://snap.stanford.edu/seismic/.
+It contains $166,783$ tweets generating $34,617,704$ retweets, along with metadata such as tweet ID, posting time, retweet time, and follower counts.
+For each tweet, we focus on the number of retweets received within the first $30$ seconds after posting: fast retweeting is often indicative of automated or bot-driven activity, as opposed to human behavior.
+The underlying retweet intensity can thus be viewed as a latent measure of a tweet’s instantaneous virality, or more generally its propensity to trigger immediate attention.
+Identifying tweets with unusually high intensity in real time could help detect automated activity or early-stage virality.
+Since these counts arrive rapidly in a streaming fashion, they motivate the need for a sequential, online estimation method such as the proposed QB-EB.
+Table 2 summarizes the empirical distribution of fast retweet counts, e.g., 40,259 tweets received none, 28,339 received one, etc.
+We apply NP-EB, NP-ML, NP-MD, P-FB, and QB-EB to estimate the intensity for each tweet, quantifying its fast-retweet potential.
+Results are shown in Table 2.
+Table 2: Counts $n_{y}$ of the number of tweets with $y$ fast retweets, and corresponding estimates.
 $y$ 0 1 2 3 4 5 6 7
@@
 P-EB 0.445 1.335 2.226 3.116 4.007 4.897 5.788 6.678
-NP-QB 0.607 0.818 1.632 3.076 4.007 4.519 5.023 5.707
-We refer to Supplementary Material D.2 for a version of Table 2 with more values of $y$, which shows that NP-QB competes with NP-ML, NP-MD and P-EB, and all these methods outperform NP-EB.
-The Supplementary Material D.2 includes another application to same Twitter dataset, but with respect to the number of followers generated by fast retweets.
-See also Supplementary Material D.3 for an application to the benchmark auto accident dataset (Efron and Hastie, 2021, Table 6.1).
+QB-EB 0.607 0.818 1.632 3.076 4.007 4.519 5.023 5.707
+We refer to D.2 for another application of the Twitter dataset, with respect to the followers generated by retweets, and for an application to the benchmark accident dataset (Efron and Hastie, 2021, Table 6.1).
 4 Discussion
-The Supplementary Material E contains an extension of our sequential $g$-modeling to the multidimensional Poisson compound decision problem (Brown et al., 1985; Johnstone, 1986).
-Further extensions may be considered with respect to compound decision problems beyond the Poisson setting, with applications to high-dimensional problem, e.g.
-exponential families (Zhang, 2003).
-Sequential $f$-modeling is also an interesting direction for future research (Brown et al., 2013).
+An extension of the sequential $g$-modeling methodology to the multidimensional Poisson compound decision problem (Brown et al., 1985) is presented in Section E.
+While our focus is on the Poisson setting, the methodology is readily extendable to other discrete linear exponential families (Banerjee et al., 2021), as well as to the Gaussian setting (Jiang and Zhang, 2009; Soloff et al., 2024).
 Appendix
@@
 C Proofs for Section 2
-C.1 Proof of Theorem 2.1
+C.1 Proof of Proposition 2.1
+We first prove that there exists a probability mass function $f\in\mathcal{S}_{\eta}$ such that $KL(p_{G}||p_{f})<2\eta.$ Since the $k$-moment of the Poisson distribution with parameter $\lambda$ is greater than $\lambda^{k}$, then the assumption $\sum_{y\in\mathbb{N}_{0}}y^{k}p_{G}(y)\leq m_{k}$ implies that
+$\int_{0}^{\infty}\theta^{k}G(\mathrm{d}\theta)\leq m_{k}.$
+For every $i=1,\dots,d_{\eta}$, let $w_{i}=G((\vartheta_{i-1},\vartheta_{i}])$ with $\vartheta_{0}=0$. Let $F$ be the probability measure
+$F=\sum_{i=1}^{d_{\eta}}w_{i}\delta_{\vartheta_{i}}+G(\vartheta_{d_{\eta}},+\infty)\delta_{\vartheta_{d_{\eta}}},$
+and let $f$ denote its density with respect to the counting measure on $\Theta_{\eta}$. Then
+$F=\sum_{i=1}^{d_{\eta}}\delta_{\vartheta_{i}}\int_{I_{i}}G(\mathrm{d}\theta)+\delta_{\vartheta_{d_{\eta}}}\int_{I_{d_{\eta}+1}}G(\mathrm{d}\theta),$
+where $I_{i}=(\vartheta_{i-1},\vartheta_{i}]$ for $i=1,\dots,d_{\eta}$ and $I_{d_{\eta}+1}=(\vartheta_{d_{\eta}},+\infty)$.
+Since the Kullback-Leibler divergence is convex, then, denoting $M=d_{\eta}\eta$, we can write that
+$KL(p_{G}||p_{f})$ $\leq\sum_{y\in\mathbb{N}_{0}}\sum_{i=1}^{d_{\eta}}\int_{I_{i}}\log\left(\frac{e^{-\theta}\theta^{y}}{e^{-\vartheta_{i}}\vartheta_{i}^{y}}\right)\frac{e^{-\theta}\theta^{y}}{y!}G(\mathrm{d}\theta)+\sum_{y\in\mathbb{N}_{0}}\int_{M}^{\infty}\log\left(\frac{e^{-\theta}\theta^{y}}{e^{-M}M^{y}}\right)\frac{e^{-\theta}\theta^{y}}{y!}G(\mathrm{d}\theta)$
+$<\int_{0}^{M}\eta G(\mathrm{d}\theta)+\int_{M}^{\infty}\sum_{y\in\mathbb{N}_{0}}y\log(\theta)\frac{e^{-\theta}\theta^{y}}{y!}G(\mathrm{d}\theta)$
+$\leq\eta+\int_{M}^{\infty}\theta\log(\theta)G(\mathrm{d}\theta)$
+$\leq\eta+\frac{\log M}{M^{k-1}}\int_{M}^{\infty}\theta^{k}G(\mathrm{d}\theta)$
+$\leq 2\eta.$
+We now prove that the minimization problem $\mbox{\sl argmin}_{f\in\mathcal{S}_{\eta}}KL(p_{G}||p_{f})$ has a unique solution.
+Representing each probability mass function $f$ with support $\Theta_{\eta}$ as a vector $\mathbf{f}$ taking values in the simplex $\Delta_{\eta}=\{\mathbf{v}=(\mathbf{v}_{1},\dots,\mathbf{v}_{d_{\eta}-1}):\mathbf{v}_{i}\geq 0,\sum_{i=1}^{d_{\eta}-1}\mathbf{v}_{i}\leq 1\},$ we can see $\ell(\mathbf{f})=KL(p_{G^{*}}||p_{f})$ as a strictly convex function defined on a convex and compact set.
+Thus the global minimum $\mathbf{g}_{\eta}$ exists and is unique.
+$\square$
+C.2 Proof of Theorem 2.2
 For every $n\geq 0$ and every $j=1,\dots,d$,
@@
 […]
-C.2 Proof of Theorem 2.2
+C.3 Proof of Theorem 2.3
 For every $z\in\mathbb{N}_{0}$ let $\tilde{g}(\vartheta_{j}\mid z)=\frac{k(z\mid\vartheta_{j})\tilde{g}(\vartheta_{j})}{p_{\tilde{g}}(z)}$.
@@
 Since this is true for every $\mathbf{u}$ with $||\mathbf{u}||=1$, then $b_{n}^{1/2}(g_{n}-\tilde{g})$ converges in the sense of almost sure conditional convergence to $\mathcal{N}(0,V_{\tilde{g}})$.
-Now suppose that $\alpha_{n}=(\alpha+n)^{-\gamma}$. With $b_{n}=(2\gamma-1)n^{2\gamma-1}$, the conditions (F.20) hold. Indeed, $b_{n}\uparrow+\infty$,
-$\sum_{n=1}^{\infty}\alpha_{n}^{4}b_{n}^{2}=\sum_{n=1}^{\infty}\left(\frac{1}{\alpha+n}\right)^{4\gamma}(2\gamma-1)^{2}\alpha^{2\gamma}n^{4\gamma-2}<+\infty,$
-$b_{n}\sup_{k\geq n}\alpha_{k}^{2}=(2\gamma-1)n^{2\gamma-1}\left(\frac{1}{\alpha+n}\right)^{2\gamma}\rightarrow 0,$
-and
-$b_{n}\sum_{k\geq n}{\alpha_{k}^{2}}=(2\gamma-1)n^{2\gamma-1}\sum_{k\geq n}\left(\frac{1}{\alpha+k}\right)^{2\gamma}\rightarrow 1.$
+∎
 of Theorem 2.3.
-The proof is based on Theorem C.2 and Theorem B.5. Let us write
+The proof is based on Theorem C.2 and Theorem B.5. We first prove that (C.5) holds under the assumptions of Theorem 2.3. We can write that
+$b_{n}\uparrow+\infty,\quad\sum_{n\geq 1}\alpha_{n}^{4}b_{n}^{2}<+\infty,\quad b_{n}\sum_{k\geq n}\alpha_{k}^{2}\rightarrow 1.$
+In particular, $\alpha_{n}^{4}b_{n}^{2}\rightarrow 0$, as $n\rightarrow\infty$. Since $(\alpha_{n})$ is non increasing, then
+$b_{n}\sup_{k\geq n}\alpha_{k}^{2}=b_{n}\alpha_{n}^{2}\rightarrow 0.$
+Thus, (C.5) holds. Let us write
 $\hat{\theta}_{g}(y)=h_{y}(\mathbf{g}),$
@@
 $=\hat{\theta}_{\tilde{g}}(y)^{2}\sum_{z\in\mathbb{N}_{0}}p_{\tilde{g}}(z)\left[\sum_{i=1}^{d}\frac{\tilde{g}(\vartheta_{i}\mid z)}{\tilde{g}(\vartheta_{i})}(\tilde{g}(\vartheta_{i}\mid y+1)-\tilde{g}(\vartheta_{i}\mid y))\right]^{2}.$
-C.3 Proof of Theorem 2.4
-The proof of Theorem 2.4 is based on the following lemma C.3 and on the rate of convergence of $g_{n}$ to $g$, under i.i.d.
+C.4 Proof of Theorem 2.4
+The proof of Theorem 2.4 is based on lemmas C.3, C.7 and on the rate of convergence of $g_{n}$ to $g_{\eta}^{*}$, under i.i.d.
 observations.
@@
 We provide for completeness the proof in Lemma C.5.
+In this section, we will occasionally represent probability mass functions $g$ on $\Theta_{\eta}$ as vectors $\mathbf{g}$, taking values either in the simplex $\Delta_{\eta}=\{\mathbf{v}=(\mathbf{v}_{1},\dots,\mathbf{v}_{d_{\eta}-1}):\mathbf{v}_{i}\geq 0\;(i=1,\dots,d_{\eta}-1),\;\sum_{i=1}^{d_{\eta}-1}\mathbf{v}_{i}\leq 1\}$ or in $\mathbb{R}^{d_{\eta}}$, depending on the context, which will be specified when needed.
+Similarly, the Poisson kernel is represented as a vector in $\Delta_{\eta}$ or $\mathbb{R}^{d_{\eta}}$, by defining $\mathbf{k}_{y}(\vartheta_{j})=k(y\mid\vartheta_{j})$.
+We denote component-wise multiplication of vectors by $\circ$.
 Lemma C.3.
-Under the assumptions of Theorem 2.4, there exists a constant $C=C(d,\vartheta_{1},\vartheta_{d})$ such that for every $n\geq 1$
-${\rm Regret}(g_{n},g^{*})\leq C||\mathbf{g}_{n}-\mathbf{g}^{*}||^{2},$
-where $\mathbf{g}_{n}$ and $\mathbf{g}^{*}$ are the representations of $g_{n}$ and $g^{*}$ in $\mathbb{R}^{d}$.
+If
+$\sum_{y\in\mathbb{N}_{0}}d_{\eta}^{2y}p_{G^{*}}(y)<+\infty,$ (C.6)
+then the $d_{\eta}$-KL-discretization $g_{\eta}^{*}$ of $G^{*}$ satisfies for every $i=1,\dots,d_{\eta}$ the equation in $g$:
+$g(\vartheta_{i})\left(\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{i})}{p_{g}(y)}p_{G^{*}}(y)-1\right)=0.$ (C.7)
 Proof.
-We can write that
-$\sum_{y\in\mathbb{N}_{0}}(\hat{\theta}_{g_{n}}(y)-\hat{\theta}_{g^{*}}(y))^{2}p_{g^{*}}(y)$
-$=\sum_{y\in\mathbb{N}_{0}}\left(\frac{\sum_{i=1}^{d}\vartheta_{i}^{y+1}e^{-\vartheta_{i}}g_{n}(\vartheta_{i})}{\sum_{i=1}^{d}\vartheta_{i}^{y}e^{-\vartheta_{i}}g_{n}(\vartheta_{i})}-\frac{\sum_{i=1}^{d}\vartheta_{i}^{y+1}e^{-\vartheta_{i}}g^{*}(\vartheta_{i})}{\sum_{i=1}^{d}\vartheta_{i}^{y}e^{-\vartheta_{i}}g^{*}(\vartheta_{i})}\right)^{2}\frac{1}{y!}\sum_{i=1}^{d}\vartheta_{i}^{y}e^{-\vartheta_{i}}g^{*}(\vartheta_{i})$
-$\leq 2\sum_{y\in\mathbb{N}_{0}}\frac{1}{y!}\frac{\left(\sum_{i=1}^{d}\vartheta_{i}^{y+1}e^{-\vartheta_{i}}\left(g_{n}(\vartheta_{i})-g^{*}(\vartheta_{i})\right)\right)^{2}}{\sum_{i=1}^{d}\vartheta_{i}^{y}e^{-\vartheta_{i}}g^{*}(\vartheta_{i})}$
-$+2\sum_{y\in\mathbb{N}_{0}}\frac{1}{y!}\frac{\left(\sum_{i=1}^{d}\vartheta_{i}^{y+1}e^{-\vartheta_{i}}g_{n}(\vartheta_{i})\right)^{2}}{\left(\sum_{i=1}^{d}\vartheta_{i}^{y}e^{-\vartheta_{i}}g_{n}(\vartheta_{i})\right)^{2}\sum_{i=1}^{d}\vartheta_{i}^{y}e^{-\vartheta_{i}}g^{*}(\vartheta_{i})}\left(\sum_{i=1}^{d}\vartheta_{i}^{y}e^{-\vartheta_{i}}(g_{n}(\vartheta_{i})-g^{*}(\vartheta_{i})\right)^{2}$
-$\leq 2\sum_{y\in\mathbb{N}_{0}}\frac{1}{y!}\frac{\sum_{i=1}^{d}\vartheta_{i}^{2y+2}e^{-2\vartheta_{i}}||g_{n}-g^{*}||^{2}}{\sum_{i=1}^{d}\vartheta_{i}^{y}e^{-\vartheta_{i}}g^{*}(\vartheta_{i})}+2\sum_{y\in\mathbb{N}_{0}}\frac{1}{y!}\frac{\vartheta_{d}^{2}\sum_{i=1}^{d}\vartheta_{i}^{2y}e^{-2\vartheta_{i}}||g_{n}-g^{*}||^{2}}{\sum_{i=1}^{d}\vartheta_{i}^{y}e^{-\vartheta_{i}}g^{*}(\vartheta_{i})}$
-$\leq 4\vartheta_{d}^{2}||\mathbf{g}_{n}-\mathbf{g}^{*}||^{2}\sum_{y\in\mathbb{N}_{0}}\frac{1}{y!}\frac{\sum_{i=1}^{d}\vartheta_{i}^{2y}e^{-2\vartheta_{i}}}{\sum_{i=1}^{d}\vartheta_{i}^{y}e^{-\vartheta_{i}}g^{*}(\vartheta_{i})}$
-$\leq C||\mathbf{g}_{n}-\mathbf{g}^{*}||^{2},$
-with $C=4d\vartheta_{d}^{2}e^{\vartheta_{d}^{2}/\vartheta_{1}-2\vartheta_{1}+\vartheta_{d}}$. Thus,
-${\rm Regret}(g_{n},g^{*})\leq C||\mathbf{g}_{n}-\mathbf{g}^{*}||^{2}.$
+Representing probability mass functions on $\Theta_{\eta}$ as vectors in $\Delta_{\eta},$ we can see $\ell(\mathbf{f})=KL(p_{G^{*}}||p_{f})$ as a function on $\Delta_{\eta}$.
+The assumption (C.6) implies that
+$\int_{0}^{\infty}e^{\theta(d_{\eta}^{2}-1)}G^{*}(d\theta)$ $=\int_{0}^{\infty}\sum_{y\in\mathbb{N}_{0}}e^{\log(d_{\eta}^{2})y}e^{-\theta}\frac{\theta^{y}}{y!}G^{*}(d\theta)$ (C.8)
+$=\sum_{y\in\mathbb{N}_{0}}e^{2y\log d_{\eta}}p_{G^{*}}(y)<+\infty.$
+Thus, for every $i=1,\dots,d_{\eta}$,
+$\frac{\partial}{\partial\mathbf{g}_{i}}\ell(\mathbf{g})=\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{i})}{p_{\mathbf{g}}(y)}p_{G^{*}}(y)\leq e^{\vartheta_{1}-\vartheta_{d_{\eta}}}\sum_{y\in\mathbb{N}_{0}}\left(\frac{\vartheta_{d_{\eta}}}{\vartheta_{1}}\right)^{y}\int_{\Theta}\frac{e^{-\theta}\theta^{y}}{y!}{G^{*}}(\mathrm{d}\theta)\leq e^{\vartheta_{1}-\vartheta_{d_{\eta}}}\int e^{\theta(d_{\eta}-1)}G^{*}(\mathrm{d}\theta)<+\infty$
+by (C.8) Analogously, for every $i=1,\dots,d_{\eta}$,
+$\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{i})^{2}}{p_{\mathbf{g}}(y)^{2}}p_{G^{*}}(y)\leq e^{\vartheta_{1}-\vartheta_{d_{\eta}}}\int e^{\theta(d_{\eta}^{2}-1)}G^{*}(\mathrm{d}\theta)<+\infty.$ (C.9)
+Thus, $\ell(\mathbf{g})$ is twice continuously differentiable on $\Delta^{d}$. Denoting by $k_{y}(\vartheta_{i})=k(y\mid\vartheta_{i})$ for $i=1,\dots,\vartheta_{d_{\eta}}$, we can write
+$\nabla_{\mathbf{g}}KL(p_{G^{*}}||p_{\mathbf{g}})=-\sum_{y\in\mathbb{N}_{0}}\frac{k_{y}-k(y\mid\vartheta_{d_{\eta}})}{p_{\mathbf{g}}(y)}p_{G^{*}}(y).$
+The $(i,j)$ element of the Hessian matrix $\mathbf{M}$ of $\ell(\mathbf{g})$ is
+$\mathbf{M}_{i,j}$ $=\sum_{y\in\mathbb{N}_{0}}\frac{(k(y\mid\vartheta_{i})-k(y\mid\vartheta_{d_{\eta}}))(k(y\mid\vartheta_{j})-k(y\mid\vartheta_{d_{\eta}}))}{p_{\mathbf{g}}(y)^{2}}p_{G^{*}}(y).$
+for $i,j=1,\dots,d-1$.
+To prove (C.7), we apply the method of Lagrange multipliers for the constrained minimum of twice continuously differentiable functions.
+The solution $\mathbf{g}_{\eta}^{*}$ satisfies, for same $\lambda_{i}\geq 0$ ($i=1,\dots,d_{\eta}$) with $\lambda_{i}>0$ for at least one $i$,
+$\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{i})-k(y\mid\vartheta_{d_{\eta}})}{p_{\mathbf{g}_{\eta}^{*}}(y)}p_{G^{*}}(y)-\lambda_{i}+\lambda_{d_{\eta}}=0\quad(i=1,\dots,d_{\eta}-1),$ (C.10)
+and
+$\lambda_{1}g_{\eta}^{*}(\vartheta_{1})=0,\;\dots,\;\lambda_{d_{\eta}-1}g_{\eta}^{*}(\vartheta_{d_{\eta}-1})=0,\;\lambda_{d_{\eta}}(\sum_{i=1}^{d_{\eta}-1}g_{\eta}^{*}(\vartheta_{i})-1)=0.$ (C.11)
+Multiplying (C.10) by $g_{\eta}^{*}(\vartheta_{i})$ and taking the sum with respect to $i=1,\dots,d_{\eta}-1$, we obtain by (C.11)
+$\lambda_{d_{\eta}}=-\sum_{i=1}^{d_{\eta}-1}\sum_{y\in\mathbb{N}_{0}}g_{\eta}^{*}(\vartheta_{i})\frac{k(y\mid\vartheta_{i})}{p_{g_{\eta}^{*}}(y)}p_{G^{*}}(y)+(1-g_{\eta}^{*}(\vartheta_{d_{\eta}}))\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{d_{\eta}})}{p_{g_{\eta}^{*}}(y)}p_{G^{*}}(y).$
+Thus
+$\lambda_{d_{\eta}}-\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{d_{\eta}})}{p_{g_{\eta}^{*}}(y)}p_{G^{*}}(y)=-\sum_{i=1}^{d_{\eta}}\sum_{y\in\mathbb{N}_{0}}g_{\eta}^{*}(\vartheta_{i})\frac{k(y\mid\vartheta_{i})}{p_{g_{\eta}^{*}}(y)}p_{G^{*}}(y)=-1.$
+Substituting in (C.10), we obtain
+$\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{i})}{p_{\mathbf{g}_{\eta}^{*}}(y)}p_{G^{*}}(y)-\lambda_{i}-1=0\quad(i=1,\dots,d_{\eta}-1),$
+Thus, by (C.11),
+$g_{\eta}^{*}(\vartheta_{i})\left(\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{i})}{p_{g_{\eta}^{*}}(y)}p_{G^{*}}(y)-1\right)=0$ (C.12)
+for every $i=1,\dots,d_{\eta}-1$. Since
+$\sum_{i=1}^{d_{\eta}}\sum_{y\in\mathbb{N}_{0}}g_{\eta}^{*}(\vartheta_{i})\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{i})}{p_{g_{\eta}^{*}}(y)}p_{G^{*}}(y)=1,$
+then (C.12) holds also for $i=d_{\eta}$.
+$\square$
 Lemma C.4.
-Under the assumptions of Theorem 2.4, as $n\rightarrow\infty$ the stochastic approximation in $\Delta^{d}$:
+Under the assumptions of Theorem 2.4, as $n\rightarrow\infty$ the stochastic approximation in $\Delta_{\eta}=\{\mathbf{v}=(\mathbf{v}_{1},\dots,\mathbf{v}_{d_{\eta}-1}):\mathbf{v}_{i}\geq 0,\sum_{i=1}^{d_{\eta}-1}\mathbf{v}_{i}\leq 1\}$:
 $\mathbf{g}_{n+1}=\mathbf{g}_{n}+\alpha_{n+1}\mathbf{h}(\mathbf{g}_{n})+\alpha_{n+1}{\bm{\epsilon}}_{n+1},$ (C.13)
-with $\alpha_{n}>0$, $\sum_{n=1}^{\infty}\alpha_{n}=\infty$, $\sum_{n=1}^{\infty}\alpha_{n}^{2}<\infty$,
-$\mathbf{h}(\mathbf{g})=\mathbf{g}\circ\left(\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}}{p_{\mathbf{g}}(y)}p_{\mathbf{g}^{*}}(y)-\bf 1\right),$
-${\bm{\epsilon}}_{n+1}=\mathbf{g}_{n}\circ\left(\frac{\mathbf{k}_{Y_{n+1}}}{p_{\mathbf{g}_{n}}(Y_{n+1})}-\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}}{p_{\mathbf{g}_{n}}(y)}p_{\mathbf{g}^{*}}(y)\right),$
-and initial state $\mathbf{g}_{0}\in\mathring{\Delta}^{d}$, converges to $\mathbf{g}^{*}$, $\mathds{P}^{*}$-a.s.
+with
+$\mathbf{h}(\mathbf{g})=\mathbf{g}\circ\left(\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}}{p_{\mathbf{g}}(y)}p_{G^{*}}(y)-\bf 1\right),$
+${\bm{\epsilon}}_{n+1}=\mathbf{g}_{n}\circ\left(\frac{\mathbf{k}_{Y_{n+1}}}{p_{\mathbf{g}_{n}}(Y_{n+1})}-\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}}{p_{\mathbf{g}_{n}}(y)}p_{G^{*}}(y)\right),$
+and initial state $\mathbf{g}_{0}\in\mathring{\Delta}^{d_{\eta}}$, converges to $\mathbf{g}_{\eta}^{*}$, $\mathds{P}_{G^{*}}$-a.s.
 Proof.
-Let $J=\partial\Delta^{d}\cup\{g^{*}\}$.
-We first prove that ${\rm d}(\mathbf{g}_{n},J)\rightarrow 0$, $\mathds{P}^{*}$-a.s., as $n\rightarrow\infty$, where ${\rm d}$ denotes the Euclidean distance in $\mathbb{R}^{d-1}$.
-The result follows from Theorem 2.2.1′′ in Chen (2005), if we can prove that the following conditions hold:
-  • A1
+Let $J$ be the set of vectors in $\Delta^{d_{\eta}}$ that satisfy (C.7). According to Theorem 2.2.1 in Chen (2005), if the following conditions hold:
+  • S1
     $\alpha_{n}>0$, $\alpha_{n}\rightarrow 0$, as $n\rightarrow\infty$ and $\sum_{n=1}^{\infty}\alpha_{n}=\infty$;
-  • A2
-    $J$ is closed, and there exists a is continuously differentiable function $\mathscr{l}:\Delta^{d}\rightarrow\mathbb{R}$ such that $\sup_{\delta_{1}\leq{\rm d}(\mathbf{g},J)\leq\delta_{2}}\mathbf{h}(\mathbf{g})^{T}\nabla_{\mathbf{g}}\mathscr{l}<0$ for every $\delta_{2}>\delta_{1}>0$:
-  • A3
-    $\lim_{T\rightarrow\infty}\limsup_{n\rightarrow\infty}\frac{1}{T}||\sum_{i=n}^{m(n,T)}\alpha_{i+1}\epsilon_{i+1}||=0,$ $\mathds{P}^{*}$-a.s., where $m(n,T)=\max\{m:\sum_{i=n}^{m}\alpha_{i+1}\leq T\}$;
-  • A4
-    $\mathbf{h}$ is measurable and locally bounded.
-Let $\mathscr{l}$ be defined as
-$\mathscr{l}(\mathbf{g})=\sum_{y\in\mathbb{N}_{0}}\log\left(\frac{p_{\mathbf{g}^{*}}(y)}{p_{\mathbf{g}}(y)}\right)p_{\mathbf{g}^{*}}(y).$
-A1 holds by assumption; A3 holds since $({\bm{\epsilon}}_{n})$ is a martingale difference, and A4 is obvious.
-We now prove A2. The set $J$ is closed and $\mathscr{l}$ is continuously differentiable. Moreover, for $\mathbf{g}\in\mathring{(\Delta^{d})}$,
-$\mathbf{h}(\mathbf{g})^{T}\nabla_{\mathbf{g}}\mathscr{l}$ $=-\left(\mathbf{g}\circ\left(\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}}{p_{\mathbf{g}}(y)}p_{\mathbf{g}^{*}}(y)-\bf 1\right)\right)^{T}\left(\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}-\mathbf{k}_{y}(\vartheta_{d})\mathbf{1}}{p_{\mathbf{g}}(y)}p_{\mathbf{g}^{*}}(y)\right)$
-$=-\left(\sum_{i=1}^{d}\mathbf{g}_{i}\left(\sum_{y\in\mathbb{N}_{0}}k_{y}(\vartheta_{i})\frac{p_{\mathbf{g}^{*}}(y)}{p_{\mathbf{g}}(y)}-1\right)^{2}\right)$
-$\leq-\left(\sum_{i=1}^{d}\mathbf{g}_{i}\left(\sum_{y\in\mathbb{N}_{0}}k_{y}(\vartheta_{i})\frac{p_{\mathbf{g}^{*}}(y)}{p_{\mathbf{g}}(y)}-1\right)\right)^{2}=0,$
-where, in the inequality, we have used Jensen inequality.
-We now prove that the inequality is strict, unless $\mathbf{g}=\mathbf{g}^{*}$.
-By contradiction, assume that there exist $\mathbf{g}^{\prime}$ and $\mathbf{g}^{\prime\prime}$ with $\mathbf{g}^{\prime}\neq\mathbf{g}^{\prime\prime}$ such that $\sum_{y\in\mathbb{N}_{0}}k_{y}(\vartheta_{i})\frac{p_{\mathbf{g}^{*}}(y)}{p_{\mathbf{g}}(y)}=1$ for every $i=1,\dots,d$ both for $\mathbf{g}=\mathbf{g}^{\prime}$ and for $\mathbf{g}=\mathbf{g}^{\prime\prime}$.
+  • S2
+    $J$ is closed and there exists a is continuously differentiable function $\mathscr{l}:\Delta^{d_{\eta}}\rightarrow\mathbb{R}$ such that $\mathscr{l}(J)$ is nowhere dense in $\mathbb{R}$ and $\sup_{\delta_{1}\leq{\rm d}(\mathbf{g},J)\leq\delta_{2}}\mathbf{h}(\mathbf{g})^{T}\nabla_{\mathbf{g}}\mathscr{l}<0$ for every $\delta_{2}>\delta_{1}>0$;
+  • S3
+    $\lim_{T\rightarrow\infty}\limsup_{n\rightarrow\infty}\frac{1}{T}||\sum_{i=n}^{m(n,T)}\alpha_{i+1}\epsilon_{i+1}||=0,$ $\mathds{P}_{G^{*}}$-a.s., where $m(n,T)=\max\{m:\sum_{i=n}^{m}\alpha_{i+1}\leq T\}$;
+  • S4
+    $\mathbf{h}$ is measurable and locally bounded,
+then
+$d(\mathbf{g}_{n},J)\rightarrow 0,$ (C.14)
+$\mathds{P}_{G^{*}}$-a.s., as $n\rightarrow\infty$, where $d$ denotes the Euclidean distance in $\mathbb{R}^{d_{\eta}}$.
+S1 holds by assumption; S3 holds since $({\bm{\epsilon}}_{n})$ is a martingale difference, and H4 is obvious.
+We now prove S2. Let
+$\mathscr{l}(\mathbf{g})=\sum_{y\in\mathbb{N}_{0}}\log\left(\frac{p_{G^{*}}(y)}{p_{\mathbf{g}}(y)}\right)p_{G^{*}}(y).$
+By (C.6), the function $\mathscr{l}(\mathbf{g})=KL(p_{G^{*}}||p_{\mathbf{g}})$ is twice continuously differentiable and strictly convex in $\Delta^{d_{\eta}}$, as proved in Lemma C.3, with gradient
+$\nabla_{\mathbf{g}}\mathscr{l}(\mathbf{g})=-\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}-\mathbf{k}_{y}(\vartheta_{d_{\eta}})\mathbf{1}}{p_{g}(y)}p_{G^{*}}(y).$
+The set $J$ is closed.
+We now prove that $\mathscr{l}(J)$ is nowhere dense in $\mathbb{R}$.
+The set $J$ is made of a finite number of isolated points.
+Indeed, let $\mathbf{g}$ be a point in $J$ and let $N=\{i\in\{1,\dots,{d_{\eta}}\}:\mathbf{g}_{i}=0\}$.
+We prove that the conditions in (C.7) identify the values of $\mathbf{g}_{i}$ for $i\not\in N$.
+The property is obvious if $N$ contains ${d_{\eta}}-1$ points.
+Otherwise, assume by contradiction that there exist $\mathbf{g}^{\prime}$ and $\mathbf{g}^{\prime\prime}$ with $\mathbf{g}^{\prime}\neq\mathbf{g}^{\prime\prime}$ such that $\mathbf{g}_{i}^{\prime}=\mathbf{g}_{i}^{\prime\prime}=0$ for $i\in N$ and $\sum_{y\in\mathbb{N}_{0}}k_{y}(\vartheta_{i})\frac{p_{G^{*}}(y)}{p_{\mathbf{g}}(y)}=1$ for every $i\not\in N$ both for $\mathbf{g}=\mathbf{g}^{\prime}$ and for $\mathbf{g}=\mathbf{g}^{\prime\prime}$.
 Then, for every $a\in(0,1)$
-$1=\sum_{i=1}^{d}(a\mathbf{g}^{\prime}_{i})+(1-a)\mathbf{g}^{\prime\prime}_{i})\sum_{y\in\mathbb{N}_{0}}k_{y}(\vartheta_{i})\frac{p_{g^{*}}(y)}{ap_{\mathbf{g}^{\prime}}(y)+(1-a)p_{\mathbf{g}^{\prime\prime}}(y)}$
-$<\sum_{i=1}^{d}(a\mathbf{g}^{\prime}_{i}+(1-a)\mathbf{g}^{\prime\prime}_{i})\left(a\sum_{y\in\mathbb{N}_{0}}k_{y}(\vartheta_{i})\frac{p_{\mathbf{g}^{*}}(y)}{p_{\mathbf{g}^{\prime}}(y)}+(1-a)\sum_{y\in\mathbb{N}_{0}}k_{y}(\vartheta_{i})\frac{p_{\mathbf{g}^{*}}(y)}{p_{\mathbf{g}^{\prime\prime}}(y)}\right)$
+$1=\sum_{i\not\in N}(a\mathbf{g}^{\prime}_{i})+(1-a)\mathbf{g}^{\prime\prime}_{i})\sum_{y\in\mathbb{N}_{0}}k_{y}(\vartheta_{i})\frac{p_{G^{*}}(y)}{ap_{\mathbf{g}^{\prime}}(y)+(1-a)p_{\mathbf{g}^{\prime\prime}}(y)}$
+$<\sum_{i\not\in N}(a\mathbf{g}^{\prime}_{i}+(1-a)\mathbf{g}^{\prime\prime}_{i})\left(a\sum_{y\in\mathbb{N}_{0}}k_{y}(\vartheta_{i})\frac{p_{G^{*}}(y)}{p_{\mathbf{g}^{\prime}}(y)}+(1-a)\sum_{y\in\mathbb{N}_{0}}k_{y}(\vartheta_{i})\frac{p_{G^{*}}(y)}{p_{\mathbf{g}^{\prime\prime}}(y)}\right)$
 $=1,$
-which leads to a contradiction. Hence A2 holds, and ${\rm d}(\mathbf{g}_{n},J)\rightarrow 0$ $\mathds{P}^{*}$-a.s., as $n\rightarrow\infty$.
-To conclude, we need to show that $\mathds{P}^{*}(\liminf_{n}{\rm d}(\mathbf{g}_{n},\partial\Delta^{d})=0)=0$.
-Since ${\rm d}(\mathbf{g}^{*},\partial\Delta^{d})>0$ and ${\rm d}(\mathbf{g}_{n},\{\mathbf{g}^{*}\}\cup\partial\Delta^{d})\rightarrow 0$ $\mathds{P}-a.s.$, then $\mathds{P}^{*}(\liminf_{n}{\rm d}(\mathbf{g}_{n},\partial\Delta^{d})=0)>0$ if and only if there exists $M$ with $\mathds{P}^{*}(M)>0$ such that ${\rm d}(\mathbf{g}_{n}(\omega),\partial\Delta^{d})\rightarrow 0$ for every $\omega\in M$, as $n\rightarrow\infty$.
-Let us make this assumption as a contradiction hypothesis.
-Define the function $\mathcal{L}(\mathbf{g}_{n})=\sum_{i=1}^{d}\log(g^{*}(\vartheta_{i})/g_{n}(\vartheta_{i}))g^{*}(\vartheta_{i})$.
+which leads to a contradiction.
+Thus $J$ is made of a finite number of isolated points.
+It follows that $\{\mathscr{l}(\mathbf{g}):\mathbf{g}\in J\}$ is nowhere dense in $\mathbb{R}$.
+To conclude the proof of S2, note that, for $\mathbf{g}\in J^{c}$
+$\mathbf{h}(\mathbf{g})^{T}\nabla_{\mathbf{g}}\mathscr{l}$ $=-\left(\mathbf{g}\circ\left(\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}}{p_{\mathbf{g}}(y)}p_{G^{*}}(y)-\bf 1\right)\right)^{T}\left(\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}-\mathbf{k}_{y}(\vartheta_{d_{\eta}})\mathbf{1}}{p_{\mathbf{g}}(y)}p_{G^{*}}(y)\right)$
+$=-\left(\sum_{i=1}^{d_{\eta}}\mathbf{g}_{i}\left(\sum_{y\in\mathbb{N}_{0}}k_{y}(\vartheta_{i})\frac{p_{G^{*}}(y)}{p_{\mathbf{g}}(y)}-1\right)^{2}\right)<0.$
+Hence S2 holds, and, by (C.14)
+$\inf_{\mathbf{g}\in J}||\mathbf{g}_{n}-\mathbf{g}||\rightarrow 0,$
+$\mathds{P}_{G^{*}}$-a.s. as $n\rightarrow\infty$.
+We now prove that $\mathbf{g}_{n}$ converges $\mathds{P}$-a.s. to the minimizer $\mathbf{g}_{\eta}^{*}$ of the Kullback-Leibler divergence $KL(p_{G^{*}}||p_{g})$.
+First, notice that, since $\mathbf{g}_{n+1}-\mathbf{g}_{n}$ converges to zero $\mathds{P}$-a.s.
+and since $J$ is made of isolated points, then $\mathbf{g}_{n}$ converges $\mathds{P}$-a.s.
+to a random vector taking values in $J$.
+The minimizer $\mathbf{g}_{\eta}^{*}$ of $\mathscr{l}(\mathbf{g})$ satisfies
+$(\mathbf{g}_{\eta}^{*}-\mathbf{g})^{T}\nabla\mathscr{l}(\mathbf{g})<0$
+for every $\mathbf{g}\neq\mathbf{g}_{\eta}^{*}$. Hence
+$\sum_{i=1}^{{d_{\eta}}-1}(\mathbf{g}_{\eta,i}^{*}-\mathbf{g}_{i})\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{i})-k(y\mid\vartheta_{d_{\eta}})}{p_{g}(y)}p_{G^{*}}(y)>0,$
+which implies that
+$\sum_{i=1}^{{d_{\eta}}}\mathbf{g}_{i}^{*}\left(\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{i})}{p_{g}(y)}p_{G^{*}}(y)-1\right)>0,$
+or, equivalently
+$\sum_{i=1}^{{d_{\eta}}}(\mathbf{g}_{i}^{*}-\mathbf{g}_{i})\left(\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{i})}{p_{g}(y)}p_{G^{*}}(y)-1\right)>0.$ (C.15)
+We now show that that, for every $\delta>0$, $\mathds{P}_{G^{*}}(\liminf_{n}{\rm d}(\mathbf{g}_{n},J_{\delta})=0)=0$, where $J_{\delta}=\{\mathbf{g}\in J:d(\mathbf{g},\mathbf{g}_{\eta}^{*})>\delta\}$.
+If $\mathbf{g}\in J_{\delta}$, then there exists $\vartheta_{i}$ such that $g(\vartheta_{i})=0$ and $g_{\eta}^{*}(\vartheta_{i})\neq 0$.
+By contradiction, suppose that there exists $M$ with $\mathds{P}_{G^{*}}(M)>0$ such that ${\rm d}(\mathbf{g}_{n}(\omega),J_{\delta})\rightarrow 0$ for every $\omega\in M$, as $n\rightarrow\infty$.
+Define the function $\mathcal{L}(\mathbf{g}_{n})=\sum_{i=1}^{d_{\eta}}\log(g_{\eta}^{*}(\vartheta_{i})/g_{n}(\vartheta_{i}))g_{\eta}^{*}(\vartheta_{i})$.
 By the contradiction hypothesis, $\mathcal{L}(\mathbf{g}_{n})(\omega)\rightarrow\infty$ as $n\rightarrow\infty$, for every $\omega\in M$.
-Then $E^{*}({\mathcal{L}}(\mathbf{g}_{n}))\rightarrow\infty$ as $n\rightarrow+\infty$, where $E^{*}$ denote expectation with respect to $\mathds{P}^{*}$.
+Then $E_{G^{*}}({\mathcal{L}}(\mathbf{g}_{n}))\rightarrow\infty$ as $n\rightarrow+\infty$, where $E_{G^{*}}$ denote expectation with respect to $\mathds{P}_{G^{*}}$.
 Let $\overline{\mathbf{g}}_{n}=\eta_{n}\mathbf{g}_{n}+(1-\eta_{n})\mathbf{g}_{n+1}$, with the random variable $\eta_{n}$ taking values in $[0,1]$, be such that
@@
 $+(\mathbf{g}_{n+1}-\mathbf{g}_{n})^{T}\nabla^{2}_{\mathbf{g}}\mathcal{L}(\overline{\mathbf{g}}_{n})(\mathbf{g}_{n+1}-\mathbf{g}_{n})$
-$=\mathcal{L}(\mathbf{g}_{n})-\sum_{i=1}^{d}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})+\sum_{i=1}^{d}\frac{\mathbf{g}^{*}_{i}}{(\overline{\mathbf{g}}_{n,i})^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}.$
-Notice that
-$E^{*}\left(\sum_{i=1}^{d}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i}\right)=\alpha_{n+1}E^{*}\left(\sum_{i=1}^{d}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}(\mathbf{h}_{i}(\mathbf{g}_{n})+{\bm{\epsilon}}_{n+1,i})\right),$
+$=\mathcal{L}(\mathbf{g}_{n})-\sum_{i=1}^{d_{\eta}}\frac{\mathbf{g}_{\eta,i}^{*}}{\mathbf{g}_{n,i}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})+\sum_{i=1}^{d_{\eta}}\frac{\mathbf{g}_{\eta,i}^{*}}{(\overline{\mathbf{g}}_{n,i})^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}.$
+Denoting by $\mathcal{G}_{n}$ the sigma-algebra generated by $Y_{1},\dots,Y_{n}$, we can write that
+$E_{G^{*}}\left(\sum_{i=1}^{d_{\eta}}\frac{\mathbf{g}_{\eta,i}^{*}}{\mathbf{g}_{n,i}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})\mid\mathcal{G}_{n}\right)=\alpha_{n+1}E_{G^{*}}\left(\sum_{i=1}^{d_{\eta}}\frac{\mathbf{g}_{\eta,i}^{*}}{\mathbf{g}_{n,i}}(\mathbf{h}_{i}(\mathbf{g}_{n})+{\bm{\epsilon}}_{n+1,i})\mid\mathcal{G}_{n}\right),$
 with
-$E^{*}\left(\sum_{i=1}^{d}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}{\bm{\epsilon}}_{n+1,i}\right)=0,$
+$E_{G^{*}}\left(\sum_{i=1}^{d_{\eta}}\frac{\mathbf{g}_{\eta,i}^{*}}{\mathbf{g}_{n,i}}{\bm{\epsilon}}_{n+1,i}\mid\mathcal{G}_{n}\right)=0,$
 and
-$\liminf_{n}E^{*}\left(\sum_{i=1}^{d}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}\mathbf{h}_{i}(\mathbf{g}_{n})\right)=\liminf_{n}E^{*}\left(\sum_{i=1}^{d}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}\left(\sum_{y\in\mathbb{N}_{0}}\frac{k_{y}(\vartheta_{i})\mathbf{g}_{n,i}p_{\mathbf{g}^{*}}(y)}{p_{\mathbf{g}_{n}}(y)}-\mathbf{g}_{n,i}\right)\right)$
-$=\liminf_{n}E^{*}\left(\sum_{y\in\mathbb{N}_{0}}\frac{p_{\mathbf{g}^{*}}(y)^{2}}{p_{\mathbf{g}_{n}}(y)}-1\right)$
-$>\liminf_{n}E^{*}\left(\left(\sum_{y\in\mathbb{N}_{0}}\frac{p_{\mathbf{g}_{n}}(y)}{p_{\mathbf{g}^{*}}(y)}p_{\mathbf{g}^{*}_{i}}(y)\right)^{-1}-1\right)=0,$
-where the strict inequality holds since, under the contradiction assumption, $\mathds{P}(p_{\mathbf{g}_{n}}\not\rightarrow p_{\mathbf{g}^{*}})>0$. Thus there exists a constant $c>0$ such that
-$E^{*}\left(-\sum_{i=1}^{d}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})\right)<-c\alpha_{n+1}$
-for $n$ large enough. On the other hand, denoting by $E^{*}_{n}$ the expectation under $\mathds{P}^{*}$, conditional to $\mathcal{G}_{n}$, we can write that
-$E^{*}\left(\sum_{i=1}^{d}\frac{\mathbf{g}^{*}_{i}}{(\overline{\mathbf{g}}_{n,i})^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}\right)$
-$\leq E^{*}\left(\sum_{i=1}^{d}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}\right)+E^{*}\left(\sum_{i=1}^{d}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n+1,i}^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}\right)$
-$\leq\alpha_{n+1}^{2}E^{*}\left(\sum_{i=1}^{d}\mathbf{g}^{*}_{i}\left(\frac{k_{Y_{n+1}}(\vartheta_{i})}{p_{\mathbf{g}_{n}}(Y_{n+1})}-1\right)^{2}\right)+\alpha_{n+1}^{2}E^{*}\left(\sum_{i=1}^{d}\mathbf{g}^{*}_{i}\frac{\left(\frac{k_{Y_{n+1}}(\vartheta_{i})}{p_{\mathbf{g}_{n}}(Y_{n+1})}-1\right)^{2}}{\left(1+\alpha_{n+1}\left(\frac{k_{Y_{n+1}}(\vartheta_{i})}{p_{\mathbf{g}_{n}}(Y_{n+1})}-1\right)\right)^{2}}\right)$
-$\leq\alpha_{n+1}^{2}E^{*}\left(1+E^{*}_{n}\left(\sum_{i=1}^{d}\left(\frac{k_{Y_{n+1}}(\vartheta_{i})}{\sum_{i=1}^{d}k_{Y_{n+1}}(\vartheta_{i})\mathbf{g}_{n,i}}\right)^{2}\right)\right)$
-$\quad\quad+\frac{\alpha_{n+1}^{2}}{(1-\alpha_{n+1})^{2}}E^{*}\left(1+E^{*}_{n}\left(\sum_{i=1}^{d}\left(\frac{k_{Y_{n+1}}(\vartheta_{i})}{\sum_{i=1}^{d}k_{Y_{n+1}}(\vartheta_{i})\mathbf{g}_{n,i}}\right)^{2}\right)\right)$
-$\leq 2\frac{\alpha_{n+1}^{2}}{(1-\alpha_{n+1})^{2}}\left(1+\sum_{y\in\mathbb{N}_{0}}\sum_{i,j,k}\frac{k_{y}(\vartheta_{i})^{2}}{(\sum_{\vartheta_{k}}k_{y}(\vartheta_{k})g_{n}(\vartheta_{k}))^{2}}k_{y}(\vartheta_{j})g^{*}(\vartheta_{j})\right)$
-$\leq 2\frac{\alpha_{n+1}^{2}}{(1-\alpha_{n+1})^{2}}\left(1+\sum_{y\in\mathbb{N}_{0}}\sum_{i,j,k}\frac{k_{y}(\vartheta_{i})^{2}}{k_{y}(\vartheta_{k})^{2}}k_{y}(\vartheta_{j})g^{*}(\vartheta_{j})g_{n}(\vartheta_{k})\right)$
-$\leq 2\frac{\alpha_{n+1}^{2}}{(1-\alpha_{n+1})^{2}}\left(1+d\sum_{y\in\mathbb{N}_{0}}\ max_{i,j,k}\frac{k_{y}(\vartheta_{i})^{2}}{k_{y}(\vartheta_{k})^{2}}k_{y}(\vartheta_{j})\right)$
-$=o(\alpha_{n}),$
-where the last equality holds since
-$\sum_{y\in\mathbb{N}_{0}}\frac{k_{y}(\vartheta_{i})^{2}}{k_{y}(\vartheta_{j})^{2}}k_{y}(\vartheta_{k})=\sum_{y\in\mathbb{N}_{0}}e^{2\vartheta_{j}-2\vartheta_{i}-\vartheta_{k}}\left(\frac{\vartheta_{i}^{2}\vartheta_{k}}{\vartheta_{j}^{2}}\right)^{y}\frac{1}{y!}<\infty.$
-It follows that,
-$E^{*}\left(\mathcal{L}(\mathbf{g}_{n+1})\right)\leq E^{*}\left(\mathcal{L}(\mathbf{g}_{n})\right),$
-for $n$ large enough, which contradicts $E(\mathcal{L}(\mathbf{g}_{n}))\rightarrow\infty$. ∎
+$\liminf_{n}E_{G^{*}}\left(\sum_{i=1}^{d_{\eta}}\frac{\mathbf{g}_{\eta,i}^{*}}{\mathbf{g}_{n,i}}\mathbf{h}_{i}(\mathbf{g}_{n})\mid\mathcal{G}_{n}\right)=\liminf_{n}\sum_{i=1}^{d_{\eta}}\frac{\mathbf{g}_{\eta,i}^{*}}{\mathbf{g}_{n,i}}\left(\sum_{y\in\mathbb{N}_{0}}\frac{k_{y}(\vartheta_{i})\mathbf{g}_{n,i}p_{G^{*}}(y)}{p_{\mathbf{g}_{n}}(y)}-\mathbf{g}_{n,i}\right)$
+$=\liminf_{n}\sum_{i=1}^{d_{\eta}}\mathbf{g}_{\eta,i}^{*}\left(\sum_{y\in\mathbb{N}_{0}}\frac{k_{y}(\vartheta_{i})}{p_{\mathbf{g}_{n}}(y)}p_{G^{*}}(y)-1\right)$
+$>0,$
+by (C.15). Thus there exists a constant $c>0$ such that
+$E_{G^{*}}\left(-\sum_{i=1}^{d_{\eta}}\frac{\mathbf{g}_{\eta,i}^{*}}{\mathbf{g}_{n,i}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})\mid\mathcal{G}_{n}\right)<-c\alpha_{n+1}$
+for $n$ large enough. On the other hand,
+$E_{G^{*}}\left(\sum_{i=1}^{d_{\eta}}\frac{\mathbf{g}_{\eta,i}^{*}}{(\overline{\mathbf{g}}_{n,i})^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}\right)$
+$\leq E_{G^{*}}\left(\sum_{i=1}^{d_{\eta}}\frac{\mathbf{g}_{\eta,i}^{*}}{\mathbf{g}_{n,i}^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}\right)+E_{G^{*}}\left(\sum_{i=1}^{d_{\eta}}\frac{\mathbf{g}_{\eta,i}^{*}}{\mathbf{g}_{n+1,i}^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}\right)$
+$\leq\alpha_{n+1}^{2}E_{G^{*}}\left(\sum_{i=1}^{d_{\eta}}\mathbf{g}_{\eta,i}^{*}\left(\frac{k_{Y_{n+1}}(\vartheta_{i})}{p_{\mathbf{g}_{n}}(Y_{n+1})}-1\right)^{2}\right)+\alpha_{n+1}^{2}E_{G^{*}}\left(\sum_{i=1}^{d_{\eta}}\mathbf{g}_{\eta,i}^{*}\frac{\left(\frac{k_{Y_{n+1}}(\vartheta_{i})}{p_{\mathbf{g}_{n}}(Y_{n+1})}-1\right)^{2}}{\left(1+\alpha_{n+1}\left(\frac{k_{Y_{n+1}}(\vartheta_{i})}{p_{\mathbf{g}_{n}}(Y_{n+1})}-1\right)\right)^{2}}\right)$
+$\leq\alpha_{n+1}^{2}E_{G^{*}}\left(1+E_{G^{*}}\left(\sum_{i=1}^{d_{\eta}}\left(\frac{k_{Y_{n+1}}(\vartheta_{i})}{\sum_{i=1}^{d_{\eta}}k_{Y_{n+1}}(\vartheta_{i})\mathbf{g}_{n,i}}\right)^{2}\mid\mathcal{G}_{n}\right)\right)$
+$\quad\quad+\frac{\alpha_{n+1}^{2}}{(1-\alpha_{n+1})^{2}}E_{G^{*}}\left(1+E_{G^{*}}\left(\sum_{i=1}^{d_{\eta}}\left(\frac{k_{Y_{n+1}}(\vartheta_{i})}{\sum_{i=1}^{d_{\eta}}k_{Y_{n+1}}(\vartheta_{i})\mathbf{g}_{n,i}}\right)^{2}\mid\mathcal{G}_{n}\right)\right)$
+$\leq 2\frac{\alpha_{n+1}^{2}}{(1-\alpha_{n+1})^{2}}\left(1+\sum_{y\in\mathbb{N}_{0}}\sum_{i,k}\frac{k_{y}(\vartheta_{i})^{2}}{(\sum_{\vartheta_{k}}k_{y}(\vartheta_{k})g_{n}(\vartheta_{k}))^{2}}\int_{\Theta}k_{y}(\theta)G^{*}(\mathrm{d}\theta)\right)$
+$\leq 2\frac{\alpha_{n+1}^{2}}{(1-\alpha_{n+1})^{2}}\left(1+\sum_{i=1}^{d_{\eta}}\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{i})^{2}}{p_{g_{n}}(y)^{2}}p_{G^{*}}(y)\right)$
+$=o(\alpha_{n}).$
+where the last equality holds by (C.6), as proved in Lemma C.3. It follows that,
+$E_{G^{*}}\left(\mathcal{L}(\mathbf{g}_{n+1})\right)\leq E_{G^{*}}\left(\mathcal{L}(\mathbf{g}_{n})\right),$
+for $n$ large enough, which contradicts $E_{G^{*}}(\mathcal{L}(\mathbf{g}_{n}))\rightarrow\infty$. ∎
 Lemma C.5.
-Let $\mathbf{g}_{n}$ and $\mathbf{g}^{*}$ be the representations of $g_{n}$ ($n\geq 0$) and $g^{*}$ in $\mathbb{R}^{d}$. Under the assumptions of Theorem 2.4, for every $\delta<2-1/\gamma$
-$||\mathbf{g}_{n}-\mathbf{g}^{*}||^{2}=o(n^{-\delta})\quad\mathds{P}^{*}-a.s.$
+Let $\mathbf{g}_{n}$ and $\mathbf{g}_{\eta}^{*}$ be the representations of $g_{n}$ ($n\geq 0$) and $g_{\eta}^{*}$ in $\mathbb{R}^{d_{\eta}}$.
+Under the assumptions of Theorem 2.4 with $\gamma<1$, for every $\delta<2-1/\gamma$
+$||\mathbf{g}_{n}-\mathbf{g}_{\eta}^{*}||^{2}=o(n^{-\delta})\quad\mathds{P}_{G^{*}}-a.s.$
 Proof.
-In this proof, $\mathbf{g}_{n}$ and $\mathbf{g}^{*}$ are considered as random vectors in $\{\mathbf{v}\in\mathbb{R}^{d}:||\mathbf{v}||\leq 1\}$.
-By Lemma F.4, $\mathbf{g}_{n}$, seen as vectors in $\Delta^{d}$ converge $P^{*}$-a.s.
-to $\mathbf{g}\in\Delta^{d}$.
-This implies that $\mathbf{g}_{n}$ converges to $\mathbf{g}^{*}$ also when seen as vectors in $\mathbb{R}^{d}$.
-Moreover, $(\mathbf{g}_{n})$ satisfies the stochastic approximation in $\mathbb{R}^{d}$:
+The random vectors $\mathbf{g}_{n}$ and $\mathbf{g}_{\eta}^{*}$ take values in the compact set $\{\mathbf{v}\in\mathbb{R}^{d_{\eta}}:||\mathbf{v}||\leq 1\}$.
+By Lemma C.4, $\mathbf{g}_{n}$, seen as vectors in $\Delta^{d_{\eta}}$ converge $P_{G^{*}}$-a.s.
+to $\mathbf{g}_{\eta}^{*}\in\Delta^{d_{\eta}}$.
+This implies that $\mathbf{g}_{n}$ converges to $\mathbf{g}_{\eta}^{*}$ also when seen as vectors in $\mathbb{R}^{d_{\eta}}$.
+Moreover, $(\mathbf{g}_{n})$ satisfies the stochastic approximation in $\mathbb{R}^{d_{\eta}}$:
 $\mathbf{g}_{n+1}$ $=\mathbf{g}_{n}+\alpha_{n+1}\mathbf{g}_{n}\circ\left(\frac{\mathbf{k}_{Y_{n+1}}}{p_{\mathbf{g}_{n}}(Y_{n+1})}-\bf 1\right)$
@@
 with initial value $\mathbf{g}_{0}$, where
-$\mathbf{h}(\mathbf{g})=\mathbf{g}\circ\left(\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}}{p_{\mathbf{g}}(y)}p_{\mathbf{g}^{*}}(y)-\bf 1\right),$
+$\mathbf{h}(\mathbf{g})=\mathbf{g}\circ\left(\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}}{p_{\mathbf{g}}(y)}p_{G^{*}}(y)-\bf 1\right),$
 and
-${\bm{\epsilon}}_{n+1}=\mathbf{g}\circ\left(\frac{\mathbf{k}_{Y_{n+1}}}{p_{\mathbf{g}}(Y_{n+1})}p_{\mathbf{g}^{*}}(Y_{n+1})-\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}}{p_{\mathbf{g}}(y)}p_{\mathbf{g}^{*}}(y)\right).$
-Based on Lemma C.4, $\mathbf{g}_{n}$ converges to $\mathbf{g}^{*}$, $\mathds{P}^{*}$-a.s. The thesis follows from Theorem 3.1.1 in Chen (2005) if we can prove that the following conditions hold:
-  • B1
+${\bm{\epsilon}}_{n+1}=\mathbf{g}\circ\left(\frac{\mathbf{k}_{Y_{n+1}}}{p_{\mathbf{g}}(Y_{n+1})}-\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}}{p_{\mathbf{g}}(y)}p_{G^{*}}(y)\right).$
+The thesis follows from Theorem 3.1.1 in Chen (2005) if we can prove that the following conditions hold:
+  • T1
     $\alpha_{n}\rightarrow 0$, $\sum_{n=1}^{\infty}\alpha_{n}=\infty$, $\alpha_{n+1}^{-1}-\alpha_{n}^{-1}\rightarrow 0$;
-  • B2
-    $\sum_{n=1}^{\infty}\alpha_{n}^{1-\delta/2}\epsilon_{n}$ converges, $\mathds{P}^{*}$-a.s.;
-  • B3
-    $\mathbf{h}$ is measurable and locally bounded, and is differentiable at $\mathbf{g}^{*}$. Let $\mathbf{H}$ be a matrix such that, as $\mathbf{g}\rightarrow\mathbf{g}^{*}$,
-    $\mathbf{h}(\mathbf{g})=\mathbf{H}(\mathbf{g}-\mathbf{g}^{*})+\mathbf{r}(g),\quad\mathbf{r}(g^{*})=\mathbf{0},\quad||\mathbf{r}(\mathbf{g})||=o(||\mathbf{g}-\mathbf{g}^{*}||).$
+  • T2
+    $\sum_{n=1}^{\infty}\alpha_{n}^{1-\delta/2}\epsilon_{n}$ converges, $\mathds{P}_{G^{*}}$-a.s.;
+  • T3
+    $\mathbf{h}$ is measurable and locally bounded, and is differentiable at $\mathbf{g}_{\eta}^{*}$. Let $\mathbf{H}$ be a matrix such that, as $\mathbf{g}\rightarrow\mathbf{g}_{\eta}^{*}$,
+    $\mathbf{h}(\mathbf{g})=\mathbf{H}(\mathbf{g}-\mathbf{g}_{\eta}^{*})+\mathbf{r}(g),\quad\mathbf{r}(g_{\eta}^{*})=\mathbf{0},\quad||\mathbf{r}(\mathbf{g})||=o(||\mathbf{g}-\mathbf{g}_{\eta}^{*}||).$
     All the eigenvalues of the matrix $\mathbf{H}$ have negative real parts.
 Remark C.6.
-Theorem 3.1.1 in Chen (2005) also requests the following condition: B4 There exists a continuously differentiable function $\mathscr{l}:\Delta^{d}\rightarrow\mathbb{R}$ such that $\sup_{\delta_{1}\leq{\rm d}(\mathbf{g},\mathbf{g}^{*})\leq\delta_{2}}\mathbf{h}(\mathbf{g})^{T}\nabla_{\mathbf{g}}\mathscr{l}<0$ for every $\delta_{2}>\delta_{1}>0$.
-However, a close inspection of the proof of Theorem 3.1.1 in Chen shows that B4 is employed only to ensure convergence of $\mathbf{g}_{n}$ towards $\mathbf{g}^{*}$.
+Theorem 3.1.1 in Chen (2005) also requests the following condition: T4 There exists a continuously differentiable function $\mathscr{l}:\Delta^{d_{\eta}}\rightarrow\mathbb{R}$ such that $\sup_{\delta_{1}\leq{\rm d}(\mathbf{g},\mathbf{g}_{\eta}^{*})\leq\delta_{2}}\mathbf{h}(\mathbf{g})^{T}\nabla_{\mathbf{g}}\mathscr{l}<0$ for every $\delta_{2}>\delta_{1}>0$.
+However, a close inspection of the proof of Theorem 3.1.1 in Chen shows that T4 is employed only to ensure convergence of $\mathbf{g}_{n}$ towards $\mathbf{g}_{\eta}^{*}$.
 In our case, this property has been proved in Lemma C.4.
-B1 is obvious. To prove B2, notice that $(\sum_{k=1}^{n}\alpha_{k}^{1-\delta/2}\epsilon_{k})_{n\geq 1}$ is a martingale. Moreover, there exists a constant $C$ such that, for every $i=1,\dots,d$,
-$\sup_{n}E^{*}((\sum_{k=1}^{n}\alpha_{k}^{1-\delta/2}{\bm{\epsilon}}_{k,i})^{2})$ $=\sup_{n}\sum_{k=1}^{n}\alpha_{k}^{2-\delta}E^{*}({\bm{\epsilon}}_{k,i})^{2})$
+T1 is obvious.
+To prove T2, notice that $(\sum_{k=1}^{n}\alpha_{k}^{1-\delta/2}\epsilon_{k})_{n\geq 1}$ is a martingale.
+Moreover, there exists a constant $C$ such that, for every $i=1,\dots,{d_{\eta}}$,
+$\sup_{n}E_{G^{*}}((\sum_{k=1}^{n}\alpha_{k}^{1-\delta/2}{\bm{\epsilon}}_{k,i})^{2})$ $=\sup_{n}\sum_{k=1}^{n}\alpha_{k}^{2-\delta}E_{G^{*}}({\bm{\epsilon}}_{k,i})^{2})$
 $\leq C\sum_{k=1}^{\infty}\left(\frac{1}{\alpha+k}\right)^{\gamma(2-\delta)}<\infty,$
-since $\gamma(2-\delta)>1$ and $({\bm{\epsilon}}_{n,i})$ are uniformly bounded. We now prove B3. The function $\mathbf{h}$ is continuously differentiable at $\mathbf{g}^{*}$ and
-$\mathbf{H}_{i,j}=-\mathbf{g}^{*}_{i}\sum_{y\in\mathbb{N}_{0}}\frac{k_{y}(\vartheta_{i})k_{y}(\vartheta_{j})}{p_{g^{*}}(y)}.$
-The matrix $\mathbf{H}$ can be seen as the product of the diagonal matrix ${\rm diag}(\mathbf{g}^{*})$, which is positive definite, times $-\mathbf{M}$, with $\mathbf{M}_{i,j}=\sum_{y\in\mathbb{N}_{0}}\frac{k_{y}(\vartheta_{i})k_{y}(\vartheta_{j})}{p_{g^{*}}(y)}$.
+since $\gamma(2-\delta)>1$ and $({\bm{\epsilon}}_{n,i})$ are uniformly bounded. We now prove T3. The function $\mathbf{h}$ is continuously differentiable at $\mathbf{g}_{\eta}^{*}$ and
+$\mathbf{H}_{i,j}=-\mathbf{g}_{\eta,i}^{*}\sum_{y\in\mathbb{N}_{0}}\frac{k_{y}(\vartheta_{i})k_{y}(\vartheta_{j})}{p_{g_{\eta}^{*}}(y)^{2}}p_{G^{*}}(y).$
+The matrix $\mathbf{H}$ can be seen as the product of the diagonal matrix ${\rm diag}(\mathbf{g}_{\eta}^{*})$, which is positive definite, times $-\mathbf{M}$, with $\mathbf{M}_{i,j}=\sum_{y\in\mathbb{N}_{0}}\frac{k_{y}(\vartheta_{i})k_{y}(\vartheta_{j})}{p_{g_{\eta}^{*}}(y)^{2}}p_{G^{*}}(y)$.
 We now prove that $\mathbf{M}$ is positive definite.
-Given a vector $\mathbf{u}\in\mathbb{R}^{d}$ with $||\mathbf{u}||=1$, we can write that
-$\mathbf{u}^{T}\mathbf{M}\mathbf{u}=$ $\sum_{i,j=1}^{d}\mathbf{u}_{i}\mathbf{u}_{j}\sum_{y\in\mathbb{N}_{0}}\frac{k_{y}(\vartheta_{i})k_{y}(\vartheta_{j})}{p_{g^{*}}(y)}=\sum_{y\in\mathbb{N}_{0}}\frac{1}{p_{g^{*}}(y)}\left(\sum_{i=1}^{d}\mathbf{u}_{i}k_{y}(\vartheta_{i})\right)^{2}\geq 0,$
-with equality if and only if $\sum_{i=1}^{d}\mathbf{u}_{i}k_{y}(\vartheta_{i})=0$ for every $y\in\mathbb{N}_{0}$.
+Given a vector $\mathbf{u}\in\mathbb{R}^{d_{\eta}}$ with $||\mathbf{u}||=1$, we can write that
+$\mathbf{u}^{T}\mathbf{M}\mathbf{u}=$ $\sum_{i,j=1}^{d_{\eta}}\mathbf{u}_{i}\mathbf{u}_{j}\sum_{y\in\mathbb{N}_{0}}\frac{k_{y}(\vartheta_{i})k_{y}(\vartheta_{j})}{p_{g_{\eta}^{*}}(y)^{2}}p_{G^{*}}(y)=\sum_{y\in\mathbb{N}_{0}}\frac{p_{G^{*}{(y)}}}{p_{g_{\eta}^{*}}(y)^{2}}\left(\sum_{i=1}^{d_{\eta}}\mathbf{u}_{i}k_{y}(\vartheta_{i})\right)^{2}\geq 0,$
+with equality if and only if $\sum_{i=1}^{d_{\eta}}\mathbf{u}_{i}k_{y}(\vartheta_{i})=0$ for every $y\in\mathbb{N}_{0}$.
 The only solution to the last equation in $\mathbf{u}$ is the zero vector, which is inconsistent with $||\mathbf{u}||=1$.
@@
 ∎
+Lemma C.7.
+Under the assumptions of Theorem 2.4, there exists a constant $C=C(d,\vartheta_{1},\vartheta_{d})$ such that for every $n\geq 1$
+${\rm Regret}(g_{n},g_{\eta}^{*})\leq C||\mathbf{g}_{n}-\mathbf{g}_{\eta}^{*}||^{2},$
+where $\mathbf{g}_{n}$ and $\mathbf{g}_{\eta}^{*}$ are the representations of $g_{n}$ and $g_{\eta}^{*}$ in $\mathbb{R}^{d_{\eta}}$.
+Proof.
+We can write that
+$\sum_{y\in\mathbb{N}_{0}}(\hat{\theta}_{g_{n}}(y)-\hat{\theta}_{g_{\eta}^{*}}(y))^{2}p_{g_{\eta}^{*}}(y)$
+$=\sum_{y\in\mathbb{N}_{0}}\left(\frac{\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{y+1}e^{-\vartheta_{i}}g_{n}(\vartheta_{i})}{\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{y}e^{-\vartheta_{i}}g_{n}(\vartheta_{i})}-\frac{\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{y+1}e^{-\vartheta_{i}}g_{\eta}^{*}(\vartheta_{i})}{\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{y}e^{-\vartheta_{i}}g_{\eta}^{*}(\vartheta_{i})}\right)^{2}p_{G^{*}}(y)$
+$\leq\sum_{y\in\mathbb{N}_{0}}\frac{\left(\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{y+1}e^{-\vartheta_{i}}\left(g_{n}(\vartheta_{i})-g_{\eta}^{*}(\vartheta_{i})\right)\right)^{2}}{(\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{y}e^{-\vartheta_{i}}g_{\eta}^{*}(\vartheta_{i}))^{2}}p_{G^{*}}(y)$
+$+\sum_{y\in\mathbb{N}_{0}}\frac{\left(\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{y+1}e^{-\vartheta_{i}}g_{n}(\vartheta_{i})\right)^{2}\left(\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{y}e^{-\vartheta_{i}}(g_{n}(\vartheta_{i})-g_{\eta}^{*}(\vartheta_{i})\right)^{2}}{\left(\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{y}e^{-\vartheta_{i}}g_{n}(\vartheta_{i})\right)^{2}\left(\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{y}e^{-\vartheta_{i}}g_{\eta}^{*}(\vartheta_{i})\right)^{2}}p_{G^{*}}(y)$
+$\leq 2\sum_{y\in\mathbb{N}_{0}}\frac{\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{2y+2}e^{-2\vartheta_{i}}||\mathbf{g}_{n}-\mathbf{g}_{\eta}^{*}||^{2}}{\left(\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{y}e^{-\vartheta_{i}}g_{\eta}^{*}(\vartheta_{i})\right)^{2}}p_{G^{*}}(y)+2\sum_{y\in\mathbb{N}_{0}}\frac{\vartheta_{d_{\eta}}^{2}\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{2y}e^{-2\vartheta_{i}}||\mathbf{g}_{n}-\mathbf{g}_{\eta}^{*}||^{2}}{\left(\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{y}e^{-\vartheta_{i}}g_{\eta}^{*}(\vartheta_{i})\right)^{2}}p_{G^{*}}(y)$
+$\leq 4{d_{\eta}}\vartheta_{d_{\eta}}^{2}e^{2(\vartheta_{d_{\eta}}-\vartheta_{1})}||\mathbf{g}_{n}-\mathbf{g}_{\eta}^{*}||^{2}\int_{0}^{\infty}e^{-\theta}\sum_{y\in\mathbb{N}_{0}}\frac{1}{y!}\left(\frac{\theta\vartheta_{d_{\eta}}^{2}}{\vartheta_{1}^{2}}\right)^{y}G^{*}(d\theta)$
+$\leq 4{d_{\eta}}\vartheta_{d_{\eta}}^{2}e^{2(\vartheta_{d_{\eta}}-\vartheta_{1})}||\mathbf{g}_{n}-\mathbf{g}_{\eta}^{*}||^{2}\int_{0}^{\infty}e^{\theta({d_{\eta}}^{2}-1)}G^{*}(d\theta)$
+$\leq C||\mathbf{g}_{n}-\mathbf{g}_{\eta}^{*}||^{2},$
+with $C=4{d_{\eta}}\vartheta_{d_{\eta}}^{2}e^{2(\vartheta_{d_{\eta}}-\vartheta_{1})}\int_{0}^{\infty}e^{\theta({d_{\eta}}^{2}-1)}G^{*}(d\theta)<+\infty,$ by (C.8). Thus,
+${\rm Regret}(g_{n},g_{\eta}^{*})\leq C||\mathbf{g}_{n}-\mathbf{g}_{\eta}^{*}||^{2},$
+with $C<+\infty$. ∎
 of Theorem 2.4.
-Denoting by $(\mathbf{g}_{n})$ and $\mathbf{g}^{*}$ the representations of $(g_{n})$ and $g^{*}$ in $\mathbb{R}^{d}$, we can write, by Lemma C.3 and Lemma C.5,
-${\rm Regret}(g_{n},g^{*})\leq C||\mathbf{g}_{n}-\mathbf{g}^{*}||^{2}=o(n^{-\delta}).$
+Let $\mathbf{g}_{n}$ and $\mathbf{g}_{\eta}^{*}$ denote the representation of $g_{n}$ and $g_{\eta}^{*}$ in $\Delta_{\eta}=\{\mathbf{v}=(\mathbf{v}_{1},\dots,\mathbf{v}_{d_{\eta}-1}):\mathbf{v}_{i}\geq 0,\sum_{i=1}^{d_{\eta}-1}\mathbf{v}_{i}\leq 1\}$.
+The sequence $(\mathbf{g}_{n})$ satisfies the stochastic approximation (C.13), with $\mathbf{g}_{0}\in\mathring{(\Delta_{\eta})}$.
+By Lemma C.4, $\mathbf{g}_{n}$ converges to $\mathbf{g}_{\eta}^{*}$, $\mathds{P}^{*}$-a.s.
+It follows that $g_{n}(\vartheta_{j})$ converges to $g_{\eta}^{*}(\vartheta_{j})$ $\mathds{P}^{*}$-a.s.
+for every $\vartheta_{j}\in\Theta_{\eta}$, which imples that $\hat{\theta}_{g_{n}}(y)$ converges to $\hat{\theta}_{g^{*}_{\eta}}(y)$ for every $y\in\mathbb{N}_{0}$.
+Now assume that $\gamma<1$. Denoting by $\mathbf{g}_{n}$ and $\mathbf{g}_{\eta}^{*}$ the representations of $g_{n}$ and $g_{\eta}^{*}$ in $\mathbb{R}^{d_{\eta}}$, we can write, by Lemma C.5 that
+$||\mathbf{g}_{n}-\mathbf{g}_{\eta}^{*}||^{2}=o(n^{-\delta}),$
+for every $\delta<2-1/\gamma$. By Lemma C.7,
+${\rm Regret}(g_{n},g_{\eta}^{*})\leq C||\mathbf{g}_{n}-\mathbf{g}_{\eta}^{*}||^{2}=o(n^{-\delta}).$
+$\square$
 D Numerical illustrations
 D.1 Synthetic data
-We set $n=50,100,200,500$, and generate $(Y_{1},\ldots,Y_{n})$ under the Poisson compound model, with $(\theta_{1},\ldots,\theta_{n})$ i.i.d from a Uniform prior $G$ on $[0,3]$.
-For the NP-QB method, we set Newton’s algorithm as follows: i) the parameter $\theta$ takes value in $\Theta=\{0.01,0.02,\ldots,99.99,100\}$; ii) the number $d$ of mixture components is the cardinality of $\Theta$, i.e.
-$d=10,000$; iii) the initial guess $g_{0}$ is the Uniform distribution on $\Theta$; iv) the learning rate is $\alpha_{n}=(1+n)^{-1}$.
-Results are displayed in Figure 2 and Table 3.
 […]
-Figure 2: Data $Y_{i}$’s against the $\theta_{i}$’s (black circle), and estimates of the $\theta_{i}$’s: NP-EB (green line), NP-ML (cyan line), NP-MD (magenta line), P-EB (red line) and NP-QB (blue line).
-Table 3: Values of RMSE and MAD for the NP-EB, NP-ML, NP-MD, P-EB and NP-QB estimates.
-NP-EB NP-ML NP-MD P-EB NP-QB
-RMSE   $n=100$ 1.01 0.75 1.31 0.75 0.75
-RMSE   $n=500$ 0.95 0.75 0.74 0.75 0.78
-MAD     $n=100$ 0.77 0.60 1.05 0.61 0.60
-MAD     $n=500$ 0.65 0.59 0.59 0.60 0.61
-Figure 3: Data $Y_{i}$’s against the $\theta_{i}$’s (black circle), and estimates of the $\theta_{i}$’s: NP-EB (green line), NP-ML (cyan line), NP-MD (magenta line), P-EB (red line) and NP-QB (blue line).
-Table 4: Values of RMSE and MAD for the NP-EB, NP-ML, NP-MD, P-EB and NP-QB estimates.
-NP-EB NP-ML NP-MD P-EB NP-QB
-RMSE   $n=100$ 5.83 1.99 1.71 1.83 1.75
-RMSE   $n=500$ 2.63 1.89 1.86 1.93 1.86
-MAD     $n=100$ 3.40 1.52 1.32 1.38 1.34
-MAD     $n=500$ 1.75 1.42 1.42 1.47 1.43
-D.2 Real data: Tweeter dataset
-We consider a dataset of tweets with at least 50 retweets on Twitter in the period October 7 to November 6, 2011 (30 days).
-The dataset is available at the link https://snap.stanford.edu/seismic/, and it consists of $166,783$ tweets that generate $34,617,704$ retweets.
-More specifically, the data provides information on the tweet ID, the tweet post time, the retweet post time and the number of followers of the poster/retweeter.
-For each tweet, we focus on the number of retweets generated within $30$ seconds from the tweet post time.
+Table 3: Weibull prior: grid $\Theta_{n,\eta}$ as $n$ and $\eta$ vary.
+$\eta=0.025$ $\eta=0.05$ $\eta=0.075$ $\eta=0.1$
+$n=50$ $d_{\eta}=454,169$ $d_{\eta}=104,053$ $d_{\eta}=43,755$ $d_{\eta}=23,612$
+$\vartheta_{1}=0.025$ $\vartheta_{1}=0.05$ $\vartheta_{1}=0.075$ $\vartheta_{1}=0.1$
+$\vartheta_{d_{\eta}}=11,354.225$ $\vartheta_{d_{\eta}}=52,02.65$ $\vartheta_{d_{\eta}}=3,281.625$ $\vartheta_{d_{\eta}}=2,361.2$
+$n=100$ $d_{\eta}=369,148$ $d_{\eta}=84,375$ $d_{\eta}=35,423$ $d_{\eta}=19,090$
+$\vartheta_{1}=0.025$ $\vartheta_{1}=0.05$ $\vartheta_{1}=0.075$ $\vartheta_{1}=0.1$
+$\vartheta_{d_{\eta}}=9,228.7$ $\vartheta_{d_{\eta}}=4,218.75$ $\vartheta_{d_{\eta}}=2,656.725$ $\vartheta_{d_{\eta}}=1,909$
+$n=200$ $d_{\eta}=450,573$ $d_{\eta}=103,220$ $d_{\eta}=43,402$ $d_{\eta}=23,420$
+$\vartheta_{1}=0.025$ $\vartheta_{1}=0.05$ $\vartheta_{1}=0.075$ $\vartheta_{1}=0.1$
+$\vartheta_{d_{\eta}}=11,264.325$ $\vartheta_{d_{\eta}}=5,161$ $\vartheta_{d_{\eta}}=3,255.15$ $\vartheta_{d_{\eta}}=2,342$
+$n=500$ $d_{\eta}=431,768$ $d_{\eta}=98,865$ $d_{\eta}=41,558$ $d_{\eta}=22,419$
+$\vartheta_{1}=0.025$ $\vartheta_{1}=0.05$ $\vartheta_{1}=0.075$ $\vartheta_{1}=0.1$
+$\vartheta_{d_{\eta}}=10,794.2$ $\vartheta_{d_{\eta}}=49,43.25$ $\vartheta_{d_{\eta}}=31,16.85$ $\vartheta_{d_{\eta}}=2,241.9$
+Table 4: Weibull prior: RMSE and MAD on the grid $\Theta_{n,\eta}$ as $n$ and $\eta$ vary.
+$\eta=0.025$ $\eta=0.05$ $\eta=0.075$ $\eta=0.1$
+$n=50$
+RMSE 1.572 1.575 1.578 1.581
+MAD 1.319 1.322 1.325 1.328
+$n=100$
+RMSE 1.677 1.72 1.75 1.772
+MAD 1.335 1.366 1.386 1.4
+$n=200$
+RMSE 1.335 1.336 1.336 1.337
+MAD 1.064 1.064 1.065 1.065
+$n=500$
+RMSE 1.348 1.348 1.349 1.349
+MAD 1.071 1.071 1.071 1.071
+[…]
+Table 5: Weibull prior: RMSE, MAD and CPU time (in seconds) on the grid $\Theta_{n,d}$ as $n$ and $d$ vary.
+$d=50,000$ $d=10,000$ $d=5,000$ $d=1,000$ $d=500$
+$n=50$
+RMSE 1.574 1.581 1.582 6.13 14.49
+MAD 1.323 1.328 1.329 5.706 12.518
+CPU time 1.243 0.24 0.12 0.025 0.013
+$n=100$
+RMSE 1.667 1.675 1.677 4.496 11.476
+MAD 1.335 1.335 1.335 4.101 10.536
+CPU time 1.208 0.24 0.12 0.024 0.013
+$n=200$
+RMSE 1.33 1.336 1.336 6.039 13.984
+MAD 1.062 1.064 1.065 5.733 11.949
+CPU time 1.231 0.242 0.122 0.025 0.013
+$n=500$
+RMSE 1.348 1.348 1.348 5.869 12.96
+MAD 1.071 1.071 1.072 5.553 11.152
+CPU time 1.214 0.241 0.12 0.025 0.013
+Table 6 compares the RMSE and MAD for NP-EB, NP-ML, NP-MD, P-EB and QB-EB.
+The QB-EB implements a Newton’s algorithm is initialized as follows: i) the parameter $\theta$ takes values on a grid $\Theta_{n,d}=\{\vartheta_{1},\ldots,\vartheta_{d}\}$, with $\vartheta_{1}$ and $\vartheta_{d}$ being the minimum and maximum value of $\Theta_{n,0.025}$ in Table 3, respectively, and $d=10,000$; ii) the initial guess $g_{0}$ is the Uniform distribution on $\Theta_{n,d}$; iii) the learning rate is $\alpha_{n}=(1+n)^{-0.99}$.
+This is precisely the initialization in the second column of Table 5, which also provides CPU time.
+Table 6: Weibull prior: RMSE and MAD for NP-EB, NP-ML, NP-MD, P-EB and QB-EB.
+NP-EB NP-ML NP-MD P-EB QB-EB
+RMSE   $n=50$ 3.154 1.542 1.559 1.554 1.581
+RMSE   $n=100$ 2.516 1.695 1.648 1.478 1.675
+RMSE   $n=200$ 3.236 1.273 1.248 1.262 1.336
+RMSE   $n=500$ 2.108 1.341 1.515 1.353 1.348
+MAD     $n=50$ 2.32 1.262 1.258 1.295 1.328
+MAD     $n=100$ 1.807 1.334 1.3 1.179 1.335
+MAD     $n=200$ 1.917 0.999 1.001 1.014 1.064
+MAD     $n=500$ 1.4 1.068 1.197 1.073 1.071
+[…]
+Table 7: Uniform prior: grid $\Theta_{n,\eta}$ as $n$ and $\eta$ vary.
+$\eta=0.025$ $\eta=0.05$ $\eta=0.075$ $\eta=0.1$
+$n=50$ $d_{\eta}=71,674$ $d_{\eta}=15,988$ $d_{\eta}=6,596$ $d_{\eta}=3,504$
+$\vartheta_{1}=0.025$ $\vartheta_{1}=0.05$ $\vartheta_{1}=0.075$ $\vartheta_{1}=0.1$
+$\vartheta_{d_{\eta}}=1,791.85$ $\vartheta_{d_{\eta}}=799.4$ $\vartheta_{d_{\eta}}=494.7$ $\vartheta_{d_{\eta}}=350.4$
+$n=100$ $d_{\eta}=67,264$ $d_{\eta}=14,986$ $d_{\eta}=6,177$ $d_{\eta}=3,279$
+$\vartheta_{1}=0.025$ $\vartheta_{1}=0.05$ $\vartheta_{1}=0.075$ $\vartheta_{1}=0.1$
+$\vartheta_{d_{\eta}}=1,681.6$ $\vartheta_{d_{\eta}}=749.3$ $\vartheta_{d_{\eta}}=463.275$ $\vartheta_{d_{\eta}}=327.9$
+$n=200$ $d_{\eta}=52,298$ $d_{\eta}=11,593$ $d_{\eta}=4,760$ $d_{\eta}=2,519$
+$\vartheta_{1}=0.025$ $\vartheta_{1}=0.05$ $\vartheta_{1}=0.075$ $\vartheta_{1}=0.1$
+$\vartheta_{d_{\eta}}=1,307.45$ $\vartheta_{d_{\eta}}=579.65$ $\vartheta_{d_{\eta}}=357$ $\vartheta_{d_{\eta}}=251.9$
+$n=500$ $d_{\eta}=44,591$ $d_{\eta}=9,850$ $d_{\eta}=4,034$ $d_{\eta}=2,130$
+$\vartheta_{1}=0.025$ $\vartheta_{1}=0.05$ $\vartheta_{1}=0.075$ $\vartheta_{1}=0.1$
+$\vartheta_{d_{\eta}}=1,114.775$ $\vartheta_{d_{\eta}}=492.5$ $\vartheta_{d_{\eta}}=302.55$ $\vartheta_{d_{\eta}}=213$
+Table 8: Uniform prior: RMSE and MAD on the grid $\Theta_{n,\eta}$ as $n$ and $\eta$ vary.
+$\eta=0.025$ $\eta=0.05$ $\eta=0.075$ $\eta=0.1$
+$n=50$
+RMSE 0.724 0.725 0.727 0.728
+MAD 0.615 0.616 0.617 0.618
+$n=100$
+RMSE 0.798 0.798 0.799 0.8
+MAD 0.648 0.65 0.651 0.653
+$n=200$
+RMSE 0.728 0.729 0.73 0.731
+MAD 0.619 0.62 0.621 0.622
+$n=500$
+RMSE 0.72 0.72 0.721 0.721
+MAD 0.59 0.59 0.59 0.59
+[…]
+Table 9: Uniform prior: RMSE, MAD and CPU time (in seconds) on the grid $\Theta_{n,d}$ as $n$ and $d$ vary.
+$d=50,000$ $d=10,000$ $d=5,000$ $d=1,000$ $d=500$
+$n=50$
+RMSE 0.724 0.724 0.724 0.737 1.587
+MAD 0.615 0.615 0.616 0.619 1.397
+CPU time 1.233 0.241 0.125 0.026 0.014
+$n=100$
+RMSE 0.799 0.804 0.812 0.841 1.514
+MAD 0.649 0.653 0.66 0.678 1.302
+CPU time 1.310 0.241 0.120 0.026 0.014
+$n=200$
+RMSE 0.728 0.73 0.731 0.751 1.056
+MAD 0.619 0.619 0.619 0.623 0.831
+CPU time 1.252 0.239 0.131 0.025 0.013
+$n=500$
+RMSE 0.72 0.719 0.719 0.721 0.854
+MAD 0.59 0.612 0.612 0.64 0.665
+CPU time 1.207 0.241 0.119 0.025 0.013
+Table 10 compares the RMSE and MAD for NP-EB, NP-ML, NP-MD, P-EB and QB-EB.
+The QB-EB implements a Newton’s algorithm is initialized as follows: i) the parameter $\theta$ takes values on a grid $\Theta_{n,d}=\{\vartheta_{1},\ldots,\vartheta_{d}\}$, with $\vartheta_{1}$ and $\vartheta_{d}$ being the minimum and maximum value of $\Theta_{n,0.025}$ in Table 7, respectively, and $d=10,000$; ii) the initial guess $g_{0}$ is the Uniform distribution on $\Theta_{n,d}$; iii) the learning rate is $\alpha_{n}=(1+n)^{-0.99}$.
+This is precisely the initialization considered in the second column of Table 9, which also provides the CPU time.
+Figure 2 displays the NP-EB, NP-ML, NP-MD, P-EB and QB-EB estimates.
+Table 10: Uniform prior: RMSE and MAD for NP-EB, NP-ML, NP-MD, P-EB and QB-EB.
+NP-EB NP-ML NP-MD P-EB QB-EB
+RMSE   $n=50$ 1.354 0.727 0.702 0.703 0.724
+RMSE   $n=100$ 1.253 0.818 0.841 0.763 0.804
+RMSE   $n=200$ 0.88 0.718 0.729 0.726 0.73
+RMSE   $n=500$ 0.88 0.725 0.746 0.72 0.719
+MAD     $n=50$ 0.996 0.608 0.562 0.581 0.615
+MAD     $n=100$ 0.998 0.673 0.685 0.628 0.653
+MAD     $n=200$ 0.742 0.613 0.605 0.623 0.619
+MAD     $n=500$ 0.719 0.606 0.633 0.607 0.612
+Figure 2: Data $Y_{i}$’s (Uniform prior) against the $\theta_{i}$’s (black circle), and estimates of the $\theta_{i}$’s: NP-EB (green line), NP-ML (cyan line), NP-MD (magenta line), P-EB (red line) and QB-EB (blue line).
+[…]
+Table 11: Half-Gaussian prior: grid $\Theta_{\eta}$
+$\eta=0.025$ $\eta=0.05$ $\eta=0.075$ $\eta=0.1$
+$n=50$ $d_{\eta}=29,100$ $d_{\eta}=6,362$ $d_{\eta}=2,585$ $d_{\eta}=1,355$
+$\vartheta_{1}=0.025$ $\vartheta_{1}=0.05$ $\vartheta_{1}=0.075$ $\vartheta_{1}=0.1$
+$\vartheta_{d_{\eta}}=727.5$ $\vartheta_{d_{\eta}}=318.1$ $\vartheta_{d_{\eta}}=193.875$ $\vartheta_{d_{\eta}}=135.5$
+$n=100$ $d_{\eta}=19,629$ $d_{\eta}=4,243$ $d_{\eta}=1,709$ $d_{\eta}=889$
+$\vartheta_{1}=0.025$ $\vartheta_{1}=0.05$ $\vartheta_{1}=0.075$ $\vartheta_{1}=0.1$
+$\vartheta_{d_{\eta}}=490.725$ $\vartheta_{d_{\eta}}=212.15$ $\vartheta_{d_{\eta}}=128.175$ $\vartheta_{d_{\eta}}=88.9$
+$n=200$ $d_{\eta}=16,878$ $d_{\eta}=3,631$ $d_{\eta}=1,456$ $d_{\eta}=755$
+$\vartheta_{1}=0.025$ $\vartheta_{1}=0.05$ $\vartheta_{1}=0.075$ $\vartheta_{1}=0.1$
+$\vartheta_{d_{\eta}}=421.95$ $\vartheta_{d_{\eta}}=181.55$ $\vartheta_{d_{\eta}}=109.2$ $\vartheta_{d_{\eta}}=75.5$
+$n=500$ $d_{\eta}=17,307$ $d_{\eta}=3,727$ $d_{\eta}=1,496$ $d_{\eta}=776$
+$\vartheta_{1}=0.025$ $\vartheta_{1}=0.05$ $\vartheta_{1}=0.075$ $\vartheta_{1}=0.1$
+$\vartheta_{d_{\eta}}=432.675$ $\vartheta_{d_{\eta}}=186.35$ $\vartheta_{d_{\eta}}=112.2$ $\vartheta_{d_{\eta}}=77.6$
+Table 12: Half-Gaussian prior: RMSE and MAD on the grid $\Theta_{n,\eta}$ as $n$ and $\eta$ vary.
+$\eta=0.025$ $\eta=0.05$ $\eta=0.075$ $\eta=0.1$
+$n=50$
+RMSE 0.569 0.57 0.571 0.572
+MAD 0.462 0.463 0.464 0.465
+$n=100$
+RMSE 0.633 0.633 0.627 0.624
+MAD 0.532 0.535 0.538 0.541
+$n=200$
+RMSE 0.471 0.471 0.472 0.472
+MAD 0.372 0.373 0.374 0.375
+$n=500$
+RMSE 0.536 0.537 0.537 0.538
+MAD 0.425 0.425 0.426 0.427
+[…]
+Table 13: Half-Gaussian prior: RMSE, MAD and CPU time (in seconds) on the grid $\Theta_{n,d}$ as $n$ and $d$ vary.
+$d=50,000$ $d=10,000$ $d=5,000$ $d=1,000$ $d=500$
+$n=50$
+RMSE 0.569 0.57 0.57 0.58 0.632
+MAD 0.461 0.462 0.462 0.463 0.489
+CPU time 1.234 0.242 0.134 0.027 0.013
+$n=100$
+RMSE 0.633 0.633 0.633 0.635 0.687
+MAD 0.541 0.541 0.542 0.543 0.594
+CPU time 1.226 0.241 0.12 0.024 0.013
+$n=200$
+RMSE 0.471 0.471 0.471 0.476 0.486
+MAD 0.372 0.372 0.372 0.375 0.385
+CPU time 1.208 0.241 0.118 0.024 0.013
+$n=500$
+RMSE 0.535 0.535 0.535 0.536 0.536
+MAD 0.419 0.42 0.424 0.424 0.425
+CPU time 1.223 0.241 0.12 0.025 0.013
+Table 14 compares the RMSE and MAD for NP-EB, NP-ML, NP-MD, P-EB and QB-EB.
+The QB-EB implements a Newton’s algorithm is initialized as follows: i) the parameter $\theta$ takes values on a grid $\Theta_{n,d}=\{\vartheta_{1},\ldots,\vartheta_{d}\}$, with $\vartheta_{1}$ and $\vartheta_{d}$ being the minimum and maximum value of $\Theta_{n,0.025}$ in Table 11, respectively, and $d=10,000$; ii) the initial guess $g_{0}$ is the Uniform distribution on $\Theta_{n,d}$; iii) the learning rate is $\alpha_{n}=(1+n)^{-0.99}$.
+This is precisely the initialization considered in the second column of Table 13, which also provides the CPU time.
+Figure 2 displays the NP-EB, NP-ML, NP-MD, P-EB and QB-EB estimates.
+Table 14: Half-Gaussian prior: RMSE and MAD for NP-EB, NP-ML, NP-MD, P-EB and QB-EB.
+NP-EB NP-ML NP-MD P-EB QB-EB
+RMSE   $n=50$ 1.129 0.623 0.643 0.578 0.57
+RMSE   $n=100$ 0.679 0.511 0.484 0.493 0.633
+RMSE   $n=200$ 0.611 0.477 0.491 0.483 0.471
+RMSE   $n=500$ 0.838 0.555 0.564 0.54 0.535
+MAD     $n=50$ 0.773 0.49 0.493 0.467 0.462
+MAD     $n=100$ 0.512 0.428 0.391 0.411 0.541
+MAD     $n=200$ 0.467 0.374 0.403 0.388 0.372
+MAD     $n=500$ 0.508 0.429 0.44 0.43 0.42
+Figure 3: ata $Y_{i}$’s (half-Gaussian prior) against the $\theta_{i}$’s (black circle), and estimates of the $\theta_{i}$’s: NP-EB (green line), NP-ML (cyan line), NP-MD (magenta line), P-EB (red line) and QB-EB (blue line).
+D.2 Real data: Twitter dataset
+We analyze a dataset of tweets that received at least $50$ retweets on Twitter between October 7 and November 6, 2011, available at https://snap.stanford.edu/seismic/.
+It contains $166,783$ tweets generating $34,617,704$ retweets, along with metadata such as tweet ID, posting time, retweet time, and follower counts.
+For each tweet, we focus on the number of retweets received within the first $30$ seconds after posting.
+For each tweet, we consider the number of retweets generated within $30$ seconds from the tweet post time.
 From the first two rows of Table 15, $40,259$ tweets have 00 retweets in $30$ seconds, $28,339$ tweets have $1$ retweet in $30$ seconds, etc.
@@
 P-EB 0.45 1.34 2.23 3.12 4.01 4.90 5.79 6.68
-NP-QB 0.61 0.82 1.63 3.08 4.01 4.52 5.02 5.71
+QB-EB 0.61 0.82 1.63 3.08 4.01 4.52 5.02 5.71
 $y$ 8 11 14 17 20 23 26 29
@@
 P-EB 7.57 10.24 12.91 15.58 18.25 20.93 23.60 26.27
-NP-QB 6.67 10.37 12.69 14.64 17.55 21.38 24.78 27.53
+QB-EB 6.67 10.37 12.69 14.64 17.55 21.38 24.78 27.53
 $y$ 32 35 38 41 44 47 50 53
@@
 P-EB 28.94 31.61 34.28 36.96 39.63 42.30 44.97 47.64
-NP-QB 30.26 33.32 36.59 39.71 42.50 45.05 47.54 50.09
+QB-EB 30.26 33.32 36.59 39.71 42.50 45.05 47.54 50.09
 $y$ 56 59 62 65 68 71 74 77
@@
 P-EB 50.31 52.99 55.66 58.33 61.00 63.67 66.35 69.02
-NP-QB 52.74 55.40 57.94 60.26 62.33 64.18 65.87 67.45
+QB-EB 52.74 55.40 57.94 60.26 62.33 64.18 65.87 67.45
 Figure 4: Estimates of the number of retweets expected for a tweet with $y$ retweets (in 30 seconds): NP-EB (green line), NP-ML (cyan line), NP-MD (magenta line), P-EB (red line) and NP-QB (blue line).
@@
 P-EB 0.56 30.410 60.26 90.12 119.97 149.82 179.67 209.52
-NP-QB 0.54 31.41 60.90 91.19 120.68 150.89 180.31 210.85
+QB-EB 0.54 31.41 60.90 91.19 120.68 150.89 180.31 210.85
 $y$ 240 270 300 330 360 390 420 450
@@
 P-EB 239.38 269.23 299.08 328.93 358.79 388.64 418.49 448.34
-NP-QB 240.40 271.03 300.57 330.71 360.28 390.42 418.99 442.98
+QB-EB 240.40 271.03 300.57 330.71 360.28 390.42 418.99 442.98
 Figure 5: Estimates of the number of followers expected to be generated by a tweet whose retweets have $y$ followers (in 30 seconds): NP-EB (green line), NP-ML (cyan line), NP-MD (magenta line), P-EB (red line) and NP-QB (blue line).
@@
 P-FB 0.16 0.42 0.68 0.94 1.20 1.47 1.75 2.03
-NP-QB 0.12 0.25 0.47 0.91 1.85 3.17 4.33 5.22
+QB-EB 0.12 0.25 0.47 0.91 1.85 3.17 4.33 5.22
 E An extension to the multidimensional Poisson compound decision problem
@@
 $\hat{\bm{\theta}}_{g_{n},j}(\mathbf{y})\rightarrow\hat{\bm{\theta}}_{\tilde{g},j}(\mathbf{y})=(\mathbf{y}_{j}+1)\frac{p_{\tilde{g}}(\mathbf{y}+\mathbf{e}_{j})}{p_{\tilde{g}}(\mathbf{y})}\qquad\mathds{P}\,\text{-}\,a.s.$ (E.22)
-The next result extends Theorem 2.2 to the multidimensional case.
+The next result extends Theorem 2.3 in two directions: the observations are now random vectors, and the sequence of learning rates $(\alpha_{n})$ is no longer required to be increasing.
 Theorem E.1.
@@
 Theorem 2.3 provides a tool to obtain asymptotic credible regions for $\hat{\bm{\theta}}_{g}(\mathbf{y})$, for $y\in\mathbb{N}_{0}^{k}$.
-We now consider the Bayesian regret
+We now consider the Bayesian regret under a “true” Poisson compound model for $(\mathbf{Y}_{n})_{n\geq 1}$, under which, for any $n\geq 1$: i) the distribution of $\mathbf{Y}_{n}$ given $\bm{\theta}^{\ast}_{n}$ is a product of Poisson distributions with parameters $\theta^{\ast}_{n,j}$; ii) the $\bm{\theta}^{\ast}_{n}$’s are independent and identically distributed from the “true” prior distribution, say the oracle prior $G^{\ast}$.
+By the same techniques used in the univariate case, the frequentist guarantees can be studied in this more general setting.
+We give a complete treatment under the simplifying assumption that the “true” probability distribution $G^{*}$ has support $\Theta_{d}^{k}$.
+Denoting by $g^{*}$ its probability mass function on $\Theta_{d}^{k}$, the regret becomes
 $\text{Regret}(g_{n},g^{*})=\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}||\hat{\bm{\theta}}_{g_{n}}(\mathbf{y})-\hat{\bm{\theta}}_{g^{*}}(\mathbf{y})||^{2}p_{g^{*}}(\mathbf{y}).$ (E.25)
-under a “true” Poisson compound model for $(\mathbf{Y}_{n})_{n\geq 1}$, under which, for any $n\geq 1$: i) the distribution of $\mathbf{Y}_{n}$ given $\bm{\theta}^{\ast}_{n}$ is a product of Poisson distributions with parameters $\theta^{\ast}_{n,j}$; ii) the $\bm{\theta}^{\ast}_{n}$’s are independent and identically distributed from the “true” prior distribution, say the oracle prior $G^{\ast}$ on $\Theta_{d}^{k}$, with probability mass function $g^{\ast}$ on $\Theta_{d}^{k}$.
 The asymptotic behavior of regret for large $n$ in the multidimensional case follows the same rate of convergence as in the univariate case.
@@
 Let $g_{n}$ be as in (E.19) with $\alpha_{n}=(\alpha+n)^{-\gamma}$ for $\alpha>0$ and $\gamma\in(1/2,1)$, and let $g^{\ast}$ be the probability mass function of the oracle prior $G^{\ast}$.
-Denote by $\mathds{P}^{*}$ a probability measure under which the “true” Poisson compound model for $(\mathbf{Y}_{n})_{n\geq 1}$ holds true.
+Denote by $\mathds{P}_{G^{*}}$ a probability measure under which the “true” Poisson compound model for $(\mathbf{Y}_{n})_{n\geq 1}$ holds true.
 Then, for every $\delta<2-1/\gamma$, it holds that
-$\text{Regret}(g_{n},g^{*})=o(n^{-\delta})\quad\mathds{P}^{\ast}\,\text{-}\,a.s.$
+$\text{Regret}(g_{n},g^{*})=o(n^{-\delta})\quad\mathds{P}_{G^{*}}\,\text{-}\,a.s.$
 See Supplementary Material F for the proof of Theorem E.2.
@@
 ${\bm{\epsilon}}_{n+1}=\mathbf{g}_{n}\circ\left(\frac{k_{\mathbf{Y}_{n+1}}}{p_{\mathbf{g}_{n}}(\mathbf{Y}_{n+1})}-\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{k_{\mathbf{y}}}{p_{\hat{\mathbf{g}}_{n}}(\mathbf{y})}p_{\mathbf{g}^{*}}(\mathbf{y})\right),$
-and initial state $\mathbf{g}_{0}\in\mathring{(\Delta^{D})}$, converges $\mathds{P}^{*}$-a.s. to $\mathbf{g}^{*}$ as $n\rightarrow\infty$.
+and initial state $\mathbf{g}_{0}\in\mathring{(\Delta^{D})}$, converges $\mathds{P}_{G^{*}}$-a.s. to $\mathbf{g}^{*}$ as $n\rightarrow\infty$.
 Proof.
 Let $J=\partial\Delta^{D}\cup\{\mathbf{g}^{*}\}$.
-We first prove that ${\rm d}(\mathbf{g}_{n},J)\rightarrow 0$, $\mathds{P}^{*}$-a.s., as $n\rightarrow\infty$, where ${\rm d}$ denotes the Euclidean distance in $\mathbb{R}^{D-1}$.
+We first prove that ${\rm d}(\mathbf{g}_{n},J)\rightarrow 0$, $\mathds{P}_{G^{*}}$-a.s., as $n\rightarrow\infty$, where ${\rm d}$ denotes the Euclidean distance in $\mathbb{R}^{D-1}$.
 The result follows from Theorem 2.2.1′′ in Chen (2005) if we can prove that
-  • A1
+  • S1
     $\alpha_{n}>0$, $\alpha_{n}\rightarrow 0$, as $n\rightarrow\infty$ and $\sum_{n=1}^{\infty}\alpha_{n}=\infty$,
-  • A2
+  • S2
     $J$ is closed and there exists a is continuously differentiable function $\mathscr{l}:\Delta^{D}\rightarrow\mathbb{R}$ such that $\sup_{\delta_{1}\leq{\rm d}(\mathbf{g},J)\leq\delta_{2}}\mathbf{h}(\mathbf{g})^{T}\nabla_{\mathbf{g}}\mathscr{l}<0$ for every $\delta_{2}>\delta_{1}>0$,
-  • A3
-    $\lim_{T\rightarrow\infty}\limsup_{n\rightarrow\infty}\frac{1}{T}||\sum_{i=n}^{m(n,T)}\alpha_{i+1}{\bm{\epsilon}}_{i+1}||=0,$ $\mathds{P}^{*}$-a.s., where $m(n,T)=\max\{m:\sum_{i=n}^{m}\alpha_{i+1}\leq T\}$,
-  • A4
+  • S3
+    $\lim_{T\rightarrow\infty}\limsup_{n\rightarrow\infty}\frac{1}{T}||\sum_{i=n}^{m(n,T)}\alpha_{i+1}{\bm{\epsilon}}_{i+1}||=0,$ $\mathds{P}_{G^{*}}$-a.s., where $m(n,T)=\max\{m:\sum_{i=n}^{m}\alpha_{i+1}\leq T\}$,
+  • S4
     $\mathbf{h}$ is measurable and locally bounded.
@@
 $\mathscr{l}(\mathbf{g})=\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\log\left(\frac{p_{\mathbf{g}^{*}}(\mathbf{y})}{p_{\mathbf{g}}(\mathbf{y})}\right)p_{\mathbf{g}^{*}}(\mathbf{y}).$
-A1 holds by assumption; A3 holds since $({\bm{\epsilon}}_{n})$ is a martingale difference, and A4 is obvious.
-We now prove A2.
+S1 holds by assumption; S3 holds since $({\bm{\epsilon}}_{n})$ is a martingale difference, and S4 is obvious.
+We now prove S2.
 The set $J$ is closed and $\mathscr{l}$ is continuously differentiable.
@@
 $=1,$
-which leads to a contradiction. Hence A2 holds, and ${\rm d}(\mathbf{g}_{n},J)\rightarrow 0$ $\mathds{P}^{*}$-a.s., as $n\rightarrow\infty$.
-To conclude, we need to show that $\mathds{P}^{*}(\liminf_{n}{\rm d}(\mathbf{g}_{n},\partial\Delta^{D})=0)=0$.
-Since ${\rm d}(\mathbf{g}^{*},\partial\Delta^{D})>0$ and ${\rm d}(\mathbf{g}_{n},\{\mathbf{g}^{*}\}\cup\partial\Delta^{D})\rightarrow 0$ $\mathds{P}-a.s.$, then $\mathds{P}^{*}(\liminf_{n}{\rm d}(\mathbf{g}_{n},\partial\Delta^{D})=0)>0$ if and only if there exists $M$ with $\mathds{P}^{*}(M)>0$ such that ${\rm d}(\mathbf{g}_{n}(\omega),\partial\Delta^{D})\rightarrow 0$ for every $\omega\in M$, as $n\rightarrow\infty$.
+which leads to a contradiction. Hence S2 holds, and ${\rm d}(\mathbf{g}_{n},J)\rightarrow 0$ $\mathds{P}_{G^{*}}$-a.s., as $n\rightarrow\infty$.
+To conclude, we need to show that $\mathds{P}_{G^{*}}(\liminf_{n}{\rm d}(\mathbf{g}_{n},\partial\Delta^{D})=0)=0$.
+Since ${\rm d}(\mathbf{g}^{*},\partial\Delta^{D})>0$ and ${\rm d}(\mathbf{g}_{n},\{\mathbf{g}^{*}\}\cup\partial\Delta^{D})\rightarrow 0$ $\mathds{P}-a.s.$, then $\mathds{P}_{G^{*}}(\liminf_{n}{\rm d}(\mathbf{g}_{n},\partial\Delta^{D})=0)>0$ if and only if there exists $M$ with $\mathds{P}_{G^{*}}(M)>0$ such that ${\rm d}(\mathbf{g}_{n}(\omega),\partial\Delta^{D})\rightarrow 0$ for every $\omega\in M$, as $n\rightarrow\infty$.
 Let us make this assumption as a contradiction hypothesis.
@@
 By the contradiction hypothesis, $\mathcal{L}(\mathbf{g}_{n})(\omega)\rightarrow\infty$ as $n\rightarrow\infty$, for every $\omega\in M$.
-Then $E^{*}({\mathcal{L}}(\mathbf{g}_{n}))\rightarrow\infty$ as $n\rightarrow+\infty$, where $E^{*}$ denote expectation with respect to $\mathds{P}^{*}$.
+Then $E_{G^{*}}({\mathcal{L}}(\mathbf{g}_{n}))\rightarrow\infty$ as $n\rightarrow+\infty$, where $E_{G^{*}}$ denote expectation with respect to $\mathds{P}_{G^{*}}$.
 Let $\tilde{\mathbf{g}}_{n}=\eta_{n}\mathbf{g}_{n}+(1-\eta_{n})\mathbf{g}_{n+1}$, with the random variable $\eta_{n}$ taking values in $[0,1]$, be such that
@@
 Notice that
-$E^{*}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})\right)=\alpha_{n+1}E^{*}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}(\mathbf{h}_{i}(\mathbf{g}_{n})+{\bm{\epsilon}}_{n+1,i})\right),$
+$E_{G^{*}}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})\right)=\alpha_{n+1}E_{G^{*}}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}(\mathbf{h}_{i}(\mathbf{g}_{n})+{\bm{\epsilon}}_{n+1,i})\right),$
 with
-$E^{*}\left(\sum_{i=1}^{D}\frac{\mathbf{g}_{i}}{\mathbf{g}_{n,i}}{\bm{\epsilon}}_{n+1,i}\right)=0,$
+$E_{G^{*}}\left(\sum_{i=1}^{D}\frac{\mathbf{g}_{i}}{\mathbf{g}_{n,i}}{\bm{\epsilon}}_{n+1,i}\right)=0,$
 and
-$\liminf_{n}E^{*}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}\mathbf{h}_{i}(\mathbf{g}_{n})\right)=\liminf_{n}E^{*}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}\left(\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{k_{\mathbf{y}}({\bm{\vartheta}}_{i})\mathbf{g}_{n,i}p_{\mathbf{g}^{*}}(\mathbf{y})}{p_{\mathbf{g}_{n}}(\mathbf{y})}-\mathbf{g}_{n,i}\right)\right)$
-$=\liminf_{n}E^{*}\left(\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{p_{\mathbf{g}^{*}}(\mathbf{y})^{2}}{p_{\mathbf{g}_{n}}(\mathbf{y})}-1\right)$
-$>\liminf_{n}E^{*}\left(\left(\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{p_{\mathbf{g}_{n}}(\mathbf{y})}{p_{\mathbf{g}^{*}}(\mathbf{y})}p_{\mathbf{g}^{*}}(\mathbf{y})\right)^{-1}-1\right)=0,$
-where the strict inequality holds since, under the contradiction assumption, $\mathds{P}^{*}(p_{\mathbf{g}_{n}}\not\rightarrow p_{\mathbf{g}^{*}})>0$. Thus there exists a constant $c>0$ such that
-$E^{*}\left(-\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})\right)<-c\alpha_{n+1}$
-for $n$ large enough. On the other hand, denoting by $E^{*}_{n}$ the conditional expectation given $\mathcal{G}_{n}$, we can write that
-$E^{*}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{(\tilde{\mathbf{g}}_{n,i})^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}\right)$
-$\leq E^{*}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}\right)+E^{*}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n+1,i}^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}\right)$
-$\leq\alpha_{n+1}^{2}E^{*}\left(\sum_{i=1}^{D}\mathbf{g}^{*}_{i}\left(\frac{k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})}{p_{\mathbf{g}_{n}}(\mathbf{Y}_{n+1})}-1\right)^{2}\right)+\alpha_{n+1}^{2}E^{*}\left(\sum_{i=1}^{D}\mathbf{g}^{*}_{i}\frac{\left(\frac{k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})}{p_{\mathbf{g}_{n}}(\mathbf{Y}_{n+1})}-1\right)^{2}}{\left(1+\alpha_{n+1}\left(\frac{k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})}{p_{\mathbf{g}_{n}}(\mathbf{Y}_{n+1})}-1\right)\right)^{2}}\right)$
-$\leq\alpha_{n+1}^{2}E^{*}\left(1+E^{*}_{n}\left(\sum_{i=1}^{D}\left(\frac{k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})}{\sum_{i=1}^{D}k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})\mathbf{g}_{n,i}}\right)^{2}\right)\right)$
-$\quad\quad+\frac{\alpha_{n+1}^{2}}{(1-\alpha_{n+1})^{2}}E^{*}\left(1+E^{*}_{n}\left(\sum_{i=1}^{D}\left(\frac{k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})}{\sum_{i=1}^{D}k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})\mathbf{g}_{n,i}}\right)^{2}\right)\right)$
+$\liminf_{n}E_{G^{*}}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}\mathbf{h}_{i}(\mathbf{g}_{n})\right)=\liminf_{n}E_{G^{*}}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}\left(\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{k_{\mathbf{y}}({\bm{\vartheta}}_{i})\mathbf{g}_{n,i}p_{\mathbf{g}^{*}}(\mathbf{y})}{p_{\mathbf{g}_{n}}(\mathbf{y})}-\mathbf{g}_{n,i}\right)\right)$
+$=\liminf_{n}E_{G^{*}}\left(\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{p_{\mathbf{g}^{*}}(\mathbf{y})^{2}}{p_{\mathbf{g}_{n}}(\mathbf{y})}-1\right)$
+$>\liminf_{n}E_{G^{*}}\left(\left(\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{p_{\mathbf{g}_{n}}(\mathbf{y})}{p_{\mathbf{g}^{*}}(\mathbf{y})}p_{\mathbf{g}^{*}}(\mathbf{y})\right)^{-1}-1\right)=0,$
+where the strict inequality holds since, under the contradiction assumption, $\mathds{P}_{G^{*}}(p_{\mathbf{g}_{n}}\not\rightarrow p_{\mathbf{g}^{*}})>0$. Thus there exists a constant $c>0$ such that
+$E_{G^{*}}\left(-\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})\right)<-c\alpha_{n+1}$
+for $n$ large enough. On the other hand, denoting by $\mathcal{G}_{n}$ the sigma-algebra generated by $\mathbf{Y}_{1},\dots,\mathbf{Y}_{n}$, we can write that
+$E_{G^{*}}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{(\tilde{\mathbf{g}}_{n,i})^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}\right)$
+$\leq E_{G^{*}}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}\right)+E_{G^{*}}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n+1,i}^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}\right)$
+$\leq\alpha_{n+1}^{2}E_{G^{*}}\left(\sum_{i=1}^{D}\mathbf{g}^{*}_{i}\left(\frac{k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})}{p_{\mathbf{g}_{n}}(\mathbf{Y}_{n+1})}-1\right)^{2}\right)+\alpha_{n+1}^{2}E_{G^{*}}\left(\sum_{i=1}^{D}\mathbf{g}^{*}_{i}\frac{\left(\frac{k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})}{p_{\mathbf{g}_{n}}(\mathbf{Y}_{n+1})}-1\right)^{2}}{\left(1+\alpha_{n+1}\left(\frac{k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})}{p_{\mathbf{g}_{n}}(\mathbf{Y}_{n+1})}-1\right)\right)^{2}}\right)$
+$\leq\alpha_{n+1}^{2}E_{G^{*}}\left(1+E_{G^{*}}\left(\sum_{i=1}^{D}\left(\frac{k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})}{\sum_{i=1}^{D}k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})\mathbf{g}_{n,i}}\right)^{2}\mid\mathcal{G}_{n}\right)\right)$
+$\quad\quad+\frac{\alpha_{n+1}^{2}}{(1-\alpha_{n+1})^{2}}E_{G^{*}}\left(1+E_{G^{*}}\left(\sum_{i=1}^{D}\left(\frac{k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})}{\sum_{i=1}^{D}k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})\mathbf{g}_{n,i}}\right)^{2}\mid\mathcal{G}_{n}\right)\right)$
 $\leq 2\frac{\alpha_{n+1}^{2}}{(1-\alpha_{n+1})^{2}}\left(1+\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\sum_{i,{i^{\prime}}=1}^{D}\frac{k_{\mathbf{y}}({\bm{\vartheta}}_{i})^{2}}{(\sum_{{i^{\prime\prime}}}k_{\mathbf{y}}({\bm{\vartheta}}_{i^{\prime\prime}})\mathbf{g}_{n}({\bm{\vartheta}}_{i^{\prime\prime}}))^{2}}k_{\mathbf{y}}({\bm{\vartheta}}_{i^{\prime}})\mathbf{g}^{*}({\bm{\vartheta}}_{i^{\prime}})\right)$
@@
 It follows that,
-$E^{*}\left(\mathcal{L}(\mathbf{g}_{n+1})\right)\leq E^{*}\left(\mathcal{L}(\mathbf{g}_{n})\right),$
-for $n$ large enough, which contradicts $E^{*}(\mathcal{L}(\mathbf{g}_{n}))\rightarrow\infty$. ∎
+$E_{G^{*}}\left(\mathcal{L}(\mathbf{g}_{n+1})\right)\leq E_{G^{*}}\left(\mathcal{L}(\mathbf{g}_{n})\right),$
+for $n$ large enough, which contradicts $E_{G^{*}}(\mathcal{L}(\mathbf{g}_{n}))\rightarrow\infty$. ∎
 Lemma F.5.
 Under the assumptions of Theorem 2.4, for every $\delta<2-1/\gamma$,
-$||\mathbf{g}_{n}-\mathbf{g}^{*}||^{2}=o(n^{-\delta})\quad\mathds{P}^{*}-a.s.$
+$||\mathbf{g}_{n}-\mathbf{g}^{*}||^{2}=o(n^{-\delta})\quad\mathds{P}_{G^{*}}-a.s.$
 Proof.
 In this proof, $\mathbf{g}_{n}$ and $\mathbf{g}^{*}$ are considered as random vectors in $\{\mathbf{v}\in\mathbb{R}^{D}:||\mathbf{v}||\leq 1\}$.
-By Lemma F.4, $\mathbf{g}_{n}$, seen as vectors in $\Delta^{D}$ converge $P^{*}$-a.s.
+By Lemma F.4, $\mathbf{g}_{n}$, seen as vectors in $\Delta^{D}$ converge $P_{G^{*}}$-a.s.
 to $\mathbf{g}\in\Delta^{D}$.
@@
 The thesis follows from Theorem 3.1.1 in Chen (2005) if we can prove that
-  • B1
+  • T1
     $\alpha_{n}\rightarrow 0$, $\sum_{n=1}^{\infty}\alpha_{n}=\infty$, $\alpha_{n+1}^{-1}-\alpha_{n}^{-1}\rightarrow 0$,
-  • B2
-    $\sum_{n=1}^{\infty}\alpha_{n}^{1-\delta/2}{\bm{\epsilon}}_{n}$ converges, $\mathds{P}^{*}$-a.s.
-  • B3
+  • T2
+    $\sum_{n=1}^{\infty}\alpha_{n}^{1-\delta/2}{\bm{\epsilon}}_{n}$ converges, $\mathds{P}_{G^{*}}$-a.s.
+  • T3
     $\mathbf{h}$ is measurable and locally bounded, and is differentiable at $\mathbf{g}^{*}$. Let $H$ be a matrix such that, as $\mathbf{g}\rightarrow\mathbf{g}^{*}$,
@@
 Remark F.6.
-Theorem 3.1.1 in Chen (2005) also requests the following condition: B4 There exists a is continuously differentiable function $\mathscr{l}:\{\mathbf{v}\in\mathbb{R}^{D}:||\mathbf{v}||\leq 1\}\rightarrow\mathbb{R}$ such that $\sup_{\delta_{1}\leq{\rm d}(\mathbf{g},\mathbf{g}^{*})\leq\delta_{2}}\mathbf{h}(\mathbf{g})^{T}\nabla_{\mathbf{g}}\mathscr{l}<0$ for every $\delta_{2}>\delta_{1}>0$.
-However, a close inspection of the proof of Theorem 3.1.1 in Chen (2005) shows that B4 is employed only to ensure convergence of $\mathbf{g}_{n}$ towards $\mathbf{g}^{*}$.
+Theorem 3.1.1 in Chen (2005) also requests the following condition:
+T4 There exists a is continuously differentiable function $\mathscr{l}:\{\mathbf{v}\in\mathbb{R}^{D}:||\mathbf{v}||\leq 1\}\rightarrow\mathbb{R}$ such that $\sup_{\delta_{1}\leq{\rm d}(\mathbf{g},\mathbf{g}^{*})\leq\delta_{2}}\mathbf{h}(\mathbf{g})^{T}\nabla_{\mathbf{g}}\mathscr{l}<0$ for every $\delta_{2}>\delta_{1}>0$.
+However, a close inspection of the proof of Theorem 3.1.1 in Chen (2005) shows that T4 is employed only to ensure convergence of $\mathbf{g}_{n}$ towards $\mathbf{g}^{*}$.
 In our case, this property has been proved in Lemma F.4.
 B1 is obvious.
-To prove B2, notice that $(\sum_{k=1}^{n}\alpha_{k}^{1-\delta/2}{\bm{\epsilon}}_{k})_{n\geq 1}$ is a martingale.
+To prove T2, notice that $(\sum_{k=1}^{n}\alpha_{k}^{1-\delta/2}{\bm{\epsilon}}_{k})_{n\geq 1}$ is a martingale.
 Moreover, there exoists a constant $C$ such that, for every $i=1,\dots,D$,
-$\sup_{n}E^{*}((\sum_{k=1}^{n}\alpha_{k}^{1-\delta/2}{\bm{\epsilon}}_{k,i})^{2})$ $=\sup_{n}\sum_{k=1}^{n}\alpha_{k}^{2-\delta}E^{*}({\bm{\epsilon}}_{k,i}^{2})$
+$\sup_{n}E_{G^{*}}((\sum_{k=1}^{n}\alpha_{k}^{1-\delta/2}{\bm{\epsilon}}_{k,i})^{2})$ $=\sup_{n}\sum_{k=1}^{n}\alpha_{k}^{2-\delta}E_{G^{*}}({\bm{\epsilon}}_{k,i}^{2})$
 $\leq C\sum_{k=1}^{\infty}\left(\frac{\alpha}{\alpha+k}\right)^{\gamma(2-\delta)}<\infty,$
-since $\gamma(2-\delta)>1$ and $({\bm{\epsilon}}_{n,i})$ are uniformly bounded. We now prove B3. The function $\mathbf{h}$ is continuously differentiable at $\mathbf{g}^{*}$ and
+since $\gamma(2-\delta)>1$ and $({\bm{\epsilon}}_{n,i})$ are uniformly bounded. We now prove T3. The function $\mathbf{h}$ is continuously differentiable at $\mathbf{g}^{*}$ and
 $\mathbf{H}_{i,j}=-\mathbf{g}^{*}_{i}\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{k_{\mathbf{y}}({\bm{\vartheta}}_{i})k_{\mathbf{y}}({\bm{\vartheta}}_{j})}{p_{\mathbf{g}^{*}}(\mathbf{y})}.$